### v3

# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

Abstract.
This paper delves into the fundamental trade-off between security, latency, and throughput in proof-of-work longest-chain-fork-choice protocols, also known as the Nakamoto consensus. New upper and lower bounds on the probability of violating transaction safety are derived as a function of honest and adversarial mining rates, an upper bound on block propagation delays, and transaction confirmation latency, both in time and in block depth. The results include a first closed-form finite-latency bound applicable to all delays and mining rates up to the ultimate fault tolerance. Notably, the gap between the upper and lower bounds is significantly narrower than the best gaps previously established for Bitcoin and its derivatives such as Litecoin and Dogecoin, as well as for Ethereum Classic. Furthermore, the paper reveals a fundamental trade-off between transaction throughput and confirmation latency, ultimately determined by the desired fault tolerance and the growth of block propagation delay as block size increases.
The authors are with the Department of Electrical and Computer Engineering, Northwestern University, Evanston, Illinois. Email: {shujie.cao, dguo}@northwestern.edu.
copyright: none
1. Introduction
Nakamoto consensus, the celebrated consensus mechanism used in Bitcoin, centers around proof-of-work (PoW) mining and the longest-chain-fork-choice rule. In particular, a transaction can be committed once it is confirmed by a sufficiently long chain or after a sufficient amount of time. For example, a widely adopted rule of thumb is to accept a Bitcoin transaction after six confirmations.
The safety of a transaction may be violated due to forks caused by block propagation delays and/or adversarial mining. Suppose the honest miners collectively mine $h$ blocks per second and the adversary mines $a$ blocks per second. Suppose also once a block is known to one honest miner, it is known to all honest miners within $\Delta$ seconds. In 2020, Dembo et al. (dembo2020everything, ) and Gaži et al. (gazi2020tight, ) used very different methods to determine the ultimate asymptotic fault tolerance for the Nakamoto consensus. Specifically, as long as [^1] The condition is expressed as $\beta<(1-\beta)/(1+(1-\beta)\lambda\Delta)$ in (dembo2020everything, ), where $\lambda=a+h$ denotes the total mining rate and $\beta=a/(a+h)$ denotes the fraction of adversarial mining.
(1) $\frac{1}{a}>\Delta+\frac{1}{h},$
every transaction is safe with arbitrarily high probabilityafter sufficient confirmation; conversely, if $1/a<\Delta+1/h$, no transaction is ever safe. How safe is a transaction after a given confirmation latency? Gaži, Ren, and Russell (gazi2022practical, ) provided a safety guarantee via numerically evaluating a Markov chain, which is only practical for small confirmation latencies and parameters that are significantly lower than the fault tolerance. By analyzing races between renewal processes, Li, Guo, and Ren (li2021close, ) presented a closed-form result focusing on confirmation by time. Using a reduction technique, Guo and Ren (guo2022bitcoin, ) gave a simple closed-form latency–security guarantee focusing on confirmation by depth, which is numerically effective across a wide range of parameters, although there remains a gap between their fault tolerance and (1). In this paper, we put forth a continuous-time model for block mining processes, blockchains, and safety violation. We develop a new closed-form latency–security trade-off for the Nakamoto consensus which is the first to guarantee safety for all parameter values of $a$, $h$, and $\Delta$ within the ultimate fault tolerance (1). The new results here also substantially reduce the previous gap between upper and lower bounds on the latency-security trade-off in many scenarios of interest.
adversary new lower new upper (guo2022bitcoin, ) upper (gazi2022practical, ) upper
10% 6 7 7 7
20% 13 14 17 15
30% 32 35 42 N/A
40% 122 161 187 N/A
49.2% 26720 52807 $\infty$ N/A
Table 1. Bounds on confirmation depth for 99.9% safety. $a+h=1/600$ (blocks/second); $\Delta=10$ seconds.
In Table 1 we compare our new bounds to the best prior bounds on the minimum confirmation depth that guarantees 99.9% safety. The total mining rate is assumed to be Bitcoin’s 1 block every 600 seconds. The propagation delay bound $\Delta$ is set to 10 seconds. Our results encompass various adversarial mining fractions from $10\%$ to $49.2\%$. We include bounds for 10% and 20% adversarial mining taken directly from (gazi2022practical, ), which provides no result for higher adversarial mining rates because the numerical method therein is only feasible for parameters that are significantly distant from the ultimate fault tolerance. Evidently, the new upper bound is much closer to the lower bound than prior bounds. Moreover, the new bounds apply to all parameter values satisfying (1), which in this case is up to $a/(a+h)=49.8\%$. In contrast, the fault tolerance in (guo2022bitcoin, ) is $49.16\%$, hence for $49.16\%$, it is indicated by $\infty$ in Table 1. The inverse of the right hand side of (1) is known to be the minimum growth rate of honest blocks’ heights (pass2017analysis, ; dembo2020everything, ; gazi2020tight, ). If this growth rate exceeds the adversarial mining rate (as in (1)), the chance that more adversarial blocks are mined to overwhelm honest blocks to create a violation is very small for a confirmation latency sufficiently high. However, even if fewer adversarial blocks are mined than the height growth of honest blocks (i.e., the adversary has a deficit), a violation is still possible when honest miners with different views extend different chains to balance each other’s blocks on enough heights (i.e., the deficit is covered by those “balanced heights”). Examining the number of balanced heights offers a novel approach to the latency-security problem. Upon each opportunity to have a balanced height, there is a chance that all honest miners fail to balance, rather they synchronize to a common highest chain. Since honest blocks’ height growth rate exceeds the adversarial mining rate, there is a good chance that the adversary is left behind permanently, so no alternative highest chain is ever produced to cause a violation of the target transaction. Using this insight, we show that the number of balanced heights is stochastically dominated by essentially a geometric random variable. Interestingly, the aforementioned synchronizing block fall under the category of “Nakamoto blocks” defined in (dembo2020everything, ). Without using the complicated machinery of trees and tines as in (dembo2020everything, ; gazi2022practical, ), our analysis yields finite-latency, easily computable bounds that are tighter than the best existing results. In particular, the new bounds extend to the ultimate fault tolerance established in (dembo2020everything, ; gazi2022practical, ). The tight latency-security trade-off enables a user to choose an efficient confirmation latency that aligns with their desired security level. There has been ample evidence (e.g., (fechner2022calibrating, )) that the block propagation delay is approximately an affine function of the block size. In this paper, we explore the trade-off between the total throughput in kilobytes (KB) per second and the confirmation latency (in seconds), both of which depend on the block size and mining rate. This paper provides improved guidance for selecting parameters that optimize key performance indicators. The remainder of this paper is organized as follows: Sec. 2, introduces a mathematical model for the Nakamoto consensus. Sec. 3 explains why the model is sufficient for our purposes. Sec. 4 presents the main theorems. Sec. 5 introduces the most essential elements of the proof, followed by a sketch of the analysis. Sec. 6 presents the numerical results, showcasing the practical implications. Due to space limitations, detailed proofs are relegated to the appendices.
2. Model
In this section, we put forth a mathematical model for block mining processes and blockchains, enabling a precise definition of key concepts such as confirmation, latency, and safety violations. In Sec. 3, we will delve into why the model is sufficient for the latency-security analysis.
2.1. Block and Chain
We consider a continuous-time model, where transactions can be created and blocks can be mined at arbitrary times. Throughout this paper, “by time $t$” means “during $(0,t]$”.
Definition 1 (block).
The genesis block, also referred to as block 0, is mined first. Subsequent blocks are numbered sequentially in the order they are mined, starting from block 1, block 2, and so on. From block 1 onward, each block has a unique parent block, which is mined strictly before it. The parent of block 1 is the genesis block. We let $t_{b}$ denote the time when block $b$ is mined.
Definition 2 (transaction).
Transactions are created sequentially in time as transaction 1, transaction 2, and so on. Block $b$ is associated with an array $B_{b}$, which contains the indexes of some transactions that are created strictly before time $t_{b}$. We say transaction $j$ is included in block $b$ if $j$ is in the array $B_{b}$. A transaction can be included in multiple blocks.
Definition 3 (blockchain and height).
A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$, block $b_{i-1}$ is block $b_{i}$’s parent for every $i=1,\dots,n$, and no transaction is included in two or more of those blocks. Since every chain can be uniquely identified by its last block, we also refer to the preceding chain as chain $b_{n}$. As a convention, we say the height of this chain is $n$. In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$. Evidently, the genesis block is the only block on height 00.
[…]
An example of transactions, blocks, and chains.
Figure 1. An example of transactions, blocks, and chains. Nonpublic blocks are shaded. Each adversarial block is marked by a pair of horns on top. Below each block is an array of transactions included in the block.
Fig. 1 illustrates an example of chains formed by blocks 0 to 10 along with a time axis. It depicts a snapshot at $t_{10}$ when block $10$ is mined. Each block is horizontally positioned at its mining time and is connected by a line to its unique parent to its left and to its children to its right (if any). Under each block is its associated array describing the transactions included in the block. Chain $2$’s height is $h_{2}=2$; chain $5$’s height is $h_{5}=3$. We say chain $2$ is lower than chain $5$. Even though block $6$ is mined after block $5$, chain 5 is higher than chain $6$. Block $7$ or chain $7$ is the highest at its mining time. Block $1$ is associated with $B_{1}=[2,5]$, i.e., it includes transactions $2$ and $5$. Transaction $4$ is included in both blocks 2 and 3.
We next define two crucial attributes of transactions, blocks, and chains.
Definition 4 (public).
A transaction, block, or chain may become public at a specific point in timestrictly after its creation and will remain public at all times afterward. A chain being public at $t$ is equivalent to all of its blocks being public at $t$. While all transactions included in a public block are public, a transaction that is public at time $t$ may or may not be included in any block mined by $t$. As a convention, the genesis block becomes public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$.
Every chain consists of a public prefix chain followed by some nonpublic blocks (possibly none). In Fig. 1, the snapshot at $t_{10}$, we shade blocks which are nonpublic. Chain 5 is the public prefix of chain 9. If chain 9 becomes public at a later time, then block 7 must also be public by that time. Transaction 8 may be nonpublic at $t_{10}$, while transaction 11 is public since it is included in public block 6.
Definition 5 (credible).
A chain is said to be credible at time $t$, or $t$-credible, if it is no shorter than any public chain at $t$. All blocks in a credible chain are said to be credible.
Being public or not is an inherent attribute of a chain, while credibility is a different attribute determined by the set of public chains. Either public or not, a chain is $t$-credible if it is no shorter than any public chain at time $t$. In the snapshot at $t_{10}$ depicted in Fig. 1, chains 4 and 5 are both the highest public chains, making them both $t_{10}$-credible. Chain 6 is lower than chain 5, so it is not $t_{10}$-credible. Chains 7, 8, 9 and 10 are all $t_{10}$-credible. If chain 9 becomes public at a later time, then chain 8 will no longer be credible, regardless of whether it is public at that time.
2.2. Confirmation and Safety
Definition 6 (Confirmation by time $t$).
Let the confirmation time be fixed and denoted by $t>0$. Suppose a transaction becomes public at time $s>0$. We say the transaction is confirmed (by time $t$) by chain $c$ at time $s+r$ if chain $c$ is public and credible at $(s+r)$ with $r\geq t$ and contains a block which includes the transaction.
In Fig. 1, fix confirmation time $t$. Target transaction $4$ becomes public at $s$. Transaction $4$ (in block 2) is confirmed by chain 8 at $t_{10}$, which is $t_{8}$-credible, with $t_{8}>s+t$.
Definition 7 (confirmation by depth $k$).
Let the confirmation depth be a fixed natural number denoted by $k$. Suppose a transaction becomes public at $s>0$. Let block $d$ be the highest public block at time $s+\Delta$. If there exists a credible public chain $c$, which contains a block $b$ that includes the transaction, where $h_{c}\geq h_{b}+k-1$ and $h_{c}\geq h_{d}+k$, then we say chain $c$ confirms the transaction by depth $k$.
Fix $k=3$. In Fig. 1, transaction 4 is included in block 2 ($h_{2}=2$), which is contained in chain 8 ($h_{8}=4$). The highest public block by $s$ is block 1 with $h_{1}=1$ (assuming block 2 is not yet public at $s$). Since chain 8 is public and credible at $t_{10}$, $h_{8}\geq h_{2}+k-1$, and $h_{8}\geq h_{2}+k$, we say transaction 4 is confirmed (by depth $k$) by chain 8.
With confirmation by either time or depth, we can define the violation of a target transaction’s safety as follows:
Definition 8 (safety violation of a target transaction).
Suppose a target transaction becomes public at $s>0$. If there exist $r$ and $r^{\prime}$ with $r^{\prime}\geq r$ and also:
  • a chain $c$ that is public and credible at time $r$ which confirms the transaction in block $b$, and
  • an $r^{\prime}$-credible chain which does not contain block $b$,
then we say the target transaction’s safety is violated.
In Fig. 1, target transaction $4$’s safety is violated as it is confirmed by $t_{10}$-credible and public chain 8 (in block 2), while $t_{10}$-credible chain 7 does not contain block 2. Note that if two credible chains confirm the same target transaction but in different blocks, its safety is still violated.
2.3. Adversary and Propagation Delay
Definition 9 (types of blocks).
A block is either an A-block or an H-block. The parent of an H-block must be credible when the H-block is mined.
An A-block (resp. H-block) is a mathematical concept that represents a block mined by an adversarial (resp. honest) miner. According to Definitions 5 and 9, an H-block must be higher than all public blocks at the time of its mining. There is no such height constraint on an A-block. In fact, this is the only constraint on block heights. In Fig. 1, we distinguish each A-block using a pair of horns on top. Blocks 2 and 7 are A-blocks, while all the other blocks are H-blocks.
Definition 10 (transaction’s inclusion).
Suppose a transaction becomes public at time $s$. If block $b$ is the first H-block mined after $s$, then chain $b$ must include the transaction.
The propagation delays of transactions, blocks, and chains depend on the network conditions and may also be manipulated by the adversary. To establish a strong safety guarantee, we adopt the widely used $\Delta$-synchrony model.
Definition 11 (chain delay bound $\Delta$).
If block $b$ (mined at $t_{b}$) is an H-block, then chain $b$ must be public at all times strictly [^2] The H-block $b$ is nonpublic at $t_{b}$ according to Definition 4. In the degenerate case of $\Delta=0$, it becomes public right after $t_{b}$. after $t_{b}+\Delta$.
Definition 10 guarantees that once a transaction becomes public, it will be promptly included in at least one chain. Together, Definitions 10 and 11 imply that, if a transaction becomes public at $s$, no H-block mined before time $s-\Delta$ includes it, while it may be included in an A-block mined before $s-\Delta$.
Definition 12 (Poisson mining processes).
A-blocks are mined according to a Poisson point process with rate $a$ (blocks per second). H-blocks are mined according to a Poisson point process with rate $h$ (blocks per second). The two Poisson mining processes are independent.
Let $A_{s,t}$ denote the number of A-blocks mined during $(s,t]$. Let $H_{s,t}$ denote the number of H-blocks mined during $(s,t]$. Evidently, $(A_{0,t}+H_{0,t})_{t\geq 0}$ is a Poisson point process with rate $a+h$. As a convention, let $A_{s,t}=H_{s,t}=0$ for all $s>t$. Since the probability that any two blocks are mined at exactly the same time is equal to zero, we assume without loss of generality that all blocks are mined at different times. Hence $t_{0}<t_{1}<t_{2}<\dots$.
3. Model Sufficiency
3.1. Sufficiency
The prevailing model in this line of work (see, e.g., (garay2015bitcoin, ; pass2017analysis, ; pass2017rethinking, ; garay2017bitcoin, ; dembo2020everything, ; gazi2020tight, ; saad2021revisiting, ; neu2022longest, ; gazi2023practical, ; kiffer2023security, )) is based on the universally composable security framework (canetti2000universally, ). A complete description of the model, henceforth referred to as the view-based model, requires the definition of a protocol as a set of algorithms, the execution of those algorithms that is directed by an environment, which activates $n$ either honest or corrupted parties with individual views (pass2017analysis, , Sec. 2). The protocol’s execution is no simpler than a system of $n$ interactive Turing machines. Due to its complexity, the view-based model always assumes discrete time steps in execution.
The model proposed in Sec. 2 of this paper is entirely self-contained and much simpler than the view-based model. In contrast to previous works as well as (li2021close, ; guo2022bitcoin, ), our model intentionally exclude miners and their views in definitions. In particular, with well-defined block mining processes, we are not concerned with the number of miners; neither do we keep track of which blocks an individual miner has mined or seen. Instead, our model keeps track of the set of public chains at every point in time, which represents the weakest possible view of any miner. In particular, every honest view is a superset of this weakest view, and conversely, the adversary in the view-based model can ensure at least one honest miner holds that weakest view.
Relative to the view-based model under $\Delta$-synchrony, our model has no loss of generality in the sense that, for a given realization (or sample path) of the A- and H-block mining processes, if the view-based model can produce a certain set of chains, so can our model. This is because both models respect the same constraints for chain creation described in Definitions 9 and 11. For the latency-security analysis, our model is as powerful as the view-based model (if simpler). To elaborate, consider the event in the view-based model where, first, a chain in an honest view confirms a block on height $h$, and then another chain in an honest view confirms a different block on the same height, creating a safety violation. Our model can produce the same pair of chains, each of which is credible to the weakest view at the time it confirms a conflicting block. Conversely, if our model produces a pair of conflicting credible chains, then those chains can be used to convince two honest parties in the view-based model to confirm different blocks on the same height. Once we realize the role of the weakest view, there is no need to define individual views and their complicated interactions.
3.2. Operations
For concreteness and practicality, we relate the definitions in Sec. 2 to the operations of miners (or nodes), who assemble transactions into blocks and chains by expending proof of work. The same miners also propagate those chains through a peer-to-peer network. The honest miners follow the longest-chain-fork-choice rule whereas the adversary can use an arbitrary (possibly randomized) strategy. At time $t$, each node has its own view, which includes all transactions, blocks, and chains the node has seen by $t$. A node’s view can only increase with time. The operations of each node is based on the entire history of this node’s view.
In general, an honest node will only provide goods and services for a transaction after it is considered committed. If an honest node first encounters a transaction at time $s$, it can deduce that the transaction will become public during the time window of $[s,s+\Delta]$. Here we define two commit rules corresponding to the two types of confirmation defined in Sec. 2. Specifically, if the transaction is included in a highest chain in this node’s view at time $s+r$ for some $r\geq t$, and the chain remains the highest at $s+r+\Delta$, then the node can infer that 1) the chain is credible and public at $s+r+\Delta$, and 2) the transaction has been confirmed by time $t$ in the mathematical model. In this case, the node commits the transaction by time $t$. Similarly, if block $d$ is the highest block in this node’s view at time $s+\Delta$, and there is a chain $c$ in the node’s view that remains the highest for a period of $\Delta$ or longer, which contains a block $b$ including the transaction, and $h_{c}\geq h_{b}+k-1$ as well as $h_{c}\geq h_{d}+k$, then the node can infer that the transaction has been confirmed by depth $k$ in the mathematical model. In this case, the node commits the transaction by depth $k$.
We will show that, regardless of the adversary’s strategy, violating the safety of a transaction implies that an “error event” concerning the two mining processes occurs. We then bound the probability of safety violation defined in Sec. 2, which implies a probabilistic guarantee for an honest node’s operational safety.
3.3. Delay Variations
The seemingly rigid $\Delta$-synchrony model is quite practical. First, “light” nodes do not mine blocks or fork chains, so the delays they perceive are inconsequential to ledger safety as long as honest miners satisfy the delay requirement. Second, if some blocks of honest miners suffer longer delays than $\Delta$, we simply regard those blocks as adversarial blocks to preserve $\Delta$-synchrony. In addition, potential attacks like Erebus (Erebus, ) and SyncAttack (SyncAttack, ) seek to partition the peer-to-peer network. As long as those delay violations are independent of the block mining times, the honest and adversarial mining processes remain to be Poisson. If block propagation delays have a long-tailed distribution (decker2013information, ; fechner2022calibrating, ), we can use a delay bound that applies to a high percentage of blocks. The main results here continue to apply with all out-of-sync mining power treated as being adversarial.
In Definition 10, we have assumed that every public transaction gets included in a chain very soon. In a more general model, a transaction’s inclusion can be delayed due to censorship or a lack of space in new blocks. If this is the case, we modify the confirmation rules as follows: First, block $b$ which contains the target transaction becomes public. Then, if block $b$ is in a credible public chain $c$ with $h_{c}\geq h_{b}+k-1$, we say chain $c$ confirms the transaction by depth $k$. If block $b$ is in a credible public chain $c$ at time $t_{b}+r$ with $r\geq t$, we say chain $c$ confirms the transaction by time $t$. We note that the main theorems in the next section can be easily generalized to this case, although we omit the details in this paper.
4. Main Results
Recall that $a$ and $h$ represent as the adversarial and honest mining rates, respectively, in the number of blocks mined per second. For confirmation by either time or depth, the following negative result is well understood (see, e.g., (pass2017analysis, ; dembo2020everything, ; gazi2020tight, )).
Theorem 1 (unachievable fault tolerance).
If the mining rates and the delay upper bound satisfy
(2) $\frac{1}{a}<\Delta+\frac{1}{h},$
then every transaction is unsafe in the sense that the private-mining attack on the transaction will violate its safety with probability 1.
For completeness, we prove Theorem 1 in Appendix A.
4.1. Bounds for Given Confirmation Depth
Let us define
(3) $b(\alpha,\eta)=\left(\frac{e^{4\gamma}\gamma(\eta-\alpha-\alpha\eta)}{\alpha\eta e^{\gamma}-(\alpha+\gamma)(\eta-\gamma)}\right)^{2}+1-\frac{\eta-\alpha-\alpha\eta}{\eta e^{\alpha}},$
and
(4) $\begin{split}c&(\alpha,\eta)=\left(\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma\right)\\ &\cdot\left(1-\ln\left(1+\frac{\gamma}{\alpha}\right)\bigg{/}\ln\frac{\eta e^{\alpha}(e^{\eta}-1)}{\eta+\eta e^{\alpha}(e^{\eta}-1)-\alpha(1+\eta)}\right)^{-1}\end{split}$
where $\gamma$ in (3) and (4) is a shorthand for
(5) $\gamma(\alpha,\eta)=\frac{2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}}{2}.$
The following latency–security trade-off is the first set of non-trivial bounds that apply to all parameter values up to the ultimate fault tolerance.
Theorem 2 (Upper bounding the violation probability for given confirmation depth).
Suppose the mining rates and the delay upper bound satisfy
(6) $\frac{1}{a}>\Delta+\frac{1}{h}.$
Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
(7) $b(a\Delta,h\Delta)e^{-c(a\Delta,h\Delta)\cdot k}$
where $b(\cdot,\cdot)$ and $c(\cdot,\cdot)$ are defined in (3) and (4), respectively.
Theorem 2 is proved in Appendix D, in which the following result is established as a byproduct in Appendix F.
Corollary 3 (Achievable fault tolerance).
As long as (6) holds, every transaction is arbitrarily safe after sufficiently long confirmation latency (either in depth or time).
The pair of opposing inequalities (2) and (6) determine the exact boundary between the asymptotically achievable and unachievable latency–security trade-offs. The right hand side of (2) and (6) is exactly the expected inter-arrival time of the following type of blocks:
Definition 13 (Pacer).
We say the genesis block is the 00-th pacer. After the genesis block, every pacer is the first H-block mined at least $\Delta$ time after the previous pacer. We use $P_{s,t}$ to denote the number of pacers mined in $(s,t]$.
By definition, pacers become public by their immediate successors and hence all pacers are on different heights. Since the expected pacer inter-arrival time is $\Delta+1/h$, pacers arrive at the rate of $h/(1+h\Delta)$ in the long run. As a consequence, the number of heights H-blocks occupy grows at least at that rate. This approach enables us to use the growth rate of pacers as a lower bound for the height increase of H-blocks.
Let us introduce the following probability density function (pdf):
(8) $f(t)=\frac{1}{h}e^{-h(t-\Delta)}1_{\{t>\Delta\}}.$
We also consider a delayed renewal process in general, which is like an ordinary renewal process, except that the first arrival time (delay) may have a different distribution than the other inter-arrival times.
Lemma 4 ((Kroese_difference, , Theorem 3.4)).
Let $A_{r}$ denote a Poisson point process with rate $a$. Let $Q_{r}$ denote a renewal process whose inter-arrival time has a pdf given by (8). Then the probability generating function of the maximum difference between the two processes
(9) $Y=\sup_{t\geq 0}\{A_{t}-Q_{t}\}$
is equal to
(10) $\mathcal{E}(r)=\frac{(1-r)(h-a-ha\Delta)}{h-e^{(1-r)a\Delta}(h+a-ar)r}.$
We denote the probability mass function (pmf) of $Y$ as
(11) $e(i)$ $=\frac{1}{i!}\mathcal{E}^{(i)}(0),\qquad i=0,1,\dots$
where $\mathcal{E}^{(i)}(0)$ represents the the $i$-th derivative of $\mathcal{E}(\cdot)$ evaluated at $r=0$.
Throughout this paper, we let
(12) $\epsilon$ $=1-e^{-h\Delta}$
and
(13) $\delta$ $=\left(1-a\frac{1+h\Delta}{h}\right)^{+}e^{-a\Delta}.$
Let us define
(14) $F^{(k)}(n)=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}\bar{F}_{b}(n;k,(\epsilon+\delta-\epsilon\delta))$
where $\bar{F}_{b}(\cdot;k,p)$ denotes the complementary cumulative distribution function (cdf) of the binomial distribution with parameters $(k,p)$. We also define
(15) $F(n)$ $=\lim_{k\to\infty}F^{(k)}(n)$
(16) $=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}.$
Theorem 5 (Finer upper bound for the violation probability for given confirmation depth).
Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
$\inf_{n}\bigg{\{}2-$ $\sum_{i=0}^{k-n}e(i)\sum_{j=0}^{k-n-i}e(j)\int_{0}^{\infty}F_{1}(k-n-i-j;a(t+6\Delta))$
(17) $\cdot f_{2}(t-(k)\Delta;k,h)\,dt-F^{(k)}(n-1)\bigg{\}}$
where $F^{(k)}$ is defined in (14), $e(i)$ is given by (11), $F_{1}(\cdot;\lambda)$ denotes the cdf of the Poisson distribution with mean $\lambda$, and $f_{2}(\cdot;k,h)$ denotes the pmf of the Erlang distribution with shape parameter $k$ and rate parameter $h$.
Theorem 5 is proved in Appendix G.
To ease computation, the infinite integral in (5) could be replaced by only integrating to a finite time limit, where the larger the limit, the tighter the upper bound.
A well-understood attack is the private-mining attack, which we will use to lower bound the probability of safety violation. The attack is defined precisely as follows.
Definition 14 (Private-mining attack on a target transaction).
The adversary delays the propagation of every H-block maximally (for $\Delta$ seconds). Starting from time 0, the adversary always mines in private to extend the highest chain which does not include the target transaction. The adversary makes public the private chain as soon as both of the following conditions hold: (1) the transaction has been confirmed by a chain of H-blocks; and (2) the private chain which does not contain the transaction is credible.
Theorem 6 (Lower bounding the violation probability for given confirmation depth).
Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
[…]
$\cdot\sum_{j=0}^{k-i}e(j)\bigg{(}f_{1}(0,h\Delta)\int_{0}^{\infty}F_{1}(k-(1+i+j);a(t-\Delta))$
$\cdot f_{2}(t-k\Delta;k,h)\,dt+(1-f_{1}(0,h\Delta))$
(18) $\cdot\int_{0}^{\infty}F_{1}(k-(i+j);a(t-\Delta))\cdot f_{2}(t-k\Delta;k,h)\,dt\bigg{)}$
where $e(\cdot)$, $F_{1}$,$f_{2}$ are defined as in Theorem 5 and $f_{1}(\cdot;\lambda)$ denotes the pmf of the Poisson distribution with mean $\lambda$.
Theorem 6 is proved in Appendix H.
4.2. Throughput and Related Trade-offs
An ideal protocol design should select parameters to yield high throughput, fast confirmation, and security against adversary miners. We refer to the average number of blocks mined during the maximum propagation delay, $\lambda\Delta$, as the natural fork number. Throughout the paper, let $\beta$ denote the maximum tolerable fraction of adversarial mining. The natural fork number is bounded by
(19) $\lambda\Delta<\frac{1}{\beta}-\frac{1}{1-\beta}$
due to the fault tolerance (1).
To achieve a high throughput (in KB/second or transactions/second), the designer may select a high mining rate and/or a large block size, where the latter also implies a large block propagation delay. The optimistic throughput, where all miners work to grow the highest chains and natural forks are orphaned, is equal to $\lambda B/(1+\lambda\Delta)$, where $B$ represents the maximum allowable size of a block, ensuring that the propagation delay does not exceed the upper limit $\Delta$. The following result is a straightforward consequence of (19).
Theorem 7.
If the blocks are no larger than $B$ KB, the block propagation delays can be as large as $\Delta$ seconds, in order to tolerate adversarial mining fraction up to $\beta$, the throughput cannot exceed
[…]
Empirically, the block propagation delay is approximately an affine function of the block size (decker2013information, ; fechner2022calibrating, ) (even with a coded design (zhang2022speeding, )). That is, the propagation delay bound can be modeled as
(21) $\Delta(B)=\frac{B}{r}+\nu$
where $B$ is the block size in KB and $r$ and $\nu$ are constants. Here, $r$ represents a fundamental limitation on the end-to-end transmission rate of the peer-to-peer network. Based on (20), this rate and the desired fault tolerance determines an upper bound for the optimistic throughput:
(22) $\frac{\lambda B}{1+\lambda\Delta}<\frac{1-2\beta}{1-\beta-\beta^{2}}\,r.$
With a close upper bound on the probability of safety violation as a function of the system parameters and the confirmation latency, e.g., in the form of $p(\lambda\Delta,\beta,k)$ where a user may select the confirmation depth $k$ to achieve a desired safety. It is well understood that a user making a very large transaction is generally willing to tolerate a longer confirmation latency. The theorems in this paper allow this choice to be made much more accurately than before.
It is also meaningful to consider optimizing the throughput by selecting parameters that guarantee a certain safety at a desired confirmation latency (in time) under a worst-case assumption of adversarial mining. In particular, we can formulate the following problem:
(23a) $\text{maximize}_{\lambda,B,k}\quad$ $\frac{\lambda B}{1+\lambda\,\Delta(B)}$
(23b) subject to $p(\lambda\,\Delta(B),\beta,k)\leq q$
(23c) $k\leq\frac{(1-\beta)\lambda}{1+(1-\beta)\lambda\Delta}d$
where $q$ denotes the maximum tolerable probability of violation and $d$ denotes the expected confirmation time. Here (23c) constrains $k$ such that at the slowest growth rate, gaining $k$ heights takes no more than $d$ units of time on average. In Sec. 6, we give such a numerical example under assumption (21) with empirically determined constants.
4.3. Bounds for Given Confirmation Time
Theorem 8 (Upper bound for given confirmation time).
Suppose the mining rates and the delay upper bound satisfy (1). Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by time $t$ can be violated with a probability no higher than
(24) $\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+G(n,t)\},$
where $F$ is defined in (15), and
(25) $\begin{split}&G(n,t)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ &\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
where $e(\cdot)$ and $f_{1}$ are defined the same as in Theorem 6, and $F_{2}(\cdot;n,h)$ denotes the cdf of the Erlang distribution with shape parameter $n$ and rate parameter $h$.
Theorem 8 is proved in Appendix I. In lieu of the infinite sums in (25), a computationally feasible bound is obtained by replacing $G(n,t)$ with its following upper bound
(26) $\begin{split}\overline{G}(n,t)&=1-F_{1}(i_{0};a(t+2\Delta))F_{E}(j_{0})F_{E}(k_{0})\\ &+\sum_{i=0}^{i_{0}}\sum_{j=0}^{j_{0}}\sum_{k=0}^{k_{0}}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ \times(1-&F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
where $F_{1}$, $F_{E}$ are the corresponding cdfs of variables with pmf of $f_{1}(i;a(t+2\Delta))$, $e(j)$ in (11), respectively. And $i_{0}$, $j_{0}$ and $k_{0}$ are positive integers. The larger these integers are, the tighter the upper bound.
For completeness, we include the following converse result, which is a variation of (li2021close, , Theorem 3.6) due to the slightly different definition of safety violation here.
Theorem 9 (Lower bound for given confirmation time).
Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
(27) $\begin{split}&\left(1-\frac{a}{h}\right)e^{(h-a)(t+2\Delta)}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}e(j)\left(\frac{a}{h}\right)^{k}F_{1}(k;h(t+2\Delta))\\ &\;\times(1-F_{2}(t+\Delta-(j+k)\Delta;j+k+1,h))-e^{-h(t+2\Delta)}.\end{split}$
Theorem 9 is proved in Appendix J. Evidently, if the infinite sums in (27) are replaced by partial sums, the resulting (tighter) security level also remains unachievable.
5. Essential Elements of the Analysis
In this section, we introduce the most essential technical elements, discuss intuitions, and provide a sketch of the statistical analysis.
5.1. The Balancing Process
Definition 15 (Agree and disagree).
Two chains are said to agree on a target transaction if they both contain a common block which includes the transaction, or neither includes the transaction; otherwise, they are said to disagree on the target transaction. We also say blocks $b$ and $d$ agree (resp. disagree) on the target transaction if chains $b$ and $d$ agree (resp. disagree) on the target transaction.
Agreement about a target transaction is an equivalence relation, which typically divide all blocks into two equivalence classes. In Fig. 1, however, because transaction 4 is included in both blocks 2 and 3, its agreement divides all blocks into three equivalence classes, namely, $\{0,1,6\}$, $\{2,4,8,10\}$, and $\{3,5,7,9\}$.
Definition 16 (Balancing candidate, candidacy).
Given a target transaction, if an H-block (denoted as block $b$) is the first block mined on height $n$ and disagrees on the target transaction with any height-$(n-1)$ block mined before block $b$ becomes public, then block $b$ is said to be a balancing candidate, or simply a candidate. The period from the first time the disagreement arises to the time block $b$ becomes public is referred to as its candidacy. Evidently, if any disagreeing height-$(n-1)$ block is mined before block $b$, its candidacy begins from $t_{b}$.
Definition 17 (Balanced height).
If during a balancing candidate’s candidacy, a disagreeing H-block is mined on the candidate’s height and no adversarial block is mined on that height, then the candidate is said to be balanced. Accordingly, its height is said to be a balanced height.
We use $X_{s,t}$ to denote the total number of balanced candidates whose candidacy is finished during $(s,t]$. With slight abuse notation, we let
(28) $X=\lim_{t\to\infty}X_{0,t}.$
[…]
Figure 2. Two balancing candidates $b$ and $b^{\prime}$ at different times. Blocks on the same height are connected by dashed lines. Left: $b$ becomes a balancing candidate upon its mining; Right: $b^{\prime}$ becomes a candidate after $c$ is mined on a lower height to create a disagreement. The duration of each candidacy is marked red on the time axis. The duration of the first of those candidacies is exactly $\Delta$. The duration of the second one is strictly less than $\Delta$.
A candidacy is a small window of opportunity to balance the corresponding candidate (and thus create a balanced height). Figs. 2 illustrates two different balancing candidacies at different times. Examples of balancing can be found in Fig. 1. With regard to target transaction 4, blocks 4 and 5 balance each other on height 3; blocks 9 and 10 balance each other on height 5. No other heights are balanced. In particular, while H-blocks 3 and 6 are on the same height, they do not balance each other due to the existence of block 2 as an A-block on the same height. We shall show that the total number of heights that can be balanced (denoted as $X$) is essentially stochastically dominated by a geometric random variable.
5.2. Violation Event under Confirmation by Depth
Definition 18 (Pre-mining lead).
The pre-mining lead (of the adversary) at time $t$, denoted by $L_{t}$, is the height of the highest block mined by time $t$ minus the height of the highest H-block mined by time $t$.
The notion of pre-mining lead is to account for the potential lead the adversary may already have when the target transaction becomes public at around time $s$. In this work, $s$ is given so that the adversary cannot mount an adaptive attack (by influencing $s$) depending on the lead. Nonetheless, the adversary can try to create and maintain some lead (if possible), which is essentially characterized by a birth-death process.
Definition 19 (Jumper).
We say the genesis block is the 00-th jumper. After the genesis block, every jumper is the first H-block mined on its height. We use $J_{s,t}$ to denote the number of jumpers mined in $(s,t]$.
In the remainder of this section, we assume $s$ is a given time. We focus on the case of confirmation by depth $k$.
It is easy to see that the height growth of H-blocks is no less than the number of jumpers mined. Intuitively, for the transaction’s safety to be violated, there must exists an interval, during which the number of jumpers mined is fewer than the sum of the lead, the number of balanced heights, and the number of A-blocks mined. This is made precise in the following lemma.
Let $\tau$ denote the first time a block is mined that is $k$ or more blocks higher than the highest public block at time $s+3\Delta$.
Lemma 1.
Suppose a target transaction becomes public during $(s,s+2\Delta]$. If this target transaction’s safety is violated, then the following event must occur:
(29) $L_{s}+X_{s,\tau}+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k.$
Proof.
Suppose target transaction is public at time $s^{\prime}$, block $b^{\prime}$ is the highest public block at time $s^{\prime}+\Delta$. Let $\tau^{\prime}$ denote the first time a block is mined at height $h_{b^{\prime}}+k$ or above. The highest public H-block at time $d$ is at least $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$. If a violation occurs at time $d$, there must exist two chains disagreeing with each other on the target transaction, where both chains are no shorter than $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$. At time $s^{\prime}$, since the highest H-block is no higher than $h_{b^{\prime}}$, the highest A-block is no higher than $h_{b^{\prime}}+L_{s^{\prime}}$. Since $X_{s^{\prime},d}$ heights are balanced, A-blocks need to be mined at least on the remaining $k-X_{s^{\prime},d}$ heights after $h_{b^{\prime}}+L_{s^{\prime}}$. Along with A-blocks mined during $(s^{\prime},d]$, a violation implies the following,
(30) $h_{b^{\prime}}+L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}$ $\geq h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}.$
By Definition 18 of the pre-mining lead,
(31) $L_{s}+X_{s,d}+A_{s,d}\geq L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}.$
By definition of $\tau$, we have
(32) $k+J_{\tau^{\prime},d-\Delta}\geq k+J_{\tau,d-\Delta}.$
One way a violation occurs is for a balancing candidate to arise (regardless of whether the height is subsequently balanced). Thus, if a violation occurs at time $d>\tau^{\prime}$, we must have
(33) $X_{\tau^{\prime},d}=0.$
Thus, we have
(34) $L_{s}+X_{s,\tau}+A_{s,d}\geq k+J_{\tau,d-\Delta}.$
Therefore, a violation can only occur if (29) occurs. ∎
Based on Lemma 1, we can characterize the latency-security trade-off by examining the competition between adversarial and honest nodes. While mining processes are completely memoryless, the height growths of honest and adversarial blocks are generally intertwined in a complex manner. In particular, when honest nodes have split views, the adversary may utilize some honest blocks to construct or maintain a competing chain, thereby perpetuating the split views among honest nodes. In general, the random variables in (29) are dependent. The challenge we must address is to decompose the event in Lemma 1 into separate, independent components.
5.3. A Sketch of the Statistical Analysis
The remaining task is to analyze the probability of the event described by (29). We need to understand the distribution of the individual elements as well as to account for their dependency.
Consider the balanced heights first. As a convention, let $(x)^{+}=\max\{0,x\}$.
Theorem 2 (Bounding the number of balanced heights).
For every integer $n\geq 0$, the number of balanced heights up to the $k$-th candidacy, denoted as $X^{(k)}$, satisfies
(35) $P(X^{(k)}\leq n)\geq F^{(k)}(n)$
where $F^{(k)}$ is defined in (14). As a special case, the total number of balanced heights defined in (28) satisfies
(36) $P(X\leq n)\geq F(n)$
where
If (2) holds, then $\delta=0$, so $F(n)=0$ and (36) is vacuous.
Theorem 2 is proved in Appendix B. The key insight here is that the adversary can only take advantage of a bounded number of H-blocks to balance each other even in the long run. In fact, with probability 1, one last balancing opportunity will be missed and the adversary is then behind future pacers permanently. The final balancing candidate serves the role of the “Nakamoto block” in (dembo2020everything, ) and is a stronger notion than the “convergence opportunity” in e.g., (pass2017rethinking, ). This observation underlies the safety guarantee of Theorem 2.
Let us define
(37) $Y^{(k)}=L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}.$
While $Y^{(k)}$ and $X_{s,\tau}$ are dependent, we can upper bound the probability of the event (29), i.e., $X_{s,\tau}+Y^{(k)}\geq 0$, using a union bound:
Lemma 3.
Let $\bar{F}_{X}(\cdot)$ denote the complementary cdf of the random variable $X$. All integer-valued random variables $X$ and $Y$ satisfy
(38) $P(X+Y\geq 0)\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X}(n-1)+\bar{F}_{Y}(-n)\}.$
In addition, given two distributions of integer-valued random variables for which the right-hand side of (38) is strictly less than 1, there exists a joint distribution of $X$ and $Y$ that satisfy (38) with equality.
The proof of Lemma 3 is relegated to Appendix C.
It suffices to upper bound the marginal complementary cdf’s of $X_{s,\tau}$ and $Y^{(k)}$. Since all $X_{s,\tau}$ balanced heights must be within the first $k$ heights above $h_{b}$, we have
(39) $P(X_{s,\tau}>n-1)\leq P(X^{(k)}>n-1)$
where the distribution of $X^{(k)}$ is given in (14). Thus, we have
$P(Y^{(k)}$ $+X_{s,\tau}\geq 0)$
(40) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X_{s,\tau}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}$
(41) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X^{(k)}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}.$
The remaining task is to understand the distribution of $Y^{(k)}$. A challenge here is that the jumper process depends on the H-block process. To decouple the dependence, we use the pacer process to lower bound the jumper process:
Lemma 4 (There are more jumpers than pacers).
(42) $J_{s,t}\geq P_{s+\Delta,t},\quad\forall t>s.$
Proof.
We first show that for every pacer block $p$, there must exist a jumper mined during $(t_{p}-\Delta,t_{p}]$. This claim becomes trivial if block $p$ is the first H-block mined on its height; otherwise, the first H-block mined on $h_{p}$ must be mined strictly after $t_{p}-\Delta$. Because pacers are mined at least $\Delta$ time apart, the lemma is established. ∎
By Lemma 4, we have
(43) $Y^{(k)}$ $\leq L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-P_{\tau+\Delta,d-\Delta}\}$
(44) $=L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P_{\tau+\Delta,\tau+\Delta+c}\}.$
It is easy to see that the lead $L_{s}$, the A-block process from $s$ onward, and the pacer process from $\tau+\Delta$ onward are mutually independent. Also, the supremum in (44) can be viewed as the extremum of the difference of two renewal processes, whose distribution is readily available from Lemma 4.
Once the distribution of $Y^{(k)}$ is bounded, we can combine it with Theorem 2 to upper bound the probability of violation using (41). A detailed proof of Theorem 2 is given in Appendix D.
5.4. Operational Implications
A brief discussion on the operational implications of the theorems presented in Sec. 2 is warranted. First, the requirement for “a sufficiently large $s$” in the theorems serves the purpose of ensuring that the maximum random pre-mining lead is in a steady state. This assumption is easily applicable in most practical scenarios.
More importantly, we define $s$ as “given” in the sense that it is not influenced by mining processes. Specifically, it is not selected by the adversary. If the adversary were to choose $s$, they could make a target transaction public once they have established a substantial lead (this might take an exceedingly long time, but it would eventually occur with a probability of 1). By doing so, a safety violation would be guaranteed. In this paper, we eliminate such adaptive attacks by assuming that the transaction becomes public within a fixed time window that is beyond the adversary’s control.
Operationally, we have the flexibility to select an arbitrary number of intervals $(s,s+2\Delta]$. If an honest node first encounters a transaction during $(s,s+\Delta]$ (who will also propagate this transaction), then the node can be sure that the transaction becomes public during the interval $(s,s+2\Delta]$. The node can be confident about the probabilistic safety guarantees for the transaction due to the theorems given in Sec. 4.
Following the preceding outline of the statistical analysis, we prove the main theorems in detail in the appendices. In Sec. 6, we present the numerical results derived from these theorems. These results demonstrate the practicality of our latency-security trade-off bounds and their utility as guidance of parameter selection for Nakamoto-style protocols.
6. Numerical Result
In this section, the system parameters are chosen to represent two different settings in most cases. The first setting is modeled after Bitcoin, where the total mining rate $h+a=1/600$ blocks/second. Bitcoin blocks’ propagation delays vary and have been decreasing over the years. As of October 2023, the 90th percentile of block propagation delay is about 4 seconds on average (bitcoinmoniter, ). We use $\Delta=10$ seconds as an upper bound. The second setting is modeled after Ethereum Classic (ETC), where $h+a=1/13$ blocks/second and $\Delta=2$ seconds. We also include results for Litecoin and Dogecoin parameters in one table.
6.1. Security Latency Trade-off
Figure 3. Bitcoin’s safety versus confirmation time.
Figure 4. ETC’s safety versus confirmation time.
In Fig. 3, we first show the achievable and unachievable probability of safety violation as a function of the confirmation time (in hours) under the Bitcoin setting. For three different adversarial mining fractions ($\beta=1/10$, $1/7$, and $1/4$), we plot the upper and lower bounds calculated using Theorems 8 and 9, respectively. Also included for comparison is an achievable trade-off from (li2021close, , Eqn. (5)).
In Fig. 4, we plot the latency-security trade-off of ETC for two different adversarial mining fractions ($\beta=1/7$ and $1/4$). The gap between the upper and lower bounds is smaller than the gap derived in (li2021close, ).
In Fig. 5, we show Bitcoin’s safety bounds as a function of confirmation depth $k$ calculated using Theorems 25, and 6, respectively. Also included are the achievable and unachievable bounds from (guo2022bitcoin, , Eqns. (7) and (8)) for comparison. The new upper bound (5) of Theorem 5 yields substantially tighter result. Similarly, ETC’s safety bounds as a function of confirmation depth $k$ is plotted in Fig. 6. For the ETC parameters, the achievable bounds in this paper significantly outperform the achievable bound of (guo2022bitcoin, , Eqn. (7)).
Figure 5. Bitcoin’s safety versus confirmation depth.
Figure 6. ETH’s safety versus confirmation depth.
$\Delta$ 00 s $2.5$ s $10$ s $40$ s $160$ s $640$ s
$\beta=10\%$ $7$ $7$ $7$ $10$ $21$ $52$
$\beta=20\%$ $14$ $14$ $14$ $26$ $35$ $117$
$\beta=30\%$ $34$ $35$ $35$ $63$ $91$ $633$
$\beta=40\%$ $150$ $154$ $161$ $277$ $552$infty
Table 2. The confirmation depth needed to guarantee $99.9\%$ safety for various block delays and adversarial fractions.
For a fixed $99.9\%$ safety guarantee, using the Bitcoin parameters, we compute the confirmation depth needed according to Theorem 5 for various block propagation delays and adversarial mining fractions. Table 2 shows that the confirmation depth rises slowly with the delay bound when the adversarial fraction is relatively low; it rises more quickly when the adversarial fraction is relatively high.
$\bar{F}(0)$ $\bar{F}(1)$ $\bar{F}(2)$ $\bar{F}(3)$ $\bar{F}(4)$
Bitcoin $0.006$ $0.0001$ $2\times 10^{-6}$ $4\times 10^{-8}$ $8\times 10^{-10}$
ETC $0.06$ $0.008$ $0.001$ $2\times 10^{-4}$ $3\times 10^{-5}$
Table 3. The complementary cdf, $\bar{F}(n)$.
Table 3 shows the probability of having more than a given number of balanced heights with 25% adversarial mining. For example, the probability of having more than four balanced heights is less than $10^{-9}$ for Bitcoin. Hence the bound in Lemma 3 is fairly tight for our purposes.
$\lambda^{-1}$ (s) $\Delta$ (s) $10^{-3}$ $10^{-6}$ $10^{-9}$
Bitcoin (guo2022bitcoin, ) $600$ 10 $\leq 25$ $\leq 51$ $\leq 78$
Bitcoin new $600$ 10 $\leq 23$ $\leq 47$ $\leq 71$
Litecoin (guo2022bitcoin, ) $150$ 10 $\leq 36$ $\leq 74$ $\leq 110$
Litecoin new $150$ 10 $\leq 32$ $\leq 63$ $\leq 99$
Dogecoin (guo2022bitcoin, ) $60$ 10 $\leq 85$ $\leq 171$ $\leq 258$
Dogecoin new $60$ 10 $\leq 52$ $\leq 127$ $\leq 211$
ETC (guo2022bitcoin, ) $13$ 2 $\leq 73$ $\leq 151$ $\leq 230$
ETC new $13$ 2 $\leq 47$ $\leq 114$ $\leq 193$
Table 4. Bounds of confirmation depth (latency) for different security levels with $25\%$ adversarial mining power.
Table 4 presents the depth needed to guarantee different security levels (probability of target transaction’s safety being violated) including $10^{-3}$, $10^{-6}$, $10^{-9}$, for Bitcoin, Litecoin, Dogecoin, and ETC, respectively. The table shows that the new upper bound due to Theorem 5 is often substantially smaller than the previous upper bound from (guo2022bitcoin, ).
6.2. Confirmation versus Fault Tolerance
In this subsection, we plot the confirmation depth or time that guarantees a desired safety level as a function of the fraction of adversarial mining, while holding other parameters fixed. This also makes it easier to see the maximum fault tolerance different bounding methods can apply to. We assume the ETC parameters.
Refer to captionRefer to caption
Figure 7. Confirmation latency that guarantees 99.9% safety for ETC assuming different fractions of adversarial mining. Top: Confirmation in depth. Bottom: Confirmation in time (minutes).
Fig. 7 shows new and previous bounds on the confirmation depth/time that guarantees 99.9% safety as a function of the adversarial fraction of mining power. The top graph uses confirmation by depth (Theorems 25, and 6), whereas the bottom graph uses confirmation by time (Theorems 8 and 9). We include two points reported in (gazi2022practical, ), where the numerical evaluation becomes impractical for much higher adversarial mining rates. Also included in the graph are achievable infinite-latency fault tolerances established in (garay2015bitcoin, )(pass2017analysis, )(li2021close, ) and the ultimate tolerance established in (dembo2020everything, )(gazi2020tight, ), and also in this paper. The new upper and lower bounds are significantly closer than all previous bounds. Even our new closed-form upper bound is quite competitive.
6.3. Throughput-Latency Trade-Off
In this subsection, we evaluate the best throughput we can guarantee for optimized parameters including block size, mining rate, and confirmation depth. We fix the safety guarantee to be $99.9\%$. We guarantee up to $\beta=40\%$ adversarial mining, which implies that $\lambda\Delta<5/6$ due to (19). We assume the block propagation delay and the block size satisfy the affine relationship (21). Using empirically determined delay bounds of 10 seconds for 1.6 MB Bitcoin blocks and 2 seconds for 200 KB Bitcoin Cash blocks, [^3] The 90th percentile of Bitcoin Cash propagation time is 1 second (fechner2022calibrating, ). we infer that $r=178.6$ KB/second and $\nu=0.9$ (seconds).
Figure 8. Achievable and unachievable throughput-latency trade-off.
For every given expected confirmation latency $d$, we can solve the optimization problem (23) with $q=0.001$ and given natural fork rates $\lambda\Delta=3/4$, $7/12$, $5/12$, and $1/4$, respectively. For each given $\lambda\Delta$, we determine the confirmation depth $k$ needed to satisfy (23b) according to Theorem 2, which then allows us to determine the throughput for every choice of the average delay in time according to (23). Fig. 8 plots the throughput-latency trade-off for several different values of $\lambda\Delta$. The upper envelop (in green dashed line) of all such curves represents the guaranteed throughput latency trade-off. Also plotted in is the unachievable bound of Theorem 6. In particular, the maximum throughput is approximately $81$ KB/second due to (22). We observe that the choice of $\lambda\Delta=7/12$ yields a large fraction of the maximum possible throughput for practical latencies of $30$ minutes or more.
6.4. Guidance to Selecting Parameters
Bitcoin’s block size ($1.6$ MB) and mining rate ($1$ every $10$ minutes) have proven to be safe choices, with a maximum throughput of about $2.8$ KB/s. The block size has long been an issue of debate. Using a larger block size while holding other parameters unchanged may increase the throughput (and perhaps lower transaction fees), but at the same time degrade the safety due to larger propagation delays and increased forking (and perhaps also tendency of mining centralization). As another important factor, increasing the mining rate increases natural forks, but also reduces the expected time to the same confirmation depth.
The theorems in this paper provide the most accurate guidance to date to selecting parameters for desired system-level performance. For example, if the block size is increased to $2$ MB, the propagation upper bound will be around $\Delta=12.5$ seconds according to (21). If the total mining rate is increased to $\lambda=1/30$ block/seconds, the throughput increases to over $40$ KB/s according to (23a). At $\lambda\Delta=5/12$, the ultimate fault tolerance is $44.44\%$ adversarial mining; in this case, at $30$% adversarial mining, the latency for achieving $10^{-6}$ safety is less than $3.5$ hours, while original Bitcoin takes more than $18$ hours. Thus the throughput, latency, and safety are simultaneously improved for a minor sacrifice in the fault tolerance. In fact, $44.44\%$ adversarial mining is practically not tolerable using the original Bitcoin parameters due to the excessively long delays of thousands of block confirmations.
7. Conclusion
In this paper, we have developed a concise mathematical model for blockchain systems based on the Nakamoto consensus. We have obtained a new set of closed-form latency-security bounds, which are the first such bounds to apply to all parameter values. These bounds are not only easily computed but also demonstrate superior numerical performance across a broad spectrum of parameter configurations. Our analysis not only provides an intuitive proof of the ultimate fault tolerance but also offers new insights for the practical application the Nakamoto consensus.
Existing safety guarantees for longest-chain-fork-choice protocols have primarily been established through “block counting” methods, which calculates, the probability that one species of blocks ismined more than another species during some periods longer than the confirmation latency. Such analyses essentially empower the adversary to use non-causal knowledge of future blocks (up until confirmation) and advance them to build a winning chain. In contrast, the production of a balanced height depends on blocks mined within an opportunity window of $\Delta$ seconds; future blocks cannot be used to create balancing candidates or balanced heights. This requirement eliminates certain non-causal attack strategies, resulting in improvements and the precise determination of fault tolerance.
In today’s Bitcoin and future open blockchain networks, miners play a crucial role in investing in infrastructure to mitigate block propagation delays. Additionally, constant monitoring of delays and vigilant of observation of potential network partitions are crucial aspects of their operations. During periods of outage or network disruption, an accurate estimation of delays enables miners and users to adapt their confirmation latency to stay safe. Subsequently, as delays normalize, forks are easily resolved, and nodes can revert to utilizing smaller confirmation latency specifically designed for standard conditions. The theorems presented in this paper serve as useful tools, empowering all stakeholders to make informed economic decisions regarding their operations.
Appendix A Proof of Theorem 1
Figure 9. Illustration of mining times of pacers and A-blocks after time $s$.
Under a confirmation rule of latency $t$, assume a target transaction is public by time $s$. Within this proof, we count block arrivals from time $s$, i.e., the first A-block is the first A-block mined after $s$, and the first pacer is the first pacer mined after $s$. Let $X_{0}$ denote the time from $s$ to the mining time of the first A-block. For $i=1,2,\dots$, let $X_{i}$ denote the inter-arrival time between the $i$-th A-block and the $(i+1)$-st A-block. Then, $X_{0},X_{1},\dots$ are independent and identically distributed (i.i.d.) exponential random variables with mean $\frac{1}{a}$. Let $Y_{0}$ denote the time from $s$ to the mining time of the first pacer. For $i=1,2,\dots$, let $Y_{i}$ denote the inter-arrival time between the $i$-th pacer and the $(i+1)$-st pacer. It is easy to see that $Y_{1},Y_{2},\dots$ are i.i.d. random variables with mean $\frac{1}{h}+\Delta$, and $Y_{0}$ has mean less than $\frac{1}{h}+\Delta$. An illustration of these random variables is given in Fig. 9.
Let $E_{j}$ denote the event that the $j$-th A-block arrives before the $j$-th pacer but after time $s+t$. The probability of safety violation is at least
(45) $P(E_{j})=P(\sum_{i=0}^{j}X_{i}<\sum_{i=0}^{j}Y_{i},\sum_{i=0}^{j}X_{i}>t)$
(46) $\geq 1-P(\sum_{i=0}^{j}(X_{i}-Y_{i})\geq 0)-P(\sum_{i=0}^{j}X_{i}\leq t)$
$\geq 1-P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)-P(X_{0}-Y_{0}\geq n)$
(47) $\quad-P(\sum_{i=0}^{j}X_{i}\leq t)$
where (47) holds for every integer $n$. For every $\epsilon>0$, there exists $n$ such that $P(X_{0}-Y_{0}\geq n)<\epsilon/3$. For sufficiently large $j$, we also have $P(\sum_{i=0}^{j}X_{i}\leq t)<\epsilon/3$. By (2), we have $E[X_{1}-Y_{1}]<0$. Thus, by the law of large numbers,
(48) $P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)<\frac{\epsilon}{3}.$
We conclude that $P(E_{j})\rightarrow 1$ as $j\rightarrow\infty$.
Appendix B Proof of Theorem 2
B.1. Critical Events
Figure 10. Illustration of all possible state transitions over time.
As a convention, we refer to the genesis block as the 0-th balancing candidate. We describe several types of events as follows: For every $i\in\{0,1,\cdots\}$, let
  • $C_{i}$ denotes the event that $i$ or more candidates are eventually mined;
  • $B_{i}$ denotes the event that the $i$-th candidate is mined and subsequently balanced;
  • $U_{i}$ denotes the event that the $i$-th candidate is mined but is not balanced by the end of its candidacy;
Evidently, $B_{i}\cup U_{i}=C_{i}$. Also, $C_{0},C_{1},\dots$ is decreasing in the sense that $C_{i+1}\subset C_{i}$. By definition, $C_{i}$ is prerequisite for $C_{i+1}$. E.g., “3 or more” means $\{3,4,...\}$ and is included in “2 or more”, which means $\{2,3,4,...\}$. For convenience, we define
(49) $A_{i}=C_{i}\cap\bar{C}_{i+1}$
which denotes the event that the $i$-th candidate is the last candidate ever mined.
As time progresses, the events are seen to occur in natural order as illustrated in Fig. 10. The process begins from $U_{0}$ as the genesis block, which is also the 0-th candidate, cannot be balanced. If a first candidate is ever mined, the process transitions to $C_{1}$, otherwise it terminates at $A_{0}$. In particular, for every $i=1,2,\dots$, if $C_{i}$ ever occurs, then either $B_{i}$ or $U_{i}$ occurs. If $B_{i}$ occurs, then with probability 1 the next candidate will be mined, i.e., $C_{i+1}$ must occur. If $U_{i}$ occurs instead, then there is some chance $C_{i+1}$ occurs as well as some chance that no more candidates are mined, so that the process terminates at $A_{i}$.
Conditioned on the $k$-th candidate is mined, whose candidacy is no longer than $\Delta$, the probability that it is also balanced is no higher than the probability that an honest block is mined over a period of $\Delta$. Hence
(50) $P(B_{k}|C_{k})\leq\epsilon$
where $\epsilon$ is given by (12).
If $U_{k}$ occurs, then at the end of the $k$-th candidacy, the candidate is not balanced, so the candidate becomes public and is strictly higher than all disagreeing blocks. As such, all honest mining power will extend chains that include the target transaction and are higher than the $k$-th candidate. This implies that the adversary must extend a shorter chain to catch up with the highest public honest block in order for the next candidate to be mined. In Appendix B.2, we show that the chance that no additional candidate will ever be mined is lower bounded as
(51) $P(A_{k}|U_{k})\geq\delta$
where $\delta$ is given by (13). The lower bound (51) is positive only when (1) holds.
B.2. Proof of (51)
Here we wish to lower bound $P(A_{k}|U_{k})$, which is the probability that no more candidate is ever mined after that the $k$-th candidate is not balanced by the end of its candidacy. We refer to this as the honest miners’ escape. There is no chance to escape if $a>h/(1+h\Delta)$. Otherwise, an escape occurs if (but not only if) the following events occur: 1) during the $k$-th candidacy, no A-block is mined, so that when the candidate becomes public, it is higher than all A-blocks; 2) from the end of the candidacy, say time $r$, the number of A-blocks mined during $(r,r+t]$ is no more than the number of jumpers mined during $(r,r+t-\Delta]$ for every $t\geq 0$, so that the adversary never produces a disagreement about the target transaction after $r$.
Since the duration of the candidacy is at most $\Delta$, the probability of the first event is evidently lower bounded by $e^{-a\Delta}$. Conditioned on the first event, once the $k$-th candidacy ends, the next H-block is a jumper, which arrives after an exponential time. The chance of an escape is minimized by delaying the propagation of all H-blocks maximally from $r$ onward. In this case, the jumper process $J_{r,r+t-\Delta}$ is statistically identically to the pacer process $P_{\Delta,t}$. Conditioned on the first event, the second event’s probability is thus lower bounded as
(52) $P\left(\cap_{t\geq 0}\{A_{r,r+t}\leq P_{\Delta,t}\}\right)=P\left(\cap_{t\geq 0}\{A_{0,t}\leq P_{0,t}\}\right)$
(53) $=P\left(\max_{t\geq 0}\{A_{0,t}-P_{0,t}\}=0\right)$
(54) $=P(M=0).$
Based on the preceding bounds and the pmf of $M$ given by (11), we have
(55) $P(A_{k}|U_{k})$ $\geq e^{-a\Delta}P(M=0)$
(56) $=\delta.$
B.3. Partially Markov Processes
Consider a sample space in which each outcome corresponds to a (sample) path on the directional graph depicted by Fig. 10 from $U_{0}$ to some $A$ vertex. If the path ends at $A_{k}$, the outcome can be represented by a binary string $00\,b_{1}\dots b_{k-1}$ 0, where $b_{i}=1$ if $B_{i}$ is on the path (in stage $i$). The corresponding number of balanced heights is thus $X=b_{1}+\dots+b_{k-1}$.
Naturally, we let each vertex in the graph denote the event which consists of all paths that go through the vertex. Under condition (1), with probability 1 some $A_{k}$ event terminates the sample path. Thus every adversarial strategy induces a pmf on the (discrete) sample space, which must satisfy (50) and (51). Under such a pmf, the transitions in the graph are not Markov in general. In particular, we cannot say the event $B_{i}$ is independent of $B_{i-1}$ conditioned on $C_{i}$.
Let $P^{0}(\cdot)$ stand for the special pmf under which all state transitions on the directed graph are Markov with the following transition probabilities: For every $k\in\{0,1,...\}$,
(57) $P^{0}(B_{k}|C_{k})$ $=1-P^{0}(U_{k}|C_{k})=\epsilon,$
(58) $P^{0}(A_{k}|U_{k})$ $=1-P^{0}(C_{k+1}|U_{k})=\delta,$
where $\epsilon$ and $\delta$ are given by (12) and (13), respectively. These transition probabilities satisfy the prior constraints (50) and (51) with equality under condition (1).
We next prove that every pmf $P$ which satisfies (50) and (51) must also satisfy
(59) $P(X\leq n)\geq P^{0}(X\leq n).$
Towards this end, we define a sequence of pmfs, $P^{1},P^{2},\dots$, as follows: Let $0\,b_{1}\dots b_{i}*$ denote the event consisting of all sample paths with the prefix $0\,b_{1}\dots b_{i}$. Let the probability of a sample path under $P^{m}$ be defined as
(60) $P^{m}(\{0\,b_{1}\dots b_{k-1}0\})=P(\{0\,b_{1}\dots b_{k-1}0\})$
for $k=1,\dots,m$ and as
[…]
for $k=m+1,m+2,\dots$. The outcome $00\,b_{1}\dots b_{k-1}$ 0 with $k>m$ implies that $C_{m}$ occurs. It is straightforward to verify that (60), (61) defines a pmf. Evidently, $P^{m}$ is partially Markov in the sense that it is consistent with $P$ up until stage $m-1$, after which $P^{m}$ becomes Markov with the same transition probabilities as $P^{0}$. In particular, conditioned on $C_{m}$, the balanced stages before stage $m$ and the balanced stages afterward are independent.
For every pair of natural numbers $k$ and $j$, we define the following pair of random variables as functions of the sample path $00\,b_{1}\dots b_{k-1}$ 0:
(62) $L_{j}$ $=b_{1}+\dots+b_{\min(k,j)-1}$
(63) $R_{j}$ $=b_{j}+\dots+b_{k-1},$
with the convention that $L_{0}=0$. Evidently, with $01_{B_{0}}=$ 0
(64) $X=L_{m}+1_{B_{m}}+R_{m+1}$
holds for every natural number $m$.
In Appendices B.4 and B.5, we show that for every $m=0,1,\dots$ and every $n=0,1,\dots$,
(65) $P^{m+1}(X\leq n)\geq P^{m}(X\leq n)$
and
(66) $\lim_{m\to\infty}P^{m}(X\leq n)=P(X\leq n)$
hold. As the limit of an increasing sequence, $P(X\geq n)$ must satisfy (59).
For $n=0,1,\dots$, let us denote
(67) $f_{n}$ $=P^{0}(X\geq n|C_{1})$
(68) $=P^{0}(R_{1}\geq n|C_{1}).$
Then we have the following recursion
(69) $f_{n}=$ $P^{0}(R_{1}\geq n,U_{1}|C_{1})+P^{0}(R_{1}\geq n,B_{1}|C_{1})$
$=$ $P^{0}(R_{2}\geq n|C_{2})P^{0}(C_{2}|U_{1})P^{0}(U_{1}|C_{1})$
(70) $+P^{0}(R_{2}\geq n-1|B_{1})P^{0}(B_{1}|C_{1})$
(71) $=$ $(1-\delta)(1-\epsilon)f_{n}+\epsilon f_{n-1}.$
Solving the recursion with the fact that $f_{0}=1$ yields
(72) $f_{n}=\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}.$
It is easy to verify that
(73) $P^{0}(X\leq n)$ $=1-P^{0}(R_{1}\geq n+1|C_{1})P(C_{1})$
(74) $=F(n)$
where $F(\cdot)$ is defined in (15).
With confirmation by depth, there can be at most $k$ balanced heights before violation. In the remainder of this section, we are concerned with the number of balanced candidates up to the $k$-th candidacy, denoted as $X^{(k)}$. Conditioning on whether $C_{k+1}$ is visited, we can lower bound the cdf of $X^{(k)}$:
$P(X^{(k)}\leq n)$
(75) $\geq P^{0}(X^{(k)}\leq n)$
(76) $=P^{0}(X^{(k)}\leq n,X\geq n+1)+P^{0}(X\leq n)$
(77) $=P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})+P^{0}(X\leq n)$
where (75) can be easily seen from (65), and (77) holds because $X^{(k)}\leq n$ and $X\geq n+1$ implies the occurrence of $C_{k+1}$. Since we already know $P^{0}(X\leq n)$ by (74), what left is to calculate $P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$. By law of total probability, we have
(78) $\begin{split}&P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})\\ &=\sum_{i=0}^{n}P^{0}(X^{(k)}=i,C_{k+1})P^{0}(X-X^{(k)}\geq n+1-i|C_{k+1}).\end{split}$
Due to Markov property on $P^{0}$, $X-X^{(k)}\geq n+1-i$ conditioning on $C_{k+1}$ is identically distributed as $X\geq n+1-i$ conditioning on $C_{1}$. $P^{0}(X^{(k)}=i,C_{k+1})$ equals the sum of probabilities of all possible paths from $U_{0}$ to $C_{k+1}$ which traverse $i$ of the B states. For every positive integer $m$, conditioned on $C_{m}$, the probability of going through $B_{m}$ to arrive at $C_{m+1}$ is $\epsilon$, while the probability of not going through $B_{m}$ to arrive at $C_{m+1}$ is $(1-\epsilon)(1-\delta)$. Thus, the number of B states visited between $C_{1}$ and $C_{k+1}$ is a binomial random variable. We have
$P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$
(79) $=\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}((1-\epsilon)(1-\delta))^{k-i}(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-i+1}$
(80) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}(\frac{\epsilon(1-\epsilon)(1-\delta)}{\epsilon+\delta-\epsilon\delta})$
$=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{k}$
(81) $\qquad\sum_{i=0}^{n}\binom{k}{i}(\epsilon+\delta-\epsilon\delta)^{i}(1-\epsilon)(1-\delta)^{k-i}$
(82) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n+1}F_{b}(n;k,(\epsilon+\delta-\epsilon\delta)),$
where $F_{b}(\cdot;k,p)$ denotes the cdf of a binomial distribution with $k$ independent trials and success probability $p$. Plugging (82) back into (78) completes the proof of (14) and also Theorem 2.
B.4. Proof of (65)
We use $C_{m,i}$ to denote the intersection of events $C_{m}$ and $L_{m}=i$. Then
$P^{m+1}($ $R_{m+1}<c\,|\,C_{m+1,i},B_{m})$
(83) $=P^{0}(R_{m+1}<c\,|\,C_{m+1})$
(84) $=P^{m+1}(R_{m+1}<c\,|\,C_{m+1,i},U_{m}).$
Also, partially Markov process with measure $P^{m}(\cdot)$ becomes memoryless once event $C_{m}$ occurs:
(85) $P^{m}(A_{m}|C_{m,i})$ $=P^{0}(A_{m}|C_{m,i})$
(86) $P^{m}(B_{m}|C_{m,i})$ $=P^{0}(B_{m}|C_{m,i})$
(87) $P^{m}(C_{m,i})$ $=P(C_{m,i})$
(88) $=P^{m+1}(C_{m,i}).$
Consider the right-hand side of (65):
(89) $P^{m}(X\leq a)=P^{m}(X\leq a,C_{m}^{c})+P^{m}(X\leq a,C_{m})$
(90) $=P^{m}(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P^{m}(C_{m,i})P^{m}(X\leq a|C_{m,i})$
(91) $=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m}(X\leq a|C_{m,i})$
where (91) is due to the fact that $P^{m}$ is identical to $P$ up until stage $m-1$ as in (87). Similarly, by (88)
(92) $\begin{split}&P^{m+1}(X\leq a)\\ &=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m+1}(X\leq a|C_{m,i}).\end{split}$
In order to establish (65), it suffices to show
(93) $P^{m+1}(X\leq a|C_{m,i})\geq P^{m}(X\leq a|C_{m,i})$
holds for all $i\in\{0,\dots,a\}$.
Since $C_{m}=B_{m}\cup(U_{m},C_{m+1})\cup A_{m},$, we have
$P^{m+1}(X\leq a|C_{m,i})$
$=$ $P^{m+1}(X\leq a,B_{m}|C_{m,i})+P^{m+1}(X\leq a,A_{m}|C_{m,i})$
(94) $+P^{m+1}(X\leq a,U_{m},C_{m+1}|C_{m,i})$
$=$ $P^{m+1}(R_{m}\leq a-i-1,B_{m}|C_{m,i})+P^{m+1}(A_{m}|C_{m,i})$
(95) $+P^{m+1}(R_{m}\leq a-i,U_{m},C_{m+1}|C_{m,i})$
$=$ $P^{m+1}(R_{m}\leq a-i-1|C_{m+1})P^{m+1}(B_{m},C_{m+1}|C_{m,i})$
$\quad+P^{m+1}(R_{m}\leq a-i|C_{m+1})P^{m+1}(U_{m},C_{m+1}|C_{m,i})$
(96) $\quad+P^{m+1}(A_{m}|C_{m,i})$
$=$ $P^{m+1}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{m+1}(U_{m},C_{m+1}|C_{m,i})x_{2}$
(97) $\quad+P^{m+1}(A_{m}|C_{m,i})$
where
(98) $x_{1}$ $=P^{0}(R_{m}\leq a-i-1|C_{m+1})$
(99) $x_{2}$ $=P^{0}(R_{m}\leq a-i|C_{m+1}),$
and (95) is due to (64),(97) is due to (83). Similarly, with (85) and (86), we have
(100) $\begin{split}P^{m}&(X\leq a|C_{m,i})=P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}\\ &+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i}).\end{split}$
Due to (50) and (51),
(101) $P^{0}(A_{m}|C_{m,i})$ $\leq P^{m+1}(A_{m}|C_{m,i})$
(102) $P^{0}(B_{m},C_{m+1}|C_{m,i})$ $\geq P^{m+1}(B_{m},C_{m+1}|C_{m,i}).$
Now we introduce the following claim to assist the proof, which enables us to compare differently weighted sums of three numbers.
Lemma 1.
If $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$, $\lambda_{1}^{\prime}+\lambda_{2}^{\prime}+\lambda_{3}^{\prime}=1$ and $\lambda_{1}\geq\lambda_{1}^{\prime}$, $\lambda_{3}\leq\lambda_{3}^{\prime}$, $\lambda_{i}\text{ and }\lambda_{i}^{\prime}\geq 0$ for $i=1,2,3$. Then all real numbers $x_{1}\leq x_{2}\leq 1$ satisfy
(103) $\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}\leq\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
Proof.
The proof is straightforward:
$\lambda_{1}x_{1}$ $+\lambda_{2}x_{2}+\lambda_{3}$
(104) $=x_{2}+\lambda_{1}(x_{1}-x_{2})+\lambda_{3}(1-x_{2})$
(105) $\leq x_{2}+\lambda_{1}^{\prime}(x_{1}-x_{2})+\lambda_{3}^{\prime}(1-x_{2})$
(106) $=\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
Lemma 1 implies
$\quad P^{m+1}(X\leq a|C_{m,i})$
$\geq P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+$
(107) $\quad P^{0}(A_{m}|C_{m,i})$
(108) $=P^{m}(X\leq a|C_{m,i}),$
which completes the proof of (65).
B.5. Proof of (66)
By the definition of limit, proving (66) equals to prove $\forall\lambda\geq 0,a\in\{0,1,\dots\},\exists N_{0},s.t.\forall N>N_{0}$, $|P^{N}(X\leq a)-P(X\leq a)|<\lambda$.
Conditioning on the occurrence of event $C_{N}$,
$P^{N}(X\leq a)$ $=P^{N}(C_{N}^{c})P^{N}(X\leq a|C_{N}^{c})$
(109) $\quad+P^{N}(C_{N})P^{N}(X\leq a|C_{N}),$
(110) $P(X\leq a)$ $=P(C_{N}^{c})P(X\leq a|C_{N}^{c})+P(C_{N})P(X\leq a|C_{N}).$
By definition of (60) and (61), $P(X\leq a|C_{N}^{c})=P^{N}(X\leq a|C_{N}^{c})$ and $P(C_{N}^{c})=P^{N}(C_{N}^{c})$. Let $p_{1}=P^{N}(X\leq a|C_{N})$ and $p_{2}=P(X\leq a|C_{N}))$. We have,
$|P^{N}$ $(X\leq a)-P(X\leq a)|$
(111) $=P(C_{N})|p_{1}-p_{2}|$
(112) $\leq(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1}|p_{1}-p_{2}|$
(113) $<\lambda,$
where (112) is due to the following
(114) $P(C_{N})\leq P^{N-1}(C_{N})\leq P^{N-2}(C_{N})\leq\cdots$ $\leq P^{0}(C_{N}),$
(114) can be derived by definition of (60) and (61) as
(115) $P^{i}(C_{N})$ $=P^{0}(C_{N}|C_{i+1})P^{i}(C_{i+1}|C_{i})P^{i}(C_{i})$
(116) $\leq P^{0}(C_{N}|C_{i+1})P^{0}(C_{i+1}|C_{i})P^{i-1}(C_{i}).$
Then, by the memoryless property of the Markov process with measure $P^{0}(\cdot)$we have
(117) $P^{0}(C_{N})$ $=P^{0}(C_{1}|U_{0})\prod_{i=1}^{N-1}P^{0}(C_{i+1}|C_{i})$
(118) $=(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1},$
where for $i=\{1,2,\cdots,N-1\}$
(119) $P^{0}(C_{i+1}|C_{i})$ $=1-P^{0}(A_{i}|C_{i})$
(120) $=1-\delta(1-\epsilon).$
Thus $\lim\limits_{N\rightarrow\infty}P^{N}(X\leq a)=P(X\leq a)$. Hence (66) holds.
Appendix C Proof of Lemma 3
For every integer $n$, we have
(121) $P(X+Y\geq 0)$ $\leq P(X>-n\text{ or }Y>n-1)$
(122) $\leq P(X>-n)+P(Y>n-1)).$
Taking the infimum over $n$ on both sides yields the upper bound in Lemma 3.
We prove the other direction in the following. Denote $\phi(n)=\bar{F}_{X}(-n)+\bar{F}_{Y}(n-1)$, and we have $0\leq\phi\leq 2$. Since $\phi(n)\to 1$ as $|n|\to\infty$, $\inf_{n}\phi(n)\leq 1$. If $\inf_{n}\phi(n)=1$, the bound is trivial. Hence in the rest of the proof we may assume that $\phi(a)=\inf_{n}\phi(n)<1$ for some $a\in\mathbb{Z}$.
Then for all $k\in\mathbb{Z}$, we have
(123) $\bar{F}_{X}(-a+k)+\bar{F}_{Y}(a-k-1)\geq\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$
Or equivalently, for $k\in\mathbb{N}$,
(124) $P(-a<X\leq-a+k)\leq P(a-k-1<Y\leq a-1),$
(125) $P(a-1<Y\leq a+k-1)\leq P(-a-k<X\leq-a).$
To attain the bound, we must construct $X_{0}$ and $Y_{0}$ such that $X_{0}+Y_{0}\geq 0$ whenever $X_{0}>-a$ or $Y_{0}>a-1$, and we need to ensure the two cases are disjoint.
Set a random variable $U$, which has a uniform distribution over $[0,1]$. Given $F$, define
(126) $F^{-1}(x)=\sup\{a\in\mathbb{R}|F(a)<x\}.$
We define $X_{0}$ and $Y_{0}$ to be the following:
(127) $X_{0}=F_{X}^{-1}(U+F_{X}(-a)-\lfloor U+F_{X}(-a)\rfloor)$
(128) $Y_{0}=F_{Y}^{-1}(F_{Y}(a-1)-U-\lfloor F_{Y}(a-1)-U\rfloor)$
$X_{0}$ and $Y_{0}$ have cdf $F_{X}$ and $F_{Y}$ since they are rearrangements of $F^{-1}_{X}$ and $F^{-1}_{Y}$. Now, we show that $X_{0}(u)+Y_{0}(u)\geq 0$ for $u\leq 1-F_{X}(-a)$ $a.e.$ The same holds for $u>F_{Y}(a-1)$, and we omit this similar proof here.
(129) $F^{-1}_{X}(x)=\sup\{a\in\mathbb{R}|F_{X}(a)<x\}.$
implies that
(130) $F_{x}\circ F_{x}^{-1}(x)>x\quad a.e.,$
since $X$ is integer-valued. Consider
(131) $S_{b}=\left\{u\leq 1-F_{X}(-a)|X_{0}(u)\leq b,Y_{0}(u)\leq-b-1\right\}$
And,
(132) $F_{X}(b)$ $\geq F_{X}\circ X_{0}(u)$
(133) $=F_{X}\circ F^{-1}_{X}(u+F_{X}(-a))$
(134) $>u+F_{X}(-a)\quad a.e.,$
Accordingly,
(135) $F_{Y}(-b-1)$ $\geq F_{Y}\circ Y_{0}(u)$
(136) $=F_{Y}\circ F^{-1}_{Y}(F_{Y}(a-1)-u)$
(137) $>F_{Y}(a-1)-u\quad a.e.$
Adding up (134) and (137) yields
(138) $F_{X}(b)+F_{Y}(-b-1)>F_{X}(-a)+F_{Y}(a-1)\quad a.e.$
However, $\phi(a)$ is maximal, leading to $P(S_{b})=0$
(139) $P(u\leq 1-F_{X}(-a)|X_{0}(u)+Y_{0}(u)<0)=0$
Finally,
$P(X_{0}+Y_{0})$ $\geq 1-F_{X}(-a)+1-F_{Y}(a-1)$
(140) $\quad=\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$
Thus $X_{0}$ and $Y_{0}$ attain the optimal bound, which concludes the lemma.
Appendix D Proof of Theorem 2
This appendix finished the proof of Theorem 2 in Sec. 5.3. We utilize Theorem 2, proved in Appendix B, to derive a closed-form upper bound on the probability of a safety violation for confirmation by depth $k$. We produce two upper bounds, both of which decay exponentially with $k$. The first bound has a weaker exponent, but it is easy to verify its decay within the fault tolerance (1). The second bound has a steeper exponent whose expression is more complicated.
D.1. Upper Bounding $Y$
Recall that $Y^{(k)}$ is defined in (37). Due to the renewal nature of the pacer process, we have $P_{\tau+\Delta,\tau+\Delta+c}$ stochastically dominating $P^{\prime}_{0,c}$, where $P^{\prime}$ is an independent pacer process of the same statistics as $P$. Hence
$P(Y^{(k)}>-n)$
(141) $\leq P\!\big{(}L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}$
(142) $\leq P\!\big{(}L_{s}-k+A_{s,\tau+2\Delta}+\sup_{c\geq 0}\{A_{\tau+2\Delta,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}.$
The lead $L_{s}$ depends only on blocks mined by $s$. The supreme term in (142) depends only on the mining processes after $\tau+2\Delta>s$. Hence the supreme term in (142) is independent of $L_{s}-k+A_{s,\tau+2\Delta}$.
Let us define
(143) $M=\sup_{t\geq 0}\{A^{\prime}_{0,t}-P^{\prime}_{0,t}\}$
where $A^{\prime}$ is a Poisson process identically distributed as $A$. By Lemma 4, the pmf of $M$ is given by (11). Evidently, the supreme term in (142) is identically distributed as $M$. We can thus rewrite (142) as
(144) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,\tau+2\Delta}+M>-n\right).$
Let the first H-block mined after $s+3\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let us also define $s+T+2\Delta$ to be the first time $k$ new pacers are mined. Since the height of the $k$-th new pacer is at least $h_{b}+k$, we have $s+T+2\Delta\geq\tau$ by $\tau$’s definition. We then relax (144) to
(145) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,s+T+4\Delta}+M>-n\right).$
While $\tau$ is generally dependent on $L_{s}$, we have $T$ independent of $L_{s}$ due to the memoryless nature of mining.
[…]
(146) $P(Y^{(k)}>-n)$ $\leq P\left(L+A_{0,2\Delta}-k+A_{2\Delta,T+6\Delta}+M>-n\right)$
(147) $=P\left(L+M+A_{0,T+6\Delta}-k+n>0\right).$
In (147), $L$, $M$, and $A_{0,T+4\Delta}$ are mutually independent. We shall derive their MGFs shortly and use them to upper bound the probability on the right hand side of (147). We take two separate approaches: In Appendix F, we let $n=ck$ with a specific choice of $c>0$ and establish Corollary 3 by showing that as long as (1) holds, (39) and (147) both vanish as $k\to\infty$.
D.2. Chernoff Bound
In this appendix, obtain an explicit exponential bound and thus establish Theorem 2. (This analysis can be thought of as a refinement of the analysis in Appendix F.) One step is to optimize $\nu$ introduced for the Chernoff bound. A second step is to select $n$ carefully to balance the two exponents corresponding to $P(X^{(k)}\geq n)$ and $P(Y^{(k)}>-n)$, respectively.
First, by Theorem 2,
(148) $P(X^{(k)}\geq n)$ $\leq P(X\geq n)$
(149) $\leq(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}$
where $\epsilon$ and $\delta$ are defined in (12) and (13), respectively. Note that this even holds for non-integer $n$.
We can use Chernoff inequality again to upper bound $P(Y^{(k)}>-n)$. Recall that $\bar{a}$ and $\bar{h}$ are defined as $a\Delta$ and $h\Delta$, respectively. Using the same computation in Appendix F, we have
(150) $P(Y^{(k)}>-n)$ $\leq P(M+L+A_{0,T+6\Delta}+n-k>0)$
(151) $\leq e^{(e^{\nu}-1)6\bar{a}}\left(q(\nu)\right)^{k}e^{\nu n}E[e^{\nu M}]E[e^{\nu L}]$
where
(152) $q(\nu)=\frac{e^{(e^{\nu}-1)\bar{a}-\nu}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}.$
We select $\nu$ to minimize $q(\nu)$, which is not difficult by setting the derivative $q^{\prime}(\nu)=0$. The unique valid solution $\nu^{*}$ satisfies
(153) $e^{\nu^{*}}=1+\gamma(\bar{a},\bar{h})/\bar{a}$
where
(154) $\gamma(\alpha,\eta)=\frac{1}{2}\left(2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}\right).$
Letting $n=\zeta k$ with $\zeta$ to be decided later, we have
(155) $P(X^{(k)}\geq\zeta k)\leq(1-\delta)e^{\zeta k\,\ln\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)}.$
Using lemma 4, we evaluate the MGFs of $M$ and $L$ as
(156) $E[e^{\nu^{*}L}]$ $=E[e^{\nu^{*}M}]$
(157) $=\mathcal{E}\left(1+\frac{\gamma(\bar{a},\bar{h})}{\bar{a}}\right).$
Letting $n=\zeta k$ and plugging (153) back into (151) and also evaluating the MGFs of $M$ and $L$ , we obtain
(158) $P(Y^{(k)}>-n)\leq f(\bar{a},\bar{h})e^{(\zeta\nu^{*}-\xi(\bar{a},\bar{h}))k}$
where
(159) $f(\alpha,\eta)=\frac{e^{6\gamma}\gamma^{2}(\alpha-\eta+\alpha\eta)^{2}}{((\alpha+\gamma)(\gamma-\eta)+\alpha\eta e^{\gamma})^{2}}$
and
(160) $\xi(\alpha,\eta)$ $=-\gamma-\ln\frac{\eta\alpha}{(\alpha+\gamma)(\eta-\gamma)}$
(161) $=\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma.$
To strike a balance between the exponents of $P(Y^{(k)}>-\zeta k)$ and $P(X^{(k)}\geq\zeta k)$, we let
(162) $\zeta\nu^{*}-\xi(\bar{a},\bar{h})=\zeta\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}).$
It gives us
(163) $\zeta=\frac{\xi(\bar{a},\bar{h})}{\ln(1+\frac{\gamma}{\bar{a}})-\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})},$
such that
(164) $P(Y^{(k)}>-\zeta k)\leq f(\bar{a},\bar{h})(e^{-c(\bar{a},\bar{h})})^{k},$
and
(165) $P(X^{(k)}\geq\zeta k)\leq(1-\delta)(e^{-c(\bar{a},\bar{h})})^{k},$
with
(166) $c(\bar{a},\bar{h})$ $=\xi(\bar{a},\bar{h})-\nu^{*}\zeta$
(167) $=\xi(\bar{a},\bar{h})(1-\ln(1+\frac{\gamma}{\bar{a}})(\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}))^{-1})^{-1},$
which can be seen from the exponential parameter of (158).
Denote
(168) $b(\bar{a},\bar{h})$ $=f(\bar{a},\bar{h})+1-\delta$
(169) $=\frac{e^{8\gamma}\gamma^{2}(\bar{a}-\bar{h}+\bar{a}\bar{h})^{2}}{((\bar{a}+\gamma)(\gamma-\bar{h})+\bar{a}\bar{h}e^{\gamma})^{2}}+1-\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)e^{-\bar{a}}.$
We have
(170) $P(X^{(k)}\geq\zeta k)+P(Y^{(k)}>-\zeta k)\leq b(\bar{a},\bar{h})e^{-c(\bar{a},\bar{h})k},$
which completes the proof of Theorem 2.
In Appendices G and H, we evaluate the error events in better accuracy than the Chernoff bounds allow us to develop a pair of upper and lower bounds for the probability of safety violation. The results are quite easy to evaluate numerically.
Appendix E Bounding the Lead
The lead increases by at most one upon the mining of an A-block and decreases by at least one upon the mining of a jumper unless the lead is already zero, in which case it stays at 0 with the mining of a jumper. Note that $L_{0}=0$. On every sample path, there exists $u\in[0,s]$ with $L_{u}=0$ such that
(171) $L_{s}-L_{u}$ $\leq A_{u,s}-J_{u,s}.$
If $L_{s}=0$, $u=s$ satisfies (171). If $L_{s}>0$, $u$ is in the last interval of zero lead before $s$. In general, $L_{u}=0$ and (171) imply that
(172) $L_{s}$ $\leq\sup_{q\in[0,s]}\{A_{q,s}-J_{q,s}\}$
(173) $\leq\sup_{q\in[0,s]}\{A_{q,s}-P_{q+\Delta,s}\}$
(174) $\leq\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{s-r,s}\}+A_{s-2\Delta,s}.$
where we have invoked Lemma 4 to obtain (173) and also used a change of variable $r=s-\Delta-q$ to obtain (174).
Let us introduce a reverse pacer process, $P^{s}_{0,r}$, which denotes the number of pacers mined during $(s-r,s]$, i.e.,
(175) $P^{s}_{0,r}=P_{s-r,s}.$
Due to the nature of the pacer process, one can show that $P^{s}_{0,r}$ stochastically dominates $P_{0,r-\Delta}$ as a process. Since $A$ is a memoryless Poisson process, $A_{s-\Delta-r,s-2\Delta}$ is statistically the same process as $A_{0,r-\Delta}$. Recall also $A^{\prime}$ is statistically identical to $A$. We can use (174) to show that $L_{s}$ is stochastically dominated in the sense that for every real number $x$,
$P(L_{s}\geq x)$
(176) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{0,r-\Delta}\}+A_{s-2\Delta,s}\geq x)$
(177) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{0,r-\Delta}-P_{0,r-\Delta}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
(178) $\leq P(\sup_{q\geq 0}\{A_{0,q}-P_{0,q}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
(179) $=P(L+A^{\prime}_{t,t+2\Delta}\geq x).$
Appendix F Proof of Corollary 3
This proof follows Appendix D.1. It suffices to show that for every $\epsilon>0$, there exists $c>0$ such that
(180) $P(X^{(k)}\geq ck)$ $\leq\epsilon/2$
(181) $P(Y^{(k)}>-ck)$ $\leq\epsilon/2$
for sufficiently large $k$. By Theorem 2, (180) holds for large enough $k$ for every $c>0$. The remainder of this appendix is devoted to the proof of (181).
By (147) and the Chernoff bound, we have for $\nu>0$,
(182) $P(Y^{(k)}>-ck)$ $\leq P(L+M+A_{0,T+6\Delta}>(1-c)k)$
(183) $\leq E[e^{\nu(M+L+A_{0,T+6\Delta}+(c-1)k)}]$
(184) $=E[e^{\nu(A_{0,T+6\Delta}+(c-1)k)}]E[e^{\nu M}]E[e^{\nu L}]$
where $E[e^{\nu L}]=E[e^{\nu M}]=\mathcal{E}(e^{\nu})$ is finite for every $\nu>0$ by Lemma 4, where $\mathcal{E}$ is defined as (10).
Only the first term on the right hand side of (184) depends on $k$. Given $t\geq 0$, $A_{0,t}$ is a Poisson random variable with mean $at$, whose MGF is
(185) $E[e^{\nu A_{0,t}}]=e^{(e^{\nu}-1)at}.$
[…]
(186) $E[e^{\eta Z}]=\frac{1}{1-\eta/h}.$
Using (185) and (186), we can write
$E[$ $e^{\nu(A_{0,T+4\Delta}+(c-1)k)}]$
(187) $=E[E[e^{\nu(A_{0,T+6\Delta})}|T]]\cdot e^{\nu(c-1)k}$
(188) $=E[e^{(e^{\nu}-1)a(T+6\Delta)}]\cdot e^{\nu(c-1)k}$
(189) $=E[e^{(e^{\nu}-1)aT}]\cdot e^{(e^{\nu}-1)6a\Delta}\cdot e^{\nu(c-1)k}$
(190) $=\left(E[e^{(e^{\nu}-1)aZ}]\right)^{k}\cdot e^{(e^{\nu}-1)(6+k)a\Delta}\cdot e^{\nu(c-1)k}$
(191) $=e^{(e^{\nu}-1)6a\Delta}\left(\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}\right)^{k}$
(192) $=e^{(e^{\nu}-1)6a\Delta}(g(\nu,c))^{k}$
where
(193) $g(\nu,c)$ $=\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}$
(194) $=\frac{e^{(e^{\nu}-1)\bar{a}}e^{\nu(c-1)}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}$
where $\bar{a}=a\Delta$ and $\bar{h}=h\Delta$.
In the remainder of this appendix, we will specify $\nu$ and $c$ as functions of $(\bar{a},\bar{h})$ which satisfy
(195) $g(\nu,c)<1,$
so that the right hand side of (192) and thus also of (184) vanishes as $k\to\infty$. Throughout this appendix, we denote
(196) $c=\frac{1}{4}\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)$
(197) $\theta$ $=\bar{a}(e^{\nu}-1).$
We have $c>0$ for mining rate and propagation delay satisfy the fault tolerance as (1). It is also quite obvious that $12c<$ 1. Now we choose
(198) $\theta=\frac{1}{2}\min\{1,\frac{\bar{a}}{2},\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}\}$
which is positive and yields the corresponding $\nu=\ln(1+\frac{\theta}{a})>0$. We rewrite (193) into
(199) $g(\nu,c)=\frac{e^{\theta}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
By (198), we have $\theta<1$ and $\frac{\theta}{\bar{a}}<\frac{1}{2}$, so that
(200) $g(\nu,c)$ $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
(201) $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+(1-2c)\frac{\theta}{\bar{a}})}$
(202) $=\frac{1+\theta+\theta^{2}}{1+\theta+\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}}$
where we have used the fact that $(1-2c)/\bar{a}-1/\bar{h}=1+2c/\bar{a}$. Since $\theta$ is strictly less than $\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}$, we have
(203) $\theta^{2}<\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}.$
Together, (202) and (203) guarantee (195). With the chosen $\nu$ and $c$, for sufficiently large $k$, (181) holds. The proof of Corollary 3 is thus complete.
Appendix G Proof of Theorem 5
To derive a tighter upper bound in the case of confirmation by depth, we evaluate $P(Y^{(k)}>-n)$ without resorting to the Chernoff bound. Starting from (147) in Appendix D, we use the law of total probability to write
$P($ $A_{0,T+6\Delta}\leq-n+k-(M+L)))$
$=$ $\sum_{i=0}^{k-n+1}P(L=i)(\sum_{j=0}^{k-n+1-i}P(M=j)$
(204) $\quad P(A_{0,T+6\Delta}\leq-n+k-(i+j)))$
$=$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
[…]
$=$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
[…]
Note the pdf of the $k$-st new pacer’s mining time $T$ , $f_{T}(t)$ can be driven by taking the derivative of its cdf , which is
(207) $F_{2}(t-k\Delta;k,h),$
where $F_{2}$ denotes the cdf of Erlang distribution. Now we only have to calculate the last term $P(A^{\prime}_{0,t+6\Delta}\leq-n+k-(i+j))$, which equals
(208) $F_{1}(-n+k-(i+j);a(t+6\Delta))).$
Finally, we have
$P(Y^{(k)}\leq-n)$
$\geq$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
[…]
$=$ $\sum_{i=0}^{k-n+1}e(i)\sum_{j=0}^{k-n+1-i}e(j)\int_{0}^{\infty}F_{1}(-n+k-(i+j);a(t+6\Delta))$
(210) $\cdot f_{2}(t-k\Delta;k,h)\,dt.$
Thus, according to (41), the probability of violation is upper bounded by
(211) $\inf_{n}\{2-P(Y^{(k)}\leq-n)-P(X^{(k)}\leq n-1)\},$
which equals (5). This completes the proof of Theorem 5.
Appendix H Proof of Theorem 6
Consider the same attack as the private-mining attack in Definition 14 except that we let the propagation delays of all H-blocks be zero before time $s-\Delta$ for simplicity. The adversary mines privately from $s-\Delta$ and delays all new H-blocks maximally. Let the first H-block mined after $s-\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on. Let $P^{\prime}_{r}$ denote the number of new pacers mined from $s-\Delta$ up until time $s+r-\Delta$. Evidently, $P^{\prime}_{r}$ is identically distributed as $P_{0,r+\Delta}$.
A violation occurs if there exists $d\geq s+\tau$ such that 1) the lead $L_{s-\Delta}$ plus the number of A-blocks $A_{s-\Delta,d}$ is no fewer than $P^{\prime}_{d-s+\Delta}$, the number of new pacers mined between $s-\Delta$ and $d-\Delta$ (any additional ones mined after that is not yet public); and 2) at least one pacer arrives by time $d$ to carry the target transaction, i.e., $P^{\prime}_{d-s+\Delta}\geq 1$. Hence the probability of violation is lower bounded by
(212) $P$ $\left(\cup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right),$
where $\tau$ is the time it takes for $k$ such new pacers to be mined after $s$. And the probability of (212) can be further bounded from below by
$P\left(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
(213) $\quad-P\left(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
The second term is easily determined as
(214) $P(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{\tau+\Delta}=0)$
(215) $=e^{-h(\tau+2\Delta)}.$
Since $L_{s-\Delta}$, $A_{s-\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we have the following similar in Appendix D,
$P(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(216) $=P(L_{s-\Delta}+\max_{d\geq s+\tau}\{A_{s-\Delta,d}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
(217) $=P(L_{s-\Delta}+\max_{d\geq+\tau}\{A^{\prime}_{0,d-s-\Delta}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
(218) $\geq P(L_{s-\Delta}+A^{\prime}_{0,\tau-\Delta}+\max_{d\geq\tau}\{A^{\prime}_{\tau-\Delta,d-\Delta}-P^{\prime}_{0,d-\tau}\}-P^{\prime}_{\Delta}-k\geq 0)$
(219) $=P(P^{\prime\prime}_{\Delta,2\Delta}\leq A^{\prime}_{0,\tau-\Delta}+M+L_{s-\Delta}-k),$
where (218) is due to that there are $k$ new pacers during $(s,\tau]$. $M$ in (219) is defined in (143). $P^{\prime}$ and $P$ are independent identically distributed. Thus, (219) can be written as
(220) $\begin{split}P&(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})\\ &\geq 1-P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-M-L_{s-\Delta}).\end{split}$
Now, we want to calculate $P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)$ Conditioning on $P^{\prime\prime}_{\Delta,2\Delta}$,
(221) $\begin{split}P(&A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)\\ &=P(A^{\prime}_{0,\tau-\Delta}\leq k-1-L_{s-\Delta}-M)f_{1}(0,h\Delta)\\ &\;\;+P(A_{0,\tau}\leq k-L_{s-\Delta}-M)(1-f_{1}(0,h\Delta)).\end{split}$
We have
$P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$
(222) $=$ $\sum_{i=0}^{k}\sum_{j=0}^{k-i}P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)P(M=j)P(L_{s-\Delta}=i),$
where $P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$ can be calculated as
[…]
where $\tau$ has cdf $F_{2}(t-k\Delta;k,h)$, and
(224) $P(A_{0,t-\Delta}\leq k-i-j)=F_{1}(k-i-j;a(t-\Delta)).$
Finally, the probability that the target transaction’s safety is violated is no large than (6). To ease computation, we can also relax the infinite integral to finite integral as
[…]
(225) $\quad+\bar{F}_{2}(T;k,h).$
Appendix I Proof of Theorem 8
There are ample similarities between confirmation by time and confirmation by depth. For example, the distribution of the adversarial lead $L_{s}$ remains the same. Also, we shall express the error event in terms of the race between the pacers and the A-blocks, where the latter is helped by the lead and the balanced heights. However, the starting time of the error events of interest is different than in the case of confirmation by depth.
I.1. An Enlarged Violation Event
Definition 20 (Unbalanced jumper).
We let $\hat{J}_{s,t}$ denote the number of jumpers mined during $(s,t-\Delta]$ which are unbalanced.
The sequence of jumpers includes all candidates, some of which may be balanced. Whether a jumper mined at time $s$ is balanced or unbalanced is known by time $s+\Delta$.
Lemma 1.
Suppose $r^{\prime}\geq r\geq s\geq 0$. If an $r$-credible chain and an $r^{\prime}$-credible chain disagree on a target transaction that is public at time $s$, then
(226) $L_{s}+A_{s,r^{\prime}}\geq\hat{J}_{s,r}$
Proof.
If $s\geq r-\Delta$, we have $\hat{J}_{s,r}=0$, so (226) holds trivially. We assume $s<r-\Delta$ in the remainder of this proof. Let block $b$ be the first H-block mined after time $s$. Since the target transaction is public at time $s$, Definition 10 implies that there must be disagreements on all heights of jumpers mined from $s$ to $r-\Delta$, i.e., on no fewer than $J_{s,r-\Delta}$ heights. Let us denote the highest honest block before time $s$ as height $\eta$. Every unbalanced jumper mined during $(s,r-\Delta]$ must meet either one of the following two conditions:
(1) All honest blocks on their height are in agreement, so all blocks on that height in a disagreeing branch must be adversarial;
(2) Some honest blocks on their height disagree, but at least one A-block is mined on that height by time $r^{\prime}$.
In either case, at least one A-block must be mined by $r^{\prime}$ on the height of every unbalanced jumper mined during $(s,r-\Delta]$. And the number of those unbalanced jumpers mined during $(s,r-\Delta]$ is $\hat{J}_{s,r}$.
The adversary needs to mine at least on $\hat{J}_{s,r}$ heights of unbalanced jumpers. Due to the pre-mining phase, at time $s$, the highest A-block has a height $\eta+L_{s}$, which is $L_{s}$ higher than the highest honest block, which means at most $L_{s}$ of A-blocks can be mined before time $s$ on the heights of unbalanced jumpers. The rest A-blocks can only be mined after time $s$, the number of which is $A_{s,r^{\prime}}$. Therefore, the disagreement about the target transaction implies the inequality (226). Hence the lemma is proved. ∎
Lemma 2.
Fix $\epsilon\in(0,\Delta)$, barring the following event
(227) $F_{s,t}^{\epsilon}=\bigcup_{d\geq t+s}\left\{L_{s}+A_{s,d}\geq J_{s,d-\epsilon-\Delta}-X\right\}$
all $r^{\prime}$-credible chains agree on the target transaction for all $r^{\prime}\geq t+s$.
Proof.
Let $r=t+s$. We first establish the result for $r^{\prime}\in[r,r+\epsilon]$ by contradiction. Then prove the lemma by induction. Fix arbitrary $r^{\prime}\in[r,r+\epsilon]$. If an $r^{\prime}$-credible chain and an $r$-credible chain disagree, then we have the following by Lemma 1:
(228) $L_{s}+A_{s,r^{\prime}}$ $\geq\hat{J}_{s,r}$
(229) $\geq\hat{J}_{s,r^{\prime}-\epsilon}$
(230) $\geq J_{s,r^{\prime}-\epsilon-\Delta}-X$
where (230) is due to that the number of balanced candidates mined during $(s,r^{\prime}-\epsilon]$ is smaller than $X$. If (230) occurs, $F_{s,t}^{\epsilon}$ also occurs. So this implies that if $F_{s,t}^{\epsilon}$ does not occur, every $r^{\prime}$-credible chain agrees with the $r$-credible chain, proving the lemma in this case.
Now suppose this lemma holds for $r^{\prime}\in[r,r+n\epsilon]$ for some positive integer $n$. We show the lemma also holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$ as the following: First, let $r^{\prime\prime}=r+n\epsilon$. A repetition of the contradiction case above with $r$ replaced by $r^{\prime\prime}$ implies that the lemma also holds for $r^{\prime}\in[r^{\prime\prime},r^{\prime\prime}+\epsilon]$. Therefore lemma holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$. The lemma is then established by induction on $n$.
By Lemma 4, we have
(231) $J_{s,d-\epsilon-\Delta}-X$ $\geq P_{s+\Delta,d-\epsilon-\Delta}-X.$
By (227) and (231), we have
(232) $F_{s,t}^{\epsilon}\subset\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\}.$
Hence
(233) $P(F_{s,t}^{\epsilon})$ $\leq P(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\})$
(234) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}+X\geq 0),$
where equality is by the definition of the union.
Let us define
(235) $Y_{t}=\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}.$
We apply Lemma 3 to obtain
(236) $P(F_{s,t}^{\epsilon})\leq\inf_{n\in\{1,2,...\}}\{P(X\geq n)+P(Y_{t}>-n)\}$
where it suffices to take the infimum over the natural number set only.
I.2. Probability Bound
We next upper bound $P(Y_{t}>-n)$ using properties of the adversarial mining process, the jumper process, as well as Kroese’s results on the difference of renewal processes. It should not be surprising that the bounds are essentially exponential in the confirmation time when evaluated numerically.
We define
(237) $T=\min\{t\in[s,s+\Delta]:P_{t}=P_{s+\Delta}\},$
which represents the last time a pacer is mined by $s+\Delta$. Evidently, $T$ is a renewal point, conditioned on which the subsequent pacer arrival times are independent of all mining times before $T$. In particular, $P_{T,T+r}$ is identically distributed as $P^{\prime}_{0,r}$ for all $r\geq 0$. Also, the lead $L_{s}$ and the pacer process $P_{T,d-\epsilon-\Delta}$ are independent conditioned on $T$. Moreover, $A_{s,d}$ is independent of $T$, $L_{s}$, and the pacer process. We evaluate $P(Y_{t}\leq-n)$ by conditioning on and then averaging over $T$:
(238) $P(Y_{t}\leq-n)=\mathbb{E}[P(Y_{t}\leq-n|T)]$
(239) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{T,d-\epsilon-\Delta}\}\leq-n|T)]$
(240) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-\epsilon-\Delta-T}\}\leq-n|T)].$
By (237), $d-\epsilon-\Delta-T\geq d-s-\epsilon-2\Delta$, so (240) implies
$P(Y_{t}\leq-n)$
(241) $\geq\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n|T)]$
(242) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n)$
where we drop conditioning on $T$ because all other random variables in (241) are independent of $T$. As a Poisson random variable, $A_{s,d}$ is identically distributed as $A^{\prime}_{0,d-s}$. Letting $q=d-s$, we rewrite (242) as
$P(Y_{t}\leq-n)$
(243) $\geq$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,q}-P^{\prime}_{0,q-\epsilon-2\Delta}\}\leq-n)$
(244) $=$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime}_{t-\epsilon-2\Delta,q-\epsilon-2\Delta}\}\leq-n).$
We further replace $A^{\prime}_{t,q}$ by $A^{\prime\prime}_{0,q-t}$ in (244). We also note that $P^{\prime}_{t-2\Delta-\epsilon,q-2\Delta-\epsilon}$ is statistically dominated by $P^{\prime\prime}_{0,q-t}$. So, we rewrite (244) as
$\quad P(Y_{t}\leq-n)$
(245) $\geq P\left(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
(246) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+\sup_{q\geq t}\{A^{\prime\prime}_{0,q-t}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
(247) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+M\leq-n\right)$
(248) $=1-P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq L_{s}+A_{0,t}^{\prime}+M+n-1)$
where
(249) $M=\sup_{r\geq 0}\{A^{\prime\prime}_{0,r}-P^{\prime\prime}_{0,r}\}.$
As the maximum difference between two renewal processes, $M$ has the pmf (11) by (Kroese_difference, , Theorem 3.4) (see also (li2021close, , Sec. 6)). given by (11).
Thus, by (236) and (248), we have
$P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(250) $+P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
Taking the limit of $\epsilon\to 0$, we use continuity to obtain
$\lim_{\epsilon\rightarrow 0}P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(251) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
The second term on the right-hand side of (251) is essentially the probability that the number of pacers mined during $(0,t-2\Delta]$ is fewer than the number of blocks mined by the adversary over a duration of $t$, plus the adversarial advantage after t, plus the adversary’s pre-mining gain, plus the number of balanced heights. The analysis thus far allows us to express this in terms of mutually independent variables $P^{\prime}_{0,t-2\Delta}$, $A^{\prime}_{0,t}$, $M$, and $L_{s}$. We analyze their individual distributions in the remainder of this subsection.
Let $L$ denote an independent and identically distributed random variable as $M$. In Appendix E, we show that $L_{s}$ is stochastically dominated by $L+A^{\prime}_{t,t+2\Delta}$. Hence (251) still holds with $L_{s}$ replaced by $L+A^{\prime}_{t,t+2\Delta}$ to yield
$\lim_{\epsilon\to 0}P(F_{s,t}^{\epsilon})\leq$ $\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(252) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)\}$
where $P^{\prime}$, $A^{\prime}$, $M$, and $L$ are independent.
We now invoke the preceding results to prove Theorem 8. Basically, we evaluate the upper bound of the probability of violation using the distributions (or their bounds) developed for $X$, $M$, and $L$. Recall the inter-pacer arrival times are $\Delta$ plus exponential variables. We have in general
(253) $P(P^{\prime}_{0,t}\leq i-1)$ $=P(i\Delta+X_{1}+\cdots+X_{i}>t)$
[…]
(254) $P(P^{\prime}_{0,t}\leq i-1)=1-F_{2}(t-i\Delta;i,h)$
where $F_{2}(\cdot;i,h)$ denotes the Erlang cdf with shape parameter $i$ and rate $h$.
By independence of those random variables, we have
$P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)$
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}P(A_{0,t+2\Delta}^{\prime}=i)P(M=j)P(L=k)$
(255) $\cdot P(P_{0,t-2\Delta}^{\prime}\leq i+j+k+n-1)$
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)$
(256) $\cdot(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h)),$
where $f_{1}(\cdot;\omega)$ denotes the pmf of a Poisson distribution with mean $\omega$. (The cdf of $P^{\prime}_{0,t}$ has been evaluated in (li2021close, , Sec. 6.1).) It is easy to see that (256) is equal to $G(n,t)$ defined in (25).
By (252), the safety of the said transaction confirmed using latency $t$ can be violated with a probability no higher than
(257) $\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+G(n,t)\},$
which is equal to (24).
Since it is numerically inconvenient to compute infinite sums, we further derive an upper bound using only finite sums, which is (slightly) looser than (256). For all positive integers $i_{0},j_{0},k_{0}$, it is easy to verify that
$\quad P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)$
$\leq 1-P(A_{0,t+2\Delta}^{\prime}\leq i_{0})P(M\leq j_{0})P(L\leq k_{0})$
$\quad+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n,A_{0,t+2\Delta}^{\prime}\leq i_{0},$
(258) $\quad M\leq j_{0},L\leq k_{0})$
(259) $\leq\overline{G}(n,t),$
where $\overline{G}(n,t)$ is calculated as (26).
Finally, we have that the safety of a given transaction confirmed using latency $t$ can be violated with a probability no higher than
$\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+\overline{G}(n,t)\}$
(260) $=\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+\overline{G}(n,t)\}.$
Appendix J Proof of Theorem 9
In this section, we lower bound the probability of safety violation achieved by a specific attack. The attack is the same as described in Appendix H, except that all H-blocks mined after $s-2\Delta$ are maximally delayed by $\Delta$, instead of after $s-\Delta$. Accordingly, the first new pacer is the first H-block mined after $s-2\Delta$.
Hence the probability that the adversary wins is lower bounded by
(261) $P$ $\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right).$
It is not difficult to see that (261) is further lower bounded by
$P\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
(262) $-P\left(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
The second term in (J) is easily determined as
(263) $P(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{t+\Delta}=0)$
(264) $=e^{-h(t+2\Delta)}.$
It remains to bound the first term in (J). Since $L_{s-2\Delta}$, $A_{s-2\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we can replace $A_{s-2\Delta,d}$ with identically distributed and independent $A^{\prime}_{0,d-s+2\Delta}$ and replace $P^{\prime}_{d-s+\Delta}$ with $P^{\prime\prime}_{0,d-s+2\Delta}$ in (261) to obtain
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(265) $=P(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A^{\prime}_{0,d-s+2\Delta}\geq P^{\prime\prime}_{0,d-s+2\Delta}\right\})$
(266) $=P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime}_{t+\Delta,q+2\Delta}\geq P^{\prime\prime}_{0,t+2\Delta}+P^{\prime\prime}_{t+2\Delta,q+2\Delta}\right\}).$
Evidently, $A^{\prime}_{t+\Delta,q+2\Delta}$ is identically distributed as $A^{\prime\prime}_{0,q-t+\Delta}$. Moreover, as a delayed renewal processes with identical inter-arrival times, $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ has statistically a larger delay than $P^{\prime}_{\Delta,q-t+\Delta}$. Thus $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ is stochastically dominated by $P^{\prime\prime}_{\Delta,q-t+\Delta}=P^{\prime\prime}_{0,q-t+\Delta}$. Hence (266) implies that
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(267) $\geq P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime\prime}_{0,q-t+\Delta}\geq P^{\prime}_{0,t+2\Delta}+P^{\prime\prime}_{0,q-t+\Delta}\right\})$
(268) $\geq P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$
where $M$ is defined in (249). The pmf of $M$ equals (11). Similar to (li2021close, ), $P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$ can be calculated using the pmfs of the relevant independent random variables. Thus, given a latency $t$, the safety of a given transaction can be violated with a probability no less than given in (27).

### v5

@@
-Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus
-Abstract.
-This paper delves into the fundamental trade-off between security, latency, and throughput in proof-of-work longest-chain-fork-choice protocols, also known as the Nakamoto consensus.
-New upper and lower bounds on the probability of violating transaction safety are derived as a function of honest and adversarial mining rates, an upper bound on block propagation delays, and transaction confirmation latency, both in time and in block depth.
-The results include a first closed-form finite-latency bound applicable to all delays and mining rates up to the ultimate fault tolerance.
-Notably, the gap between the upper and lower bounds is significantly narrower than the best gaps previously established for Bitcoin and its derivatives such as Litecoin and Dogecoin, as well as for Ethereum Classic.
-Furthermore, the paper reveals a fundamental trade-off between transaction throughput and confirmation latency, ultimately determined by the desired fault tolerance and the growth of block propagation delay as block size increases.
-The authors are with the Department of Electrical and Computer Engineering, Northwestern University, Evanston, Illinois. Email: {shujie.cao, dguo}@northwestern.edu.
-copyright: none
-1. Introduction
-Nakamoto consensus, the celebrated consensus mechanism used in Bitcoin, centers around proof-of-work (PoW) mining and the longest-chain-fork-choice rule.
-In particular, a transaction can be committed once it is confirmed by a sufficiently long chain or after a sufficient amount of time.
-For example, a widely adopted rule of thumb is to accept a Bitcoin transaction after six confirmations.
-The safety of a transaction may be violated due to forks caused by block propagation delays and/or adversarial mining.
-Suppose the honest miners collectively mine $h$ blocks per second and the adversary mines $a$ blocks per second.
-Suppose also once a block is known to one honest miner, it is known to all honest miners within $\Delta$ seconds.
-In 2020, Dembo et al.
-(dembo2020everything, ) and Gaži et al.
-(gazi2020tight, ) used very different methods to determine the ultimate asymptotic fault tolerance for the Nakamoto consensus.
-Specifically, as long as [^1] The condition is expressed as $\beta<(1-\beta)/(1+(1-\beta)\lambda\Delta)$ in (dembo2020everything, ), where $\lambda=a+h$ denotes the total mining rate and $\beta=a/(a+h)$ denotes the fraction of adversarial mining.
-(1) $\frac{1}{a}>\Delta+\frac{1}{h},$
-every transaction is safe with arbitrarily high probabilityafter sufficient confirmation; conversely, if $1/a<\Delta+1/h$, no transaction is ever safe.
-How safe is a transaction after a given confirmation latency?
-Gaži, Ren, and Russell (gazi2022practical, ) provided a safety guarantee via numerically evaluating a Markov chain, which is only practical for small confirmation latencies and parameters that are significantly lower than the fault tolerance.
-By analyzing races between renewal processes, Li, Guo, and Ren (li2021close, ) presented a closed-form result focusing on confirmation by time.
-Using a reduction technique, Guo and Ren (guo2022bitcoin, ) gave a simple closed-form latency–security guarantee focusing on confirmation by depth, which is numerically effective across a wide range of parameters, although there remains a gap between their fault tolerance and (1).
-In this paper, we put forth a continuous-time model for block mining processes, blockchains, and safety violation.
-We develop a new closed-form latency–security trade-off for the Nakamoto consensus which is the first to guarantee safety for all parameter values of $a$, $h$, and $\Delta$ within the ultimate fault tolerance (1).
+Security, Latency, and Throughput of Proof-of-Work Nakamoto Consensus
+Abstract
+This paper investigates the fundamental trade-offs between block safety, confirmation latency, and transaction throughput of proof-of-work (PoW) longest-chain fork-choice protocols, also known as PoW Nakamoto consensus.
+New upper and lower bounds are derived for the probability of block safety violations as a function of honest and adversarial mining rates, a block propagation delay limit, and confirmation latency measured in both time and block depth.
+The results include the first non-trivial closed-form finite-latency bound applicable across all delays and mining rates up to the ultimate fault tolerance.
+Notably, the gap between these upper and lower bounds is narrower than previously established bounds for a wide range of parameters relevant to Bitcoin and its derivatives, including Litecoin and Dogecoin, as well as Ethereum Classic.
+Additionally, the study uncovers a fundamental trade-off between transaction throughput and confirmation latency, ultimately determined by the desired fault tolerance and the rate at which block propagation delay increases with block size.
+I Introduction
+Nakamoto consensus is a foundational protocol in blockchain technology, widely adopted across cryptocurrencies to enable decentralized agreement on the blockchain’s state.
+In these systems, participants organize incoming transactions into blocks, with each block pointing to a unique preceding block, thereby forming a linked chain of blocks.
+Bitcoin and some of its derivatives use a proof-of-work (PoW) Nakamoto consensus mechanism, where miners continually compete to solve complex cryptographic puzzles; solving a puzzle grants a miner the right to add a new block to the system.
+The reader is referred to [1, 2] for detailed descriptions of the consensus mechanism.
+A fork occurs when two or more blocks point to the same parent block.
+Forks can arise from adversarial mining or naturally due to block propagation delays, leading honest miners to have differing views of the blockchain state.
+Nakamoto consensus is built around the longest-chain fork-choice rule, which mandates that an honest miner: 1) always mines at the tip of a longest chain in its view, and 2) immediately shares any new chain extensions with other miners through a peer-to-peer network.
+The most widely adopted rule is to commit a block once it is confirmed by a sufficient number of subsequent blocks.
+A rule of thumb for Bitcoin is six block confirmations.
+As an alternative, a block on a longest chain may be committed if a sufficient amount of time has lapsed since the block was seen.
+In either scenario, the safety of a committed block could still be violated if another chain, which excludes that block, is to emerge as a longest chain at any subsequent time.
+Since such a violation presents a double-spending opportunity with potentially grave consequences, it is crucial to make the probability of violation very small under the adversary’s strongest feasible attacks.
+[…]
+$\frac{1}{a}>\Delta+\frac{1}{h},$ (1)
+every block is safe with arbitrarily high probability after sufficient block confirmations, regardless of the adversary’s attack strategy; conversely, if $1/a<\Delta+1/h$, no block is ever safe in the sense that a well-known attack can violate its safety with probability 1, regardless of the number of block confirmations.
+If (1) answers the “capacity” question, then a “finite-blocklength” question ensues: Suppose the parameters satisfy (1), how safe is a committed block after $k$ block confirmations?
+Gaži, Ren, and Russell [5] provided a safety guarantee via numerically evaluating a Markov chain, which is only practical for small confirmation depths and parameters that are far away from the ultimate fault tolerance.
+By analyzing races between renewal processes, Li, Guo, and Ren [6] presented a closed-form result focusing on confirmation by time.
+Guo and Ren [7] introduced a rigged model, where honest blocks mined within $\Delta$ seconds of any previous blocks are flipped to become adversarial blocks.
+With this additional advantage, the adversary’s optimal attack reduces to the well-understood private mining attack.
+A simple closed-form latency-security guarantee is given in [7], although the rigging technique also weakens the fault tolerance from (1) to $\beta<1-(1/2)e^{\lambda\Delta}$, where $\lambda=a+h$ denotes the total mining rate and $\beta=a/(a+h)$ denotes the fraction of adversarial mining.
+Doger and Ulukus [8] refined the rigged model of [7] by not flipping honest blocks which see no other honest blocks within the $\Delta$ period before its mining, thereby improving the security-latency bounds.
+The same authors generalized their study based on the rigged model to the cases of different network delay assumptions in [9] and [10].
+However, all of the preceding rigged models transfer some honest mining rate to the adversary, yielding strictly positive losses in the fault tolerance, i.e., (1) is not achieved.
+In this paper, we carefully define a continuous-time model for block mining processes, blockchains, and safety violation for blocks.
+We develop a new closed-form latency-security trade-off for the Nakamoto consensus which is the first to achieve the ultimate fault tolerance (1).
 The new results here also substantially reduce the previous gap between upper and lower bounds on the latency-security trade-off in many scenarios of interest.
-adversary new lower new upper (guo2022bitcoin, ) upper (gazi2022practical, ) upper
-10% 6 7 7 7
-20% 13 14 17 15
-30% 32 35 42 N/A
-40% 122 161 187 N/A
-49.2% 26720 52807 $\infty$ N/A
-Table 1. Bounds on confirmation depth for 99.9% safety. $a+h=1/600$ (blocks/second); $\Delta=10$ seconds.
-In Table 1 we compare our new bounds to the best prior bounds on the minimum confirmation depth that guarantees 99.9% safety.
+$\beta$ new lower new upper [7] [8] [5]
+10% 6 8 7 7 7
+20% 13 16 17 15 15
+30% 32 39 42 38 N/A
+40% 122 156 187 165 N/A
+49.2% 26720 44024 $\infty$ $\infty$ N/A
+TABLE I: Bounds on confirmation depth for 99.9% safety. $\lambda=1/600$ (blocks/second); $\Delta=10$ seconds.
+In Table I we compare our new bounds to the best prior bounds on the minimum confirmation depth that guarantees 99.9% safety.
 The total mining rate is assumed to be Bitcoin’s 1 block every 600 seconds.
@@
 Our results encompass various adversarial mining fractions from $10\%$ to $49.2\%$.
-We include bounds for 10% and 20% adversarial mining taken directly from (gazi2022practical, ), which provides no result for higher adversarial mining rates because the numerical method therein is only feasible for parameters that are significantly distant from the ultimate fault tolerance.
-Evidently, the new upper bound is much closer to the lower bound than prior bounds.
-Moreover, the new bounds apply to all parameter values satisfying (1), which in this case is up to $a/(a+h)=49.8\%$.
-In contrast, the fault tolerance in (guo2022bitcoin, ) is $49.16\%$, hence for $49.16\%$, it is indicated by $\infty$ in Table 1.
-The inverse of the right hand side of (1) is known to be the minimum growth rate of honest blocks’ heights (pass2017analysis, ; dembo2020everything, ; gazi2020tight, ).
+We include bounds for 10% and 20% adversarial mining taken directly from [5], which provides no result for higher adversarial mining rates because the numerical method therein is infeasible for parameters that are close to the fault tolerance.
+Evidently, the new upper bound is much closer to the lower bound than prior bounds in challenging cases.
+Moreover, the new bounds apply to all parameter values satisfying (1), which in this case is up to $\beta=49.8\%$.
+In contrast, the fault tolerance in [7] is $49.16\%$, hence for $49.2\%$, it is indicated by $\infty$ in Table I.
+The inverse of the right-hand side of (1) is known to be the minimum growth rate of honest blocks’ heights [11, 3, 4].
 If this growth rate exceeds the adversarial mining rate (as in (1)), the chance that more adversarial blocks are mined to overwhelm honest blocks to create a violation is very small for a confirmation latency sufficiently high.
 However, even if fewer adversarial blocks are mined than the height growth of honest blocks (i.e., the adversary has a deficit), a violation is still possible when honest miners with different views extend different chains to balance each other’s blocks on enough heights (i.e., the deficit is covered by those “balanced heights”).
-Examining the number of balanced heights offers a novel approach to the latency-security problem.
+This paper develops a novel approach to the latency-security problem by examining the number of balanced heights.
 Upon each opportunity to have a balanced height, there is a chance that all honest miners fail to balance, rather they synchronize to a common highest chain.
-Since honest blocks’ height growth rate exceeds the adversarial mining rate, there is a good chance that the adversary is left behind permanently, so no alternative highest chain is ever produced to cause a violation of the target transaction.
+Since honest blocks’ height growth rate exceeds the adversarial mining rate, there is a good chance that the adversary is left behind permanently, so no alternative highest chain is ever produced to cause a violation of the target block.
 Using this insight, we show that the number of balanced heights is stochastically dominated by essentially a geometric random variable.
-Interestingly, the aforementioned synchronizing block fall under the category of “Nakamoto blocks” defined in (dembo2020everything, ).
-Without using the complicated machinery of trees and tines as in (dembo2020everything, ; gazi2022practical, ), our analysis yields finite-latency, easily computable bounds that are tighter than the best existing results.
-In particular, the new bounds extend to the ultimate fault tolerance established in (dembo2020everything, ; gazi2022practical, ).
+Interestingly, the aforementioned synchronizing block falls under the category of “Nakamoto blocks” defined in [3].
+Without using the complicated machinery of trees and tines as in [3, 5], our analysis yields tight, easily computable finite-latency security bounds which achieve and imply the ultimate fault tolerance (1).
 The tight latency-security trade-off enables a user to choose an efficient confirmation latency that aligns with their desired security level.
-There has been ample evidence (e.g., (fechner2022calibrating, )) that the block propagation delay is approximately an affine function of the block size.
-In this paper, we explore the trade-off between the total throughput in kilobytes (KB) per second and the confirmation latency (in seconds), both of which depend on the block size and mining rate.
+There has been ample evidence (e.g., [12]) that the block propagation delay is approximately an affine function of the block size.
+In this paper, we explore the trade-off between the total throughput in kilobytes (KB) per second and the confirmation latency (in minutes), both of which depend on the block size and mining rate.
 This paper provides improved guidance for selecting parameters that optimize key performance indicators.
 The remainder of this paper is organized as follows: Sec.
-2, introduces a mathematical model for the Nakamoto consensus.
+II, introduces a mathematical model for the Nakamoto consensus.
 Sec.
-3 explains why the model is sufficient for our purposes.
+III explains why the model is sufficient for our purposes.
 Sec.
-4 presents the main theorems.
+IV presents the main theorems.
 Sec.
-5 introduces the most essential elements of the proof, followed by a sketch of the analysis.
+V introduces the most essential elements of the proof, followed by a sketch of the analysis.
 Sec.
-6 presents the numerical results, showcasing the practical implications.
-Due to space limitations, detailed proofs are relegated to the appendices.
-2. Model
+VI presents the numerical results, showcasing the practical implications.
+Detailed proofs are relegated to the appendices.
+II Model
 In this section, we put forth a mathematical model for block mining processes and blockchains, enabling a precise definition of key concepts such as confirmation, latency, and safety violations.
 In Sec.
-3, we will delve into why the model is sufficient for the latency-security analysis.
-2.1. Block and Chain
-We consider a continuous-time model, where transactions can be created and blocks can be mined at arbitrary times. Throughout this paper, “by time $t$” means “during $(0,t]$”.
+III, we will delve into why the model is sufficient for the latency-security analysis.
+II-A Block and Chain
+We consider a continuous-time model, where blocks can be mined at arbitrary times. Throughout this paper, “by time $t$” means “during $(0,t]$”.
 Definition 1 (block).
@@
 The parent of block 1 is the genesis block.
-We let $t_{b}$ denote the time when block $b$ is mined.
-Definition 2 (transaction).
-Transactions are created sequentially in time as transaction 1, transaction 2, and so on.
-Block $b$ is associated with an array $B_{b}$, which contains the indexes of some transactions that are created strictly before time $t_{b}$.
-We say transaction $j$ is included in block $b$ if $j$ is in the array $B_{b}$.
-A transaction can be included in multiple blocks.
+We use $t_{b}$ to denote the time when block $b$ is mined.
 Definition 2 (blockchain and height).
-A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$, block $b_{i-1}$ is block $b_{i}$’s parent for every $i=1,\dots,n$, and no transaction is included in two or more of those blocks.
-Since every chain can be uniquely identified by its last block, we also refer to the preceding chain as chain $b_{n}$.
+A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$, and for every $i=1,\dots,n$, block $b_{i-1}$ is block $b_{i}$’s parent.
 As a convention, we say the height of this chain is $n$.
+Since every chain can be uniquely identified by its last block, we often refer to the preceding chain as chain $b_{n}$.
+Also, we say block $b_{i+1}$ extends chain $b_{i}$ and that chain $b_{i}$ is a prefix of chain $b_{n}$ for every $i=0,\dots,n-1$.
 In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$.
@@
 […]
-An example of transactions, blocks, and chains.
-Figure 1.
-An example of transactions, blocks, and chains.
-Nonpublic blocks are shaded.
-Each adversarial block is marked by a pair of horns on top.
-Below each block is an array of transactions included in the block.
+Figure 1: An example of blocks and chains. Nonpublic blocks are shaded. Each adversarial block is marked by a pair of horns on top.
 Fig.
 1 illustrates an example of chains formed by blocks 0 to 10 along with a time axis.
-It depicts a snapshot at $t_{10}$ when block $10$ is mined.
+It depicts a snapshot right after $t_{10}$ when block $10$ is mined.
 Each block is horizontally positioned at its mining time and is connected by a line to its unique parent to its left and to its children to its right (if any).
-Under each block is its associated array describing the transactions included in the block.
 Chain $2$’s height is $h_{2}=2$; chain $5$’s height is $h_{5}=3$.
@@
 Even though block $6$ is mined after block $5$, chain 5 is higher than chain $6$.
-Block $7$ or chain $7$ is the highest at its mining time.
-Block $1$ is associated with $B_{1}=[2,5]$, i.e., it includes transactions $2$ and $5$.
-Transaction $4$ is included in both blocks 2 and 3.
-We next define two crucial attributes of transactions, blocks, and chains.
-Definition 4 (public).
-A transaction, block, or chain may become public at a specific point in timestrictly after its creation and will remain public at all times afterward.
-A chain being public at $t$ is equivalent to all of its blocks being public at $t$.
-While all transactions included in a public block are public, a transaction that is public at time $t$ may or may not be included in any block mined by $t$.
-As a convention, the genesis block becomes public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$.
+Blocks $9$ and $10$ or chains $9$ and $10$ are both the highest at $t_{10}$.
+Definition 3 (public and public height).
+A chain may become public at a specific time strictly after its creation and will remain public at all times afterward.
+If a chain is public, all of its prefix chains must also be public.
+As a convention, the genesis block is public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$.
+At time $t$, the height of a highest public chain is referred to as the public height.
+We use the notion of being public to capture the situation that a chain is known to all honest nodes.
 Every chain consists of a public prefix chain followed by some nonpublic blocks (possibly none).
@@
 Chain 5 is the public prefix of chain 9.
-If chain 9 becomes public at a later time, then block 7 must also be public by that time.
-Transaction 8 may be nonpublic at $t_{10}$, while transaction 11 is public since it is included in public block 6.
+If chain 9 becomes public at a later time, then chain 7 must also be public by that time.
 Definition 4 (credible).
-A chain is said to be credible at time $t$, or $t$-credible, if it is no shorter than any public chain at $t$. All blocks in a credible chain are said to be credible.
+A chain mined by time $t$ is said to be credible at $t$, or $t$-credible, if it is no lower than the public height at $t$.
 Being public or not is an inherent attribute of a chain, while credibility is a different attribute determined by the set of public chains.
@@
 In the snapshot at $t_{10}$ depicted in Fig.
-1, chains 4 and 5 are both the highest public chains, making them both $t_{10}$-credible.
-Chain 6 is lower than chain 5, so it is not $t_{10}$-credible.
+1, chain 8 is the highest public chain.
+Chains 4, 5, 6 are lower than chain 8, so they are not $t_{10}$-credible.
 Chains 7, 8, 9 and 10 are all $t_{10}$-credible.
-If chain 9 becomes public at a later time, then chain 8 will no longer be credible, regardless of whether it is public at that time.
-2.2. Confirmation and Safety
-Definition 6 (Confirmation by time $t$).
+If chain 10 becomes public at a later time, then chain 7 will no longer be credible, regardless of whether it is public at that time.
+II-B Confirmation and Safety
+Definition 5 (confirmation by time).
 Let the confirmation time be fixed and denoted by $t>0$.
-Suppose a transaction becomes public at time $s>0$.
-We say the transaction is confirmed (by time $t$) by chain $c$ at time $s+r$ if chain $c$ is public and credible at $(s+r)$ with $r\geq t$ and contains a block which includes the transaction.
-In Fig. 1, fix confirmation time $t$. Target transaction $4$ becomes public at $s$. Transaction $4$ (in block 2) is confirmed by chain 8 at $t_{10}$, which is $t_{8}$-credible, with $t_{8}>s+t$.
-Definition 7 (confirmation by depth $k$).
+Given a time $s$, if a chain $g$ is mined by $s$ and is a prefix of a $(s+r)$-credible chain $c$ with $r\geq t$, then we say chain $c$ confirms chain $g$ by time of $t$.
+An example is depicted in Fig.
+1 with $s$ and $s+t$ marked on the time axis.
+Suppose chain 1 is mined by time $s$.
+Since chain 7 is credible at $t_{7}>s+t$, it confirms chain 1 by time of $t$.
+Likewise, chain 8 confirms chain 2 by time of $t$.
+Definition 6 (confirmation by depth).
 Let the confirmation depth be a fixed natural number denoted by $k$.
-Suppose a transaction becomes public at $s>0$.
-Let block $d$ be the highest public block at time $s+\Delta$.
-If there exists a credible public chain $c$, which contains a block $b$ that includes the transaction, where $h_{c}\geq h_{b}+k-1$ and $h_{c}\geq h_{d}+k$, then we say chain $c$ confirms the transaction by depth $k$.
-Fix $k=3$.
+If a block $g$ is included in a credible chain $b$ with $h_{b}\geq h_{g}+k-1$, then we say chain $b$ confirms block $g$ by depth of $k$ or block $g$ is $k$-confirmed by chain $b$.
 In Fig.
-1, transaction 4 is included in block 2 ($h_{2}=2$), which is contained in chain 8 ($h_{8}=4$).
-The highest public block by $s$ is block 1 with $h_{1}=1$ (assuming block 2 is not yet public at $s$).
-Since chain 8 is public and credible at $t_{10}$, $h_{8}\geq h_{2}+k-1$, and $h_{8}\geq h_{2}+k$, we say transaction 4 is confirmed (by depth $k$) by chain 8.
-With confirmation by either time or depth, we can define the violation of a target transaction’s safety as follows:
-Definition 8 (safety violation of a target transaction).
-Suppose a target transaction becomes public at $s>0$. If there exist $r$ and $r^{\prime}$ with $r^{\prime}\geq r$ and also:
-  • a chain $c$ that is public and credible at time $r$ which confirms the transaction in block $b$, and
-  • an $r^{\prime}$-credible chain which does not contain block $b$,
-then we say the target transaction’s safety is violated.
-In Fig.
-1, target transaction $4$’s safety is violated as it is confirmed by $t_{10}$-credible and public chain 8 (in block 2), while $t_{10}$-credible chain 7 does not contain block 2.
-Note that if two credible chains confirm the same target transaction but in different blocks, its safety is still violated.
-2.3. Adversary and Propagation Delay
+1, for example, block 2 is 3-confirmed by chain 8, and block 3 is 3-confirmed by chain 7 and 4-confirmed by chain 9.
+With confirmation by either time or depth, we can define the violation of a block’s safety as follows:
+Definition 7 (safety violation of a block).
+If there exist $r$ and $r^{\prime}$ with $r^{\prime}\geq r$ as well as
+  • an r-credible chain $c$ which confirms block $g$, and
+  • an $r^{\prime}$-credible chain which does not include block $g$,
+then we say block $g$’s safety is violated.
+In Fig. 1, block 3 is 3-confirmed by $t_{7}$-credible chain 7, but is not included in chain 8, which is $t_{8}$-credible. Thus block 3’s safety is violated.
+II-C Adversary and Propagation Delay
 Definition 8 (types of blocks).
-A block is either an A-block or an H-block. The parent of an H-block must be credible when the H-block is mined.
+A block other than the genesis block is either an A-block or an H-block. The parent of an H-block must be credible when the H-block is mined.
 An A-block (resp.
@@
 honest) miner.
-According to Definitions 5 and 9, an H-block must be higher than all public blocks at the time of its mining.
-There is no such height constraint on an A-block.
+According to Definitions 4 and 8, an H-block must be higher than all public chains at the time of its mining.
+There is no such constraint on an A-block.
 In fact, this is the only constraint on block heights.
@@
 Blocks 2 and 7 are A-blocks, while all the other blocks are H-blocks.
-Definition 10 (transaction’s inclusion).
-Suppose a transaction becomes public at time $s$. If block $b$ is the first H-block mined after $s$, then chain $b$ must include the transaction.
-The propagation delays of transactions, blocks, and chains depend on the network conditions and may also be manipulated by the adversary.
-To establish a strong safety guarantee, we adopt the widely used $\Delta$-synchrony model.
+The propagation delays of blocks and chains depend on the network conditions and may also be manipulated by the adversary.
+To establish a strong safety guarantee, we adopt the widely used $\Delta$-synchrony model, which is defined by the following constraint:
 Definition 9 (chain delay bound $\Delta$).
@@
 after $t_{b}+\Delta$.
-Definition 10 guarantees that once a transaction becomes public, it will be promptly included in at least one chain.
-Together, Definitions 10 and 11 imply that, if a transaction becomes public at $s$, no H-block mined before time $s-\Delta$ includes it, while it may be included in an A-block mined before $s-\Delta$.
 Definition 10 (Poisson mining processes).
@@
 Let $H_{s,t}$ denote the number of H-blocks mined during $(s,t]$.
-Evidently, $(A_{0,t}+H_{0,t})_{t\geq 0}$ is a Poisson point process with rate $a+h$.
+Evidently, $(A_{0,t}+H_{0,t})_{t\geq 0}$ is a Poisson point process with rate
+$\lambda=a+h.$ (2)
 As a convention, let $A_{s,t}=H_{s,t}=0$ for all $s>t$.
-Since the probability that any two blocks are mined at exactly the same time is equal to zero, we assume without loss of generality that all blocks are mined at different times.
-Hence $t_{0}<t_{1}<t_{2}<\dots$.
-3. Model Sufficiency
-3.1. Sufficiency
-The prevailing model in this line of work (see, e.g., (garay2015bitcoin, ; pass2017analysis, ; pass2017rethinking, ; garay2017bitcoin, ; dembo2020everything, ; gazi2020tight, ; saad2021revisiting, ; neu2022longest, ; gazi2023practical, ; kiffer2023security, )) is based on the universally composable security framework (canetti2000universally, ).
-A complete description of the model, henceforth referred to as the view-based model, requires the definition of a protocol as a set of algorithms, the execution of those algorithms that is directed by an environment, which activates $n$ either honest or corrupted parties with individual views (pass2017analysis, , Sec.
-2).
+Since the probability that any two or more blocks are mined at exactly the same time is equal to zero, we assume without loss of generality that all blocks are mined at different times.
+Hence $0=t_{0}<t_{1}<t_{2}<\dots$.
+III Model Sufficiency
+III-A Sufficiency
+The prevailing model in this line of work (see, e.g., [13, 11, 14, 15, 3, 4, 16, 17, 5, 18]) is based on the view-based model.
+Such a model defines a protocol as a set of algorithms executed in an environment that activates $n$ parties, which may be either honest or corrupted, each with its own view [11, Sec.
+2].
 The protocol’s execution is no simpler than a system of $n$ interactive Turing machines.
-Due to its complexity, the view-based model always assumes discrete time steps in execution.
+The view-based model often assumes discrete time steps in execution.
 The model proposed in Sec.
-2 of this paper is entirely self-contained and much simpler than the view-based model.
-In contrast to previous works as well as (li2021close, ; guo2022bitcoin, ), our model intentionally exclude miners and their views in definitions.
+II of this paper is entirely self-contained and much simpler than the view-based model.
+In contrast to previous works as well as [6, 7], our model intentionally excludes miners and their views in definitions.
 In particular, with well-defined block mining processes, we are not concerned with the number of miners; neither do we keep track of which blocks an individual miner has mined or seen.
-Instead, our model keeps track of the set of public chains at every point in time, which represents the weakest possible view of any miner.
+Instead, our model keeps track of the set of public chains at every point in time, which represents the weakest possible view to any miner.
 In particular, every honest view is a superset of this weakest view, and conversely, the adversary in the view-based model can ensure at least one honest miner holds that weakest view.
 Relative to the view-based model under $\Delta$-synchrony, our model has no loss of generality in the sense that, for a given realization (or sample path) of the A- and H-block mining processes, if the view-based model can produce a certain set of chains, so can our model.
-This is because both models respect the same constraints for chain creation described in Definitions 9 and 11.
-For the latency-security analysis, our model is as powerful as the view-based model (if simpler).
+This is because both models respect the same constraints for chain creation described in Definitions 8 and 9.
+For the latency-security analysis, our model is as powerful as the view-based model.
 To elaborate, consider the event in the view-based model where, first, a chain in an honest view confirms a block on height $h$, and then another chain in an honest view confirms a different block on the same height, creating a safety violation.
@@
 Once we realize the role of the weakest view, there is no need to define individual views and their complicated interactions.
-3.2. Operations
+III-B Operational Meaning
 For concreteness and practicality, we relate the definitions in Sec.
-2 to the operations of miners (or nodes), who assemble transactions into blocks and chains by expending proof of work.
-The same miners also propagate those chains through a peer-to-peer network.
-The honest miners follow the longest-chain-fork-choice rule whereas the adversary can use an arbitrary (possibly randomized) strategy.
-At time $t$, each node has its own view, which includes all transactions, blocks, and chains the node has seen by $t$.
-A node’s view can only increase with time.
-The operations of each node is based on the entire history of this node’s view.
-In general, an honest node will only provide goods and services for a transaction after it is considered committed.
-If an honest node first encounters a transaction at time $s$, it can deduce that the transaction will become public during the time window of $[s,s+\Delta]$.
-Here we define two commit rules corresponding to the two types of confirmation defined in Sec.
-2.
-Specifically, if the transaction is included in a highest chain in this node’s view at time $s+r$ for some $r\geq t$, and the chain remains the highest at $s+r+\Delta$, then the node can infer that 1) the chain is credible and public at $s+r+\Delta$, and 2) the transaction has been confirmed by time $t$ in the mathematical model.
-In this case, the node commits the transaction by time $t$.
-Similarly, if block $d$ is the highest block in this node’s view at time $s+\Delta$, and there is a chain $c$ in the node’s view that remains the highest for a period of $\Delta$ or longer, which contains a block $b$ including the transaction, and $h_{c}\geq h_{b}+k-1$ as well as $h_{c}\geq h_{d}+k$, then the node can infer that the transaction has been confirmed by depth $k$ in the mathematical model.
-In this case, the node commits the transaction by depth $k$.
-We will show that, regardless of the adversary’s strategy, violating the safety of a transaction implies that an “error event” concerning the two mining processes occurs.
-We then bound the probability of safety violation defined in Sec.
-2, which implies a probabilistic guarantee for an honest node’s operational safety.
-3.3. Delay Variations
+II to the operations of an individual miner (or node), who assembles transactions into blocks and then chains by expending proof of work.
+If the node is honest, it follows the longest-chain fork-choice rule and immediately share new blocks via the peer-to-peer network; otherwise, it can use an arbitrary (possibly randomized) attack strategy.
+At time $t$, every node has its own view, which includes all transactions, blocks, and chains the node has seen by $t$.
+The operations of this node is based on the entire history of this node’s view, which can only increase with time.
+In general, an honest node will provide goods and services for a transaction only after it commits the transaction.
+As committing a transaction has no bearing on chain production, nodes are free to choose their individual commit rules.
+A popular choice is to commit after $k$ block confirmations, where each node can select its own depth number $k$.
+In operations, after an honest node sees a new block $g$ extending one of the highest chains in the node’s view, the node commits all new transactions in all blocks in chain $g$ as soon as any chain $b$ with height $h_{b}\geq h_{g}+k-1$ becomes a highest chain in its view and has chain $g$ as a prefix.
+[^2] In case both block $g$ and its child block $c$ enter the node’s view at once, the node can commit transactions in chain $g$ first according to this rule, and then commit additional transactions in chain $c$ in the same manner.
+We emphasize that block $g$ is $k$-confirmed by chain $b$ by Definition 6.
+In other words, committed transactions must be $k$-confirmed.
+An alternative commit rule is by time of $t$: A node focuses on a block $g$ in its view at (the current) time $s$.
+If there ever is a time $r\geq t$, such that chain $g$ is a prefix of a highest chain in this node’s view at time $s+r$, then the node commits all new transactions contained in chain $g$.
+By Definition 5, since this highest chain at time $s+r$ is credible, chain $g$ is confirmed by time of $t$.
+With confirmation by either depth or time, once we have a safety guarantee established rigorously for sufficiently confirmed chains in the mathematical model in Sec.
+II, the honest node enjoys the same operational safety guarantee for every committed transaction.
+There is in fact the subtlety that block $g$ is anchored to a current view at a point in time chosen by the honest node, so that the adversary cannot directly carry out an adaptive attack of the following form: Mine in private to gain a sufficiently high lead, then pick a block to attack with certainty to cause a violation.
+III-C Delay Variations
 The seemingly rigid $\Delta$-synchrony model is quite practical.
@@
 Second, if some blocks of honest miners suffer longer delays than $\Delta$, we simply regard those blocks as adversarial blocks to preserve $\Delta$-synchrony.
-In addition, potential attacks like Erebus (Erebus, ) and SyncAttack (SyncAttack, ) seek to partition the peer-to-peer network.
+In addition, potential attacks like Erebus [19] and SyncAttack [20] seek to partition the peer-to-peer network.
 As long as those delay violations are independent of the block mining times, the honest and adversarial mining processes remain to be Poisson.
-If block propagation delays have a long-tailed distribution (decker2013information, ; fechner2022calibrating, ), we can use a delay bound that applies to a high percentage of blocks.
+If block propagation delays have a long-tailed distribution [21, 12], we can use a delay bound that applies to a high percentage of blocks.
 The main results here continue to apply with all out-of-sync mining power treated as being adversarial.
-In Definition 10, we have assumed that every public transaction gets included in a chain very soon.
-In a more general model, a transaction’s inclusion can be delayed due to censorship or a lack of space in new blocks.
-If this is the case, we modify the confirmation rules as follows: First, block $b$ which contains the target transaction becomes public.
-Then, if block $b$ is in a credible public chain $c$ with $h_{c}\geq h_{b}+k-1$, we say chain $c$ confirms the transaction by depth $k$.
-If block $b$ is in a credible public chain $c$ at time $t_{b}+r$ with $r\geq t$, we say chain $c$ confirms the transaction by time $t$.
-We note that the main theorems in the next section can be easily generalized to this case, although we omit the details in this paper.
-4. Main Results
+IV Main Results
 Recall that $a$ and $h$ represent as the adversarial and honest mining rates, respectively, in the number of blocks mined per second.
-For confirmation by either time or depth, the following negative result is well understood (see, e.g., (pass2017analysis, ; dembo2020everything, ; gazi2020tight, )).
+For confirmation by either time or depth, the following negative result is well understood (see, e.g., [11, 3, 4]).
 Theorem 1 (unachievable fault tolerance).
 If the mining rates and the delay upper bound satisfy
-(2) $\frac{1}{a}<\Delta+\frac{1}{h},$
-then every transaction is unsafe in the sense that the private-mining attack on the transaction will violate its safety with probability 1.
+$\frac{1}{a}<\Delta+\frac{1}{h},$ (3)
+then every block is unsafe in the sense that the private-mining attack will violate its safety with probability 1.
 For completeness, we prove Theorem 1 in Appendix A.
-4.1. Bounds for Given Confirmation Depth
+IV-A Bounds for Given Confirmation Depth
 Let us define
-(3) $b(\alpha,\eta)=\left(\frac{e^{4\gamma}\gamma(\eta-\alpha-\alpha\eta)}{\alpha\eta e^{\gamma}-(\alpha+\gamma)(\eta-\gamma)}\right)^{2}+1-\frac{\eta-\alpha-\alpha\eta}{\eta e^{\alpha}},$
+$b(\alpha,\eta)=\left(\frac{e^{\frac{5}{2}\gamma}\gamma(\eta-\alpha-\alpha\eta)}{\alpha\eta e^{\gamma}-(\alpha+\gamma)(\eta-\gamma)}\right)^{2}+1-\frac{\eta-\alpha-\alpha\eta}{\eta e^{\alpha}},$ (4)
 and
-(4) $\begin{split}c&(\alpha,\eta)=\left(\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma\right)\\ &\cdot\left(1-\ln\left(1+\frac{\gamma}{\alpha}\right)\bigg{/}\ln\frac{\eta e^{\alpha}(e^{\eta}-1)}{\eta+\eta e^{\alpha}(e^{\eta}-1)-\alpha(1+\eta)}\right)^{-1}\end{split}$
+$\begin{split}c(\alpha,\eta)=\left(\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma\right)\cdot\left(1-\ln\left(1+\frac{\gamma}{\alpha}\right)\bigg{/}\ln\frac{\eta e^{\alpha}(e^{\eta}-1)}{\eta+\eta e^{\alpha}(e^{\eta}-1)-\alpha(1+\eta)}\right)^{-1}\end{split}$ (5)
 where $\gamma$ in (4) and (5) is a shorthand for
-(5) $\gamma(\alpha,\eta)=\frac{2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}}{2}.$
-The following latency–security trade-off is the first set of non-trivial bounds that apply to all parameter values up to the ultimate fault tolerance.
-Theorem 2 (Upper bounding the violation probability for given confirmation depth).
+$\gamma(\alpha,\eta)=\frac{2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}}{2}.$ (6)
+The following latency-security trade-off is the first set of non-trivial safety guarantees that apply to all parameter values up to the ultimate fault tolerance.
+Theorem 2 (Safety guarantee for given confirmation depth).
 Suppose the mining rates and the delay upper bound satisfy
-(6) $\frac{1}{a}>\Delta+\frac{1}{h}.$
-Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
-(7) $b(a\Delta,h\Delta)e^{-c(a\Delta,h\Delta)\cdot k}$
+$\frac{1}{a}>\Delta+\frac{1}{h}.$ (7)
+Given a sufficiently large $s$, the safety of a block extending an $s$-credible chain and being $k$-confirmed can be violated with a probability no higher than
+$b(a\Delta,h\Delta)e^{-c(a\Delta,h\Delta)\cdot k}$ (8)
 where $b(\cdot,\cdot)$ and $c(\cdot,\cdot)$ are defined in (4) and (5), respectively.
@@
 Corollary 3 (Achievable fault tolerance).
-As long as (6) holds, every transaction is arbitrarily safe after sufficiently long confirmation latency (either in depth or time).
-The pair of opposing inequalities (2) and (6) determine the exact boundary between the asymptotically achievable and unachievable latency–security trade-offs.
+As long as (7) holds, every block is arbitrarily safe after sufficiently long confirmation latency (either in depth or time).
+The pair of opposing inequalities (3) and (7) determine the exact boundary between the asymptotically achievable and unachievable latency-security trade-offs.
 The right hand side of (3) and (7) is exactly the expected inter-arrival time of the following type of blocks:
@@
 We use $P_{s,t}$ to denote the number of pacers mined in $(s,t]$.
-By definition, pacers become public by their immediate successors and hence all pacers are on different heights.
+By definition, every pacer becomes public before the next pacer arrives, and hence all pacers are on different heights.
 Since the expected pacer inter-arrival time is $\Delta+1/h$, pacers arrive at the rate of $h/(1+h\Delta)$ in the long run.
@@
 Let us introduce the following probability density function (pdf):
-(8) $f(t)=\frac{1}{h}e^{-h(t-\Delta)}1_{\{t>\Delta\}}.$
-We also consider a delayed renewal process in general, which is like an ordinary renewal process, except that the first arrival time (delay) may have a different distribution than the other inter-arrival times.
-Lemma 4 ((Kroese_difference, , Theorem 3.4)).
+$f(t)=\frac{1}{h}e^{-h(t-\Delta)}1_{\{t>\Delta\}}.$ (9)
+Lemma 4 ([22, Theorem 3.4]).
 Let $A_{r}$ denote a Poisson point process with rate $a$.
@@
 Then the probability generating function of the maximum difference between the two processes
-(9) $Y=\sup_{t\geq 0}\{A_{t}-Q_{t}\}$
+$Y=\sup_{t\geq 0}\{A_{t}-Q_{t}\}$ (10)
 is equal to
-(10) $\mathcal{E}(r)=\frac{(1-r)(h-a-ha\Delta)}{h-e^{(1-r)a\Delta}(h+a-ar)r}.$
+$\mathcal{E}(r)=\frac{(1-r)(h-a-ha\Delta)}{h-e^{(1-r)a\Delta}(h+a-ar)r}.$ (11)
 We denote the probability mass function (pmf) of $Y$ as
-(11) $e(i)$ $=\frac{1}{i!}\mathcal{E}^{(i)}(0),\qquad i=0,1,\dots$
+$e(i)$ $=\frac{1}{i!}\mathcal{E}^{(i)}(0),\qquad i=0,1,\dots$ (12)
 where $\mathcal{E}^{(i)}(0)$ represents the the $i$-th derivative of $\mathcal{E}(\cdot)$ evaluated at $r=0$.
 Throughout this paper, we let
-(12) $\epsilon$ $=1-e^{-h\Delta}$
+$\epsilon$ $=1-e^{-h\Delta}$ (13)
 and
-(13) $\delta$ $=\left(1-a\frac{1+h\Delta}{h}\right)^{+}e^{-a\Delta}.$
+$\delta$ $=\left(1-a\frac{1+h\Delta}{h}\right)^{+}e^{-a\Delta}.$ (14)
 Let us define
-(14) $F^{(k)}(n)=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}\bar{F}_{b}(n;k,(\epsilon+\delta-\epsilon\delta))$
+$F^{(k)}(n)=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}\bar{F}_{b}(n;k,(\epsilon+\delta-\epsilon\delta))$ (15)
 where $\bar{F}_{b}(\cdot;k,p)$ denotes the complementary cumulative distribution function (cdf) of the binomial distribution with parameters $(k,p)$. We also define
-(15) $F(n)$ $=\lim_{k\to\infty}F^{(k)}(n)$
-(16) $=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}.$
+$F(n)$ $=\lim_{k\to\infty}F^{(k)}(n)$ (16)
+$=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}.$ (17)
 Theorem 5 (Finer upper bound for the violation probability for given confirmation depth).
-Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
-$\inf_{n}\bigg{\{}2-$ $\sum_{i=0}^{k-n}e(i)\sum_{j=0}^{k-n-i}e(j)\int_{0}^{\infty}F_{1}(k-n-i-j;a(t+6\Delta))$
-(17) $\cdot f_{2}(t-(k)\Delta;k,h)\,dt-F^{(k)}(n-1)\bigg{\}}$
+Given a sufficiently large $s$, the safety of a block that extends an $s$-credible chain and is confirmed by the depth of $k$ can be violated with a probability no higher than
+$\inf_{n}\bigg{\{}2-$ $\sum_{i=0}^{k-n}e(i)\sum_{j=0}^{k-n-i}e(j)\int_{0}^{\infty}F_{1}(k-n-i-j;a(t+3\Delta))\cdot f_{2}(t-(k)\Delta;k,h)\,dt-F^{(k)}(n-1)\bigg{\}}$ (18)
 where $F^{(k)}$ is defined in (15), $e(i)$ is given by (12), $F_{1}(\cdot;\lambda)$ denotes the cdf of the Poisson distribution with mean $\lambda$, and $f_{2}(\cdot;k,h)$ denotes the pmf of the Erlang distribution with shape parameter $k$ and rate parameter $h$.
@@
 A well-understood attack is the private-mining attack, which we will use to lower bound the probability of safety violation. The attack is defined precisely as follows.
-Definition 14 (Private-mining attack on a target transaction).
+Definition 12 (Private-mining attack on a target block).
 The adversary delays the propagation of every H-block maximally (for $\Delta$ seconds).
-Starting from time 0, the adversary always mines in private to extend the highest chain which does not include the target transaction.
-The adversary makes public the private chain as soon as both of the following conditions hold: (1) the transaction has been confirmed by a chain of H-blocks; and (2) the private chain which does not contain the transaction is credible.
+Starting from time 0, the adversary always mines in private to extend the highest chain which does not include the target block.
+Recall that Definition 7 allows nonpublic credible chains to cause a violation.
+However, operationally, the adversary can carry out a private-mining attack and make the private chain public as soon as both of the following conditions hold: 1.
+the block has been confirmed by a chain of H-blocks; and 2.
+the private chain which does not contain the block is credible.
+In other words, as long as a safety violation occurs by Definition 7, the adversary creates a violation operationally.
 Theorem 6 (Lower bounding the violation probability for given confirmation depth).
-Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
+Given a sufficiently large $s$, the private-mining attack violates the safety of a target block which extends an $s$-credible chain with at least the following probability:
 […]
-$\cdot\sum_{j=0}^{k-i}e(j)\bigg{(}f_{1}(0,h\Delta)\int_{0}^{\infty}F_{1}(k-(1+i+j);a(t-\Delta))$
-$\cdot f_{2}(t-k\Delta;k,h)\,dt+(1-f_{1}(0,h\Delta))$
-(18) $\cdot\int_{0}^{\infty}F_{1}(k-(i+j);a(t-\Delta))\cdot f_{2}(t-k\Delta;k,h)\,dt\bigg{)}$
 where $e(\cdot)$, $F_{1}$, $f_{2}$ are defined as in Theorem 5 and $f_{1}(\cdot;\lambda)$ denotes the pmf of the Poisson distribution with mean $\lambda$.
 Theorem 6 is proved in Appendix H.
-4.2. Throughput and Related Trade-offs
-An ideal protocol design should select parameters to yield high throughput, fast confirmation, and security against adversary miners.
+IV-B Throughput and Related Trade-offs
+An ideal protocol design should select parameters to yield high throughput, fast confirmation, and security against adversarial miners.
 We refer to the average number of blocks mined during the maximum propagation delay, $\lambda\Delta$, as the natural fork number.
 Throughout the paper, let $\beta$ denote the maximum tolerable fraction of adversarial mining.
-The natural fork number is bounded by
-(19) $\lambda\Delta<\frac{1}{\beta}-\frac{1}{1-\beta}$
-due to the fault tolerance (1).
+To operate within fault tolerance (1), the natural fork number must be bounded by
+$\lambda\Delta<\frac{1}{\beta}-\frac{1}{1-\beta}.$ (20)
+We remark that it is more convenient to discuss throughput in terms of $\beta$ and the total mining rate $\lambda$ in lieu of the adversarial and honest mining rates ($a$ and $h$).
 To achieve a high throughput (in KB/second or transactions/second), the designer may select a high mining rate and/or a large block size, where the latter also implies a large block propagation delay.
@@
 […]
-Empirically, the block propagation delay is approximately an affine function of the block size (decker2013information, ; fechner2022calibrating, ) (even with a coded design (zhang2022speeding, )).
-That is, the propagation delay bound can be modeled as
-(21) $\Delta(B)=\frac{B}{r}+\nu$
+Empirically, the block propagation delay is approximately an affine function of the block size [21, 12] (even with a coded design [23]). That is, the propagation delay bound can be modeled as
+$\Delta(B)=\frac{B}{r}+\nu$ (22)
 where $B$ is the block size in KB and $r$ and $\nu$ are constants.
 Here, $r$ represents a fundamental limitation on the end-to-end transmission rate of the peer-to-peer network.
-Based on (20), this rate and the desired fault tolerance determines an upper bound for the optimistic throughput:
-(22) $\frac{\lambda B}{1+\lambda\Delta}<\frac{1-2\beta}{1-\beta-\beta^{2}}\,r.$
-With a close upper bound on the probability of safety violation as a function of the system parameters and the confirmation latency, e.g., in the form of $p(\lambda\Delta,\beta,k)$ where a user may select the confirmation depth $k$ to achieve a desired safety.
-It is well understood that a user making a very large transaction is generally willing to tolerate a longer confirmation latency.
+Based on (21), this rate and the desired fault tolerance determine an upper bound for the optimistic throughput:
+$\frac{\lambda B}{1+\lambda\Delta}<\frac{1-2\beta}{1-\beta-\beta^{2}}\,r.$ (23)
+With a close upper bound on the probability of safety violation as a function of the system parameters and the confirmation latency, e.g., in the form of $p(\lambda\Delta,\beta,k)$, a node may select the confirmation depth $k$ to achieve a desired safety.
+In particular, the creator of a highly-valued transaction will most likely tolerate a longer-than-usual confirmation latency.
 The theorems in this paper allow this choice to be made much more accurately than before.
@@
 In particular, we can formulate the following problem:
-(23a) $\text{maximize}_{\lambda,B,k}\quad$ $\frac{\lambda B}{1+\lambda\,\Delta(B)}$
-(23b) subject to $p(\lambda\,\Delta(B),\beta,k)\leq q$
-(23c) $k\leq\frac{(1-\beta)\lambda}{1+(1-\beta)\lambda\Delta}d$
+$\text{maximize}_{\lambda,B,k}\quad$ $\frac{\lambda B}{1+\lambda\,\Delta(B)}$ (24a)
+subject to $p(\lambda\,\Delta(B),\beta,k)\leq q$ (24b)
+$k\leq\frac{(1-\beta)\lambda}{1+(1-\beta)\lambda\Delta}d$ (24c)
 where $q$ denotes the maximum tolerable probability of violation and $d$ denotes the expected confirmation time.
@@
 In Sec.
-6, we give such a numerical example under assumption (21) with empirically determined constants.
-4.3. Bounds for Given Confirmation Time
-Theorem 8 (Upper bound for given confirmation time).
+VI, we give such a numerical example under assumption (22) with empirically determined constants.
+IV-C Bounds for Given Confirmation Time
+Theorem 8 (safety guarantee for given confirmation time).
 Suppose the mining rates and the delay upper bound satisfy (1).
-Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by time $t$ can be violated with a probability no higher than
-(24) $\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+G(n,t)\},$
+Given a sufficiently large $s$, the safety of a block mined by time of $s$ and is confirmed by time of $t$ can be violated with a probability no higher than
+$\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+G(n,t)\},$ (25)
 where $F$ is defined in (16), and
-(25) $\begin{split}&G(n,t)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ &\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
+$G(n,t)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))$ (26)
 where $e(\cdot)$ and $f_{1}$ are defined the same as in Theorem 6, and $F_{2}(\cdot;n,h)$ denotes the cdf of the Erlang distribution with shape parameter $n$ and rate parameter $h$.
 Theorem 8 is proved in Appendix I. In lieu of the infinite sums in (26), a computationally feasible bound is obtained by replacing $G(n,t)$ with its following upper bound
-(26) $\begin{split}\overline{G}(n,t)&=1-F_{1}(i_{0};a(t+2\Delta))F_{E}(j_{0})F_{E}(k_{0})\\ &+\sum_{i=0}^{i_{0}}\sum_{j=0}^{j_{0}}\sum_{k=0}^{k_{0}}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ \times(1-&F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
+$\begin{split}\overline{G}(n,t)&=1-F_{1}(i_{0};a(t+2\Delta))F_{E}(j_{0})F_{E}(k_{0})\\ &+\sum_{i=0}^{i_{0}}\sum_{j=0}^{j_{0}}\sum_{k=0}^{k_{0}}f_{1}(i;a(t+2\Delta))e(j)e(k)\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$ (27)
 where $F_{1}$, $F_{E}$ are the corresponding cdfs of variables with pmf of $f_{1}(i;a(t+2\Delta))$, $e(j)$ in (12), respectively.
@@
 The larger these integers are, the tighter the upper bound.
-For completeness, we include the following converse result, which is a variation of (li2021close, , Theorem 3.6) due to the slightly different definition of safety violation here.
+For completeness, we include the following converse result, which is a variation of [6, Theorem 3.6] due to the slightly different definition of safety violation here.
 Theorem 9 (Lower bound for given confirmation time).
-Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
-(27) $\begin{split}&\left(1-\frac{a}{h}\right)e^{(h-a)(t+2\Delta)}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}e(j)\left(\frac{a}{h}\right)^{k}F_{1}(k;h(t+2\Delta))\\ &\;\times(1-F_{2}(t+\Delta-(j+k)\Delta;j+k+1,h))-e^{-h(t+2\Delta)}.\end{split}$
+Given a sufficiently large $s$, the private-mining attack violates the safety of a target block mined by time $s$ and confirmed by time of $t$ with at least the following probability:
+$\left(1-\frac{a}{h}\right)e^{(h-a)t}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}e(j)\left(\frac{a}{h}\right)^{k}F_{1}(k;ht)\times(1-F_{2}(t-(j+k)\Delta;j+k+1,h))-e^{-h(t-\Delta)}.$ (28)
 Theorem 9 is proved in Appendix J. Evidently, if the infinite sums in (28) are replaced by partial sums, the resulting (tighter) security level also remains unachievable.
-5. Essential Elements of the Analysis
+V Essential Elements of the Analysis
 In this section, we introduce the most essential technical elements, discuss intuitions, and provide a sketch of the statistical analysis.
-5.1. The Balancing Process
+V-A The Balancing Process
 Definition 13 (Agree and disagree).
-Two chains are said to agree on a target transaction if they both contain a common block which includes the transaction, or neither includes the transaction; otherwise, they are said to disagree on the target transaction.
-We also say blocks $b$ and $d$ agree (resp.
-disagree) on the target transaction if chains $b$ and $d$ agree (resp.
-disagree) on the target transaction.
-Agreement about a target transaction is an equivalence relation, which typically divide all blocks into two equivalence classes.
+Two chains that are no lower than a target block are said to agree on the target block if either they both include the target block or neither includes it; otherwise, they are said to disagree on the target block.
+Agreement on a target block is an equivalence relation, which divides all blocks higher than the target block into two equivalence classes.
 In Fig.
-1, however, because transaction 4 is included in both blocks 2 and 3, its agreement divides all blocks into three equivalence classes, namely, $\{0,1,6\}$, $\{2,4,8,10\}$, and $\{3,5,7,9\}$.
+1, agreement about target block 3 divides blocks on height 2 and above into two equivalence classes, namely, blocks $\{3,5,7,9\}$ and blocks $\{2,4,6,8,10\}$.
 Definition 14 (balancing candidate, candidacy).
-Given a target transaction, if an H-block (denoted as block $b$) is the first block mined on height $n$ and disagrees on the target transaction with any height-$(n-1)$ block mined before block $b$ becomes public, then block $b$ is said to be a balancing candidate, or simply a candidate.
+Given a target block, if an H-block (denoted as block $b$) is the first block mined on height $n$ and disagrees on the target block with any height-$(n-1)$ block mined before block $b$ becomes public, then block $b$ is said to be a balancing candidate, or simply a candidate.
 The period from the first time the disagreement arises to the time block $b$ becomes public is referred to as its candidacy.
@@
 Definition 15 (balanced height).
-If during a balancing candidate’s candidacy, a disagreeing H-block is mined on the candidate’s height and no adversarial block is mined on that height, then the candidate is said to be balanced.
-Accordingly, its height is said to be a balanced height.
-We use $X_{s,t}$ to denote the total number of balanced candidates whose candidacy is finished during $(s,t]$. With slight abuse notation, we let
-(28) $X=\lim_{t\to\infty}X_{0,t}.$
+If a disagreeing H-block is mined on the candidate’s height and no A-block is mined on that height, then the candidate is said to be balanced. Accordingly, its height is said to be a balanced height.
+Given a target block $g$, we use $X_{s,t}^{g}$ to denote the total number of balanced candidates created during $(s,t]$. We also define $X^{g}_{t}=X^{g}_{t_{g},t}$ and
+$X^{g}=\lim_{t\to\infty}X_{t}^{g}.$ (29)
 […]
-Figure 2.
-Two balancing candidates $b$ and $b^{\prime}$ at different times.
+Figure 2: Balancing candidates $b$ and $b^{\prime}$ for target block $g$ and $g^{\prime}$, respectively.
 Blocks on the same height are connected by dashed lines.
-Left: $b$ becomes a balancing candidate upon its mining; Right: $b^{\prime}$ becomes a candidate after $c$ is mined on a lower height to create a disagreement.
-The duration of each candidacy is marked red on the time axis.
-The duration of the first of those candidacies is exactly $\Delta$.
-The duration of the second one is strictly less than $\Delta$.
+Left: block $b$ becomes a balancing candidate upon its mining; its candidacy is $[t_{b},t_{b}+\Delta]$.
+Right: block $b^{\prime}$ becomes a candidate after a lower block $c$ is mined, in disagreement with block $b^{\prime}$.
+Block $b^{\prime}$ has a shorter candidacy $[t_{c},t_{b^{\prime}}+\Delta]$.
 A candidacy is a small window of opportunity to balance the corresponding candidate (and thus create a balanced height).
@@
 1.
-With regard to target transaction 4, blocks 4 and 5 balance each other on height 3; blocks 9 and 10 balance each other on height 5.
+If block 3 is the target, blocks 4 and 5 balance each other on height 3; blocks 9 and 10 balance each other on height 5.
 No other heights are balanced.
 In particular, while H-blocks 3 and 6 are on the same height, they do not balance each other due to the existence of block 2 as an A-block on the same height.
-We shall show that the total number of heights that can be balanced (denoted as $X$) is essentially stochastically dominated by a geometric random variable.
-5.2. Violation Event under Confirmation by Depth
+We shall show that the total number of heights that can be balanced is essentially stochastically dominated by a geometric random variable.
+V-B Violation Event under Confirmation by Depth
 Definition 16 (Pre-mining lead).
-The pre-mining lead (of the adversary) at time $t$, denoted by $L_{t}$, is the height of the highest block mined by time $t$ minus the height of the highest H-block mined by time $t$.
-The notion of pre-mining lead is to account for the potential lead the adversary may already have when the target transaction becomes public at around time $s$.
+The pre-mining lead (of the adversary) at time $t$, denoted by $L_{t}$, is the height of the highest block mined by time $t$ minus the public height at time $t$.
+The notion of pre-mining lead is to account for the potential lead the adversary may already have when the target block is included by a blockchain.
 In this work, $s$ is given so that the adversary cannot mount an adaptive attack (by influencing $s$) depending on the lead.
@@
 Definition 17 (Jumper).
-We say the genesis block is the 00-th jumper. After the genesis block, every jumper is the first H-block mined on its height. We use $J_{s,t}$ to denote the number of jumpers mined in $(s,t]$.
-In the remainder of this section, we assume $s$ is a given time. We focus on the case of confirmation by depth $k$.
+The first H-block mined on its height is called a jumper. We use $J_{s,t}$ to denote the number of jumpers mined in $(s,t]$.
+In the remainder of this section, we assume $s$ is a given time.
+We focus on the case of confirmation by depth $k$.
 It is easy to see that the height growth of H-blocks is no less than the number of jumpers mined.
-Intuitively, for the transaction’s safety to be violated, there must exists an interval, during which the number of jumpers mined is fewer than the sum of the lead, the number of balanced heights, and the number of A-blocks mined.
+Intuitively, for a block’s safety to be violated, there must exist an interval, during which the number of jumpers mined is fewer than the sum of the pre-mining lead, the number of balanced heights, and the number of A-blocks mined.
 This is made precise in the following lemma.
-Let $\tau$ denote the first time a block is mined that is $k$ or more blocks higher than the highest public block at time $s+3\Delta$.
 Lemma 10.
-Suppose a target transaction becomes public during $(s,s+2\Delta]$. If this target transaction’s safety is violated, then the following event must occur:
-(29) $L_{s}+X_{s,\tau}+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k.$
+Given $s$, let $\tau$ denote the arrival time of the first H-block that is at least $k$ higher than the public height at time $s$.
+If a block $g$ extends an $s$-credible chain and its safety is violated, then the following must hold:
+$L_{s}+X_{t_{f}+\Delta}^{g}+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k$ (30)
+where block $f$ is the first block mined on height $h_{g}+k-1$, whose mining time is denoted as $t_{f}$.
 Proof.
-Suppose target transaction is public at time $s^{\prime}$, block $b^{\prime}$ is the highest public block at time $s^{\prime}+\Delta$.
-Let $\tau^{\prime}$ denote the first time a block is mined at height $h_{b^{\prime}}+k$ or above.
-The highest public H-block at time $d$ is at least $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$.
-If a violation occurs at time $d$, there must exist two chains disagreeing with each other on the target transaction, where both chains are no shorter than $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$.
-At time $s^{\prime}$, since the highest H-block is no higher than $h_{b^{\prime}}$, the highest A-block is no higher than $h_{b^{\prime}}+L_{s^{\prime}}$.
-Since $X_{s^{\prime},d}$ heights are balanced, A-blocks need to be mined at least on the remaining $k-X_{s^{\prime},d}$ heights after $h_{b^{\prime}}+L_{s^{\prime}}$.
-Along with A-blocks mined during $(s^{\prime},d]$, a violation implies the following,
-(30) $h_{b^{\prime}}+L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}$ $\geq h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}.$
-By Definition 18 of the pre-mining lead,
-(31) $L_{s}+X_{s,d}+A_{s,d}\geq L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}.$
-By definition of $\tau$, we have
-(32) $k+J_{\tau^{\prime},d-\Delta}\geq k+J_{\tau,d-\Delta}.$
-One way a violation occurs is for a balancing candidate to arise (regardless of whether the height is subsequently balanced). Thus, if a violation occurs at time $d>\tau^{\prime}$, we must have
-(33) $X_{\tau^{\prime},d}=0.$
-Thus, we have
-(34) $L_{s}+X_{s,\tau}+A_{s,d}\geq k+J_{\tau,d-\Delta}.$
-Therefore, a violation can only occur if (29) occurs. ∎
+We first show that (30) is implied by a slightly stronger inequality:
+$L_{s}+X_{t_{f}+\Delta}^{g}+\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k,$ (31)
+where the domain of optimization $d\geq\tau$ in (30) is replaced by $d\geq t_{f}$.
+In the case of $t_{f}\geq\tau$, the domain in (30) is larger, yielding a higher supremum value, hence (30) holds as long as (31) holds.
+In the remaining case of $t_{f}<\tau$, we have
+$\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}$ $=\max\left\{\sup_{d>\tau}\{A_{s,d}-J_{\tau,d-\Delta}\},\sup_{\tau\geq d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\right\}$ (32)
+$=\max\left\{\sup_{d>\tau}\{A_{s,d}-J_{\tau,d-\Delta}\},A_{s,\tau}\right\}$ (33)
+$=\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}$ (34)
+where in (33) and (34) we have both used the fact that $J_{\tau,d-\Delta}=0$ for $d\leq\tau$. Hence, in this case (30) and (31) are equivalent.
+The remainder of this proof is devoted to establishing (31).
+Here we let all agreement relationship be with respect to block $g$.
+Evidently, block $g$ can not be $k$-confirmed before $t_{f}$.
+Let chain $b$ denote a highest public chain by time $s$.
+Since block $g$ extends an $s$-credible chain, we have $h_{g}>h_{b}$ and hence
+$h_{f}\geq h_{b}+k$ (35)
+by definition of block $f$ in the lemma.
+Let block $c$ denote a highest H-block mined by $t_{f}$.
+Let chain $z$ denote a highest chain that disagrees with chain $c$ at time $t_{f}$.
+Definitions of some frequently used notations in this proof are listed in Table II.
+Notation Meaning
+$s$ a given time for commitment rule
+chain $b$ a highest public chain by $s$
+block $c$ a highest H-block mined by $t_{f}$
+block $f$ the first block mined on height $h_{g}+k-1$
+block $g$ the target block
+chain $z$ a highest chain that disagrees with chain $c$ at $t_{f}$
+$\tau$ the arrival time of the first H-block that is at least $k$ higher than chain $b$
+TABLE II: Frequently used notations in the proof of Lemma 10.
+We first establish the following upper bound on the number of jumpers mined during $(\tau,t_{f}]$:
+$J_{\tau,t_{f}}\leq h_{f}-(h_{b}+k).$ (36)
+If $\tau\geq t_{f}$, $J_{\tau,t_{f}}=0$ by convention, so that (36) holds trivially due to (35).
+If $\tau<t_{f}$, the jumpers between $\tau$ and $t_{f}$ must be between heights $h_{b}+k+1$ and $h_{f}$, hence (36) must hold.
+We establish (31) in two separate cases depending on whether $h_{c}\leq h_{b}+L_{s}$.
+  • Case 1)
+    $h_{c}\leq h_{b}+L_{s}$.
+In this case no H-blocks higher than $h_{b}+L_{s}$ are mined by $t_{f}$, so there is at least one A-block on each height from $h_{b}+L_{s}+1$ to $h_{f}$ mined between $s$ and $t_{f}$, i.e.,
+    $A_{s,t_{f}}$ $\geq h_{f}-(h_{b}+L_{s})$ (37)
+    $\geq k+J_{\tau,t_{f}}-L_{s},$ (38)
+    where (38) is due to (36).
+Since $X_{t_{f}+\Delta}^{g}\geq 0$ and $\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq A_{s,t_{f}}-J_{\tau,t_{f}-\Delta}\geq A_{s,t_{f}}-J_{\tau,t_{f}}$, (38) implies (31).
+  • Case 2)
+    $h_{c}>h_{b}+L_{s}$.
+We divide the proof in this case into two parts (2-a) and (2-b), where the first part focuses on blocks mined during $(s,t_{f}]$, and the second part focuses on blocks mined after $t_{f}$.
+    • Part 2-a)
+      Since chains $c$ and $z$ disagree, there exist at least two disagreeing chains of height $\min\{h_{c},h_{z}\}$ by $t_{f}$.
+This implies that at least one A-block is mined during $(s,t_{f}]$ on every height between $h_{b}+L_{s}+1$ and $\min\{h_{c},h_{z}\}$ except those heights that are balanced by time $t_{f}$.
+The number of balanced heights is no larger than $X_{t_{f}}^{g}$.
+In addition, since no H-block is higher than $h_{c}$, at least one A-block is mined on each height from $h_{c}+1$ to $h_{f}$.
+Therefore, the number of A-blocks is lower bounded:
+      $A_{s,t_{f}}\geq(\min\{h_{c},h_{z}\}-(h_{b}+L_{s})-X_{t_{f}}^{g})+(h_{f}-h_{c}).$ (39)
+    • Part 2-b)
+      We now focus on blocks mined after $t_{f}$ in two different cases depending on the relative positions of blocks $c$ and $z$.
+      • Case $h_{z}\geq h_{c}$.
+        Since $\min\{h_{c},h_{z}\}=h_{c}$, (39) reduces to
+        $A_{s,t_{f}}$ $\geq h_{f}-(h_{b}+L_{s})-X_{t_{f}}^{g}.$ (40)
+        Comparing (36) and (40) yields
+        $L_{s}+X_{t_{f}}^{g}+A_{s,t_{f}}-J_{\tau,t_{f}}$ $\geq k.$ (41)
+        Noting that $X^{g}_{t_{f}}\leq X^{g}_{t_{f}+\Delta}$, $J_{\tau,t_{f}}\geq J_{\tau,t_{f}-\Delta}$, and reducing the domain $d\geq t_{f}$ in (31) to $d=t_{f}$ reduces the supremum, we have established (31) based on (41) for Case $h_{z}\geq h_{c}$.
+      • Case $h_{z}<h_{c}$.
+        In this case, block $c$ must agree with block $f$; otherwise chain $z$ cannot be the highest to disagree with chain $c$.
+Since block $f$ is the first block that is $k-1$ higher than block $g$, whose safety is subsequently violated, the first chain that is credible and also disagrees with chains $c$ and $f$ must be created at some time $d\geq t_{f}$ to create a safety violation.
+All jumpers mined during $(t_{f},d-\Delta]$ become public by time $d$, so the highest public block at time $d$ is at least as high as $h_{c}+J_{t_{f},d-\Delta}$.
+Because the highest disagreeing chain at $t_{f}$ is no higher than $h_{z}$, those blocks of the $d$-credible disagreeing chain between heights $h_{z}+1$ and $h_{c}+J_{t_{f},d-\Delta}$ must be mined during $(t_{f},d]$.
+Moreover, after $t_{f}+\Delta$, a violation occurs as soon as a balancing candidate is created (before it is even balanced).
+Hence at most $X_{t_{f},t_{f}+\Delta}^{g}$ of heights $(h_{z}+1,\dots,h_{c}+J_{t_{f},d-\Delta})$ can be balanced, whereas at least one A-block is mined on the remaining heights.
+Putting it altogether, we have
+        $A_{t_{f},d}$ $\geq h_{c}+J_{t_{f},d-\Delta}-h_{z}-X_{t_{f},t_{f}+\Delta}^{g}.$ (42)
+        Adding (39) with $\min\{h_{c},h_{z}\}=h_{z}$ and canceling $h_{z}$ and $h_{c}$ on both sides, we have
+        $A_{s,d}$ $\geq J_{t_{f},d-\Delta}+h_{f}-(h_{b}+L_{s})-X_{t_{f}+\Delta}^{g}$ (43)
+        $\geq J_{t_{f},d-\Delta}+J_{\tau,t_{f}}+k-L_{s}-X_{t_{f}+\Delta}^{g}$ (44)
+        $=J_{\tau,d-\Delta}+k-L_{s}-X_{t_{f}+\Delta}^{g},$ (45)
+        where (44) is due to (36). Since $\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}$ in (31) is greater or equal to $A_{s,d}-J_{\tau,d-\Delta}$, (45) implies (31).
+We caution that Lemma 10 is very general and at the same time quite delicate.
+While block $b$ in the proof is the highest public block by $s$, block $g$ can be arbitrarily higher than height $h_{b}+1$, and may or may not be mined after $s$.
+In fact, even block $f$ may or may not be mined before $s$, as there is the possibility that chain $f$ already $k$-confirms block $g$, yet all of those confirming blocks including block $g$ itself are nonpublic.
+We illustrate (30) in Lemma 10 using a pair of concrete chains in Fig 3.
+$d$$t_{f}$$\tau$$s$$b$$g$$c$$e$$z$$v$
+Figure 3: An example of (30) in Lemma 10.
+Let the depth of confirmation $k=5$.
+Blocks $b$ and $e$ are on the same height.
+Block $b$ is the highest public chain at $s$ (block $g$’s parent is not yet public).
+The lead $L_{s}=2$.
+Here $c$ and $f$ refer to the same block, i.e., $c=f$.
+At time $t_{f}$, block $g$ is $k$-confirmed by the highest chain $c$, as $h_{c}=h_{g}+k-1$.
+The highest disagreeing chain $z$ at $t_{f}$ is on height $h_{z}=h_{g}+2<h_{c}$.
+The (first) violation occurs at time $d=t_{v}$, when chain $v$ is mined, which is credible as $h_{v}=h_{c}$.
+Height $h_{g}+1$ is balanced, with $X^{g}_{t_{f}}=1$.
+A single jumper is mined during $(\tau,d-\Delta]$.
+Three A-blocks are mined during $(s,d]$.
+Therefore, we have $L_{s}+X_{t_{f}+\Delta}^{g}+A_{s,d}-J_{\tau,d-\Delta}=5=k$, i.e., (30) holds in this case.
 Based on Lemma 10, we can characterize the latency-security trade-off by examining the competition between adversarial and honest nodes.
@@
 The challenge we must address is to decompose the event in Lemma 10 into separate, independent components.
-5.3. A Sketch of the Statistical Analysis
+V-C A Sketch of the Statistical Analysis
 The remaining task is to analyze the probability of the event described by (30). We need to understand the distribution of the individual elements as well as to account for their dependency.
@@
 Theorem 11 (Bounding the number of balanced heights).
-For every integer $n\geq 0$, the number of balanced heights up to the $k$-th candidacy, denoted as $X^{(k)}$, satisfies
-(35) $P(X^{(k)}\leq n)\geq F^{(k)}(n)$
-where $F^{(k)}$ is defined in (14). As a special case, the total number of balanced heights defined in (28) satisfies
-(36) $P(X\leq n)\geq F(n)$
-where
+With respect to a given target block, the number of balanced heights up to the $k$-th candidacy, denoted as $X^{(k)}$, satisfies
+$P(X^{(k)}\leq n)\geq F^{(k)}(n)$ (46)
+for every integer $n\geq 0$, where $F^{(k)}$ is defined in (15). As a special case, the total number of balanced heights defined in (29) satisfies
+$P(X\leq n)\geq F(n).$ (47)
 If (3) holds, then $\delta=0$, so $F(n)=0$ and (47) is vacuous.
@@
 In fact, with probability 1, one last balancing opportunity will be missed and the adversary is then behind future pacers permanently.
-The final balancing candidate serves the role of the “Nakamoto block” in (dembo2020everything, ) and is a stronger notion than the “convergence opportunity” in e.g., (pass2017rethinking, ).
+The final balancing candidate serves the role of the “Nakamoto block” in [3] and is a stronger notion than the “convergence opportunity” in e.g., [14].
 This observation underlies the safety guarantee of Theorem 2.
 Let us define
-(37) $Y^{(k)}=L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}.$
-While $Y^{(k)}$ and $X_{s,\tau}$ are dependent, we can upper bound the probability of the event (29), i.e., $X_{s,\tau}+Y^{(k)}\geq 0$, using a union bound:
+$Y^{(k)}=L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}.$ (48)
+While $Y^{(k)}$ and $X_{t_{f}+\Delta}^{g}$ are dependent, we can upper bound the probability of the event (30), i.e., $X_{t_{f}+\Delta}^{g}+Y^{(k)}\geq 0$, using a union bound:
 Lemma 12.
 Let $\bar{F}_{X}(\cdot)$ denote the complementary cdf of the random variable $X$. All integer-valued random variables $X$ and $Y$ satisfy
-(38) $P(X+Y\geq 0)\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X}(n-1)+\bar{F}_{Y}(-n)\}.$
+$P(X+Y\geq 0)\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X}(n-1)+\bar{F}_{Y}(-n)\}.$ (49)
 In addition, given two distributions of integer-valued random variables for which the right-hand side of (49) is strictly less than 1, there exists a joint distribution of $X$ and $Y$ that satisfy (49) with equality.
 The proof of Lemma 12 is relegated to Appendix C.
-It suffices to upper bound the marginal complementary cdf’s of $X_{s,\tau}$ and $Y^{(k)}$. Since all $X_{s,\tau}$ balanced heights must be within the first $k$ heights above $h_{b}$, we have
-(39) $P(X_{s,\tau}>n-1)\leq P(X^{(k)}>n-1)$
+It suffices to upper bound the marginal complementary cdf’s of $X_{t_{f}+\Delta}^{g}$ and $Y^{(k)}$.
+Since all $X_{t_{f}+\Delta}^{g}$ balanced heights must be within the first $k$ heights above $h_{b}$, we have
+$P(X_{t_{f}+\Delta}^{g}>n-1)\leq P(X^{(k)}>n-1)$ (50)
 where the distribution of $X^{(k)}$ is given in (15). Thus, we have
-$P(Y^{(k)}$ $+X_{s,\tau}\geq 0)$
-(40) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X_{s,\tau}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}$
-(41) $\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X^{(k)}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}.$
+$P(Y^{(k)}$ $+X_{t_{f}+\Delta}^{g}\geq 0)$
+$\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X_{t_{f}+\Delta}^{g}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}$ (51)
+$\leq\inf_{n\in\mathbb{Z}}\{\bar{F}_{X^{(k)}}(n-1)+\bar{F}_{Y^{(k)}}(-n)\}.$ (52)
 The remaining task is to understand the distribution of $Y^{(k)}$.
@@
 Lemma 13 (There are more jumpers than pacers).
-(42) $J_{s,t}\geq P_{s+\Delta,t},\quad\forall t>s.$
+$J_{s,t}\geq P_{s+\Delta,t},\quad\forall t>s.$ (53)
 Proof.
@@
 By Lemma 13, we have
-(43) $Y^{(k)}$ $\leq L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-P_{\tau+\Delta,d-\Delta}\}$
-(44) $=L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P_{\tau+\Delta,\tau+\Delta+c}\}.$
+$Y^{(k)}$ $\leq L_{s}-k+\sup_{d\geq\tau}\{A_{s,d}-P_{\tau+\Delta,d-\Delta}\}$ (54)
+$=L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P_{\tau+\Delta,\tau+\Delta+c}\}.$ (55)
 It is easy to see that the lead $L_{s}$, the A-block process from $s$ onward, and the pacer process from $\tau+\Delta$ onward are mutually independent.
@@
 Once the distribution of $Y^{(k)}$ is bounded, we can combine it with Theorem 11 to upper bound the probability of violation using (52). A detailed proof of Theorem 2 is given in Appendix D.
-5.4. Operational Implications
-A brief discussion on the operational implications of the theorems presented in Sec.
+V-D Operational Implications
+A brief discussion on the operational implications of theorems presented in Sec.
 2 is warranted.
@@
 Specifically, it is not selected by the adversary.
-If the adversary were to choose $s$, they could make a target transaction public once they have established a substantial lead (this might take an exceedingly long time, but it would eventually occur with a probability of 1).
+If the adversary were to choose $s$, it could choose a target block one height above the highest public height once they have established a substantial lead (this might take an exceedingly long time, but it would eventually occur with a probability of 1).
 By doing so, a safety violation would be guaranteed.
-In this paper, we eliminate such adaptive attacks by assuming that the transaction becomes public within a fixed time window that is beyond the adversary’s control.
-Operationally, we have the flexibility to select an arbitrary number of intervals $(s,s+2\Delta]$.
-If an honest node first encounters a transaction during $(s,s+\Delta]$ (who will also propagate this transaction), then the node can be sure that the transaction becomes public during the interval $(s,s+2\Delta]$.
-The node can be confident about the probabilistic safety guarantees for the transaction due to the theorems given in Sec.
-4.
+In this paper, we eliminate such adaptive attacks by assuming that the target block needs to extend a credible chain that is beyond the adversary’s control.
+Operationally, we have the flexibility to select an arbitrary time $s$.
+If an honest node sees the target block mined on top of a highest block in its view at time $s$, the node can be sure that the block is mined on $s$-credible chain, and hence satisfies the conditions in the theorems given in Sec.
+IV.
 Following the preceding outline of the statistical analysis, we prove the main theorems in detail in the appendices.
 In Sec.
-6, we present the numerical results derived from these theorems.
+VI, we present the numerical results derived from these theorems.
 These results demonstrate the practicality of our latency-security trade-off bounds and their utility as guidance of parameter selection for Nakamoto-style protocols.
-6. Numerical Result
+VI Numerical Result
 In this section, the system parameters are chosen to represent two different settings in most cases.
-The first setting is modeled after Bitcoin, where the total mining rate $h+a=1/600$ blocks/second.
+The first setting is modeled after Bitcoin, where the total mining rate $\lambda=1/600$ blocks/second.
 Bitcoin blocks’ propagation delays vary and have been decreasing over the years.
-As of October 2023, the 90th percentile of block propagation delay is about 4 seconds on average (bitcoinmoniter, ).
+As of November 2024, the 90th percentile of block propagation delay is about 4 seconds on average [24].
 We use $\Delta=10$ seconds as an upper bound.
-The second setting is modeled after Ethereum Classic (ETC), where $h+a=1/13$ blocks/second and $\Delta=2$ seconds.
+The second setting is modeled after Ethereum Classic (ETC), where $\lambda=1/13$ blocks/second and $\Delta=2$ seconds.
 We also include results for Litecoin and Dogecoin parameters in one table.
-6.1. Security Latency Trade-off
-Figure 3. Bitcoin’s safety versus confirmation time.
-Figure 4. ETC’s safety versus confirmation time.
+VI-A Security Latency Trade-off
+Figure 4: Bitcoin’s safety versus confirmation time.
+Figure 5: ETC’s safety versus confirmation time.
 In Fig.
-3, we first show the achievable and unachievable probability of safety violation as a function of the confirmation time (in hours) under the Bitcoin setting.
+4, we first show the achievable and unachievable probability of safety violation as a function of the confirmation time (in hours) for the Bitcoin configuration.
 For three different adversarial mining fractions ($\beta=1/10$, $1/7$, and $1/4$), we plot the upper and lower bounds calculated using Theorems 8 and 9, respectively.
-Also included for comparison is an achievable trade-off from (li2021close, , Eqn.
-(5)).
+Also included for comparison is an achievable trade-off from [6, Eqn.
+(5)].
 In Fig.
-4, we plot the latency-security trade-off of ETC for two different adversarial mining fractions ($\beta=1/7$ and $1/4$).
-The gap between the upper and lower bounds is smaller than the gap derived in (li2021close, ).
+5, we plot the latency-security trade-off of ETC for two different adversarial mining fractions ($\beta=1/7$ and $1/4$).
+In all cases, the new bounds approximately halve the latency gaps derived in [6].
 In Fig.
-5, we show Bitcoin’s safety bounds as a function of confirmation depth $k$ calculated using Theorems 25, and 6, respectively.
-Also included are the achievable and unachievable bounds from (guo2022bitcoin, , Eqns.
-(7) and (8)) for comparison.
-The new upper bound (5) of Theorem 5 yields substantially tighter result.
+6, we show Bitcoin’s safety bounds as a function of confirmation depth $k$ calculated using Theorems 25, and 6, respectively.
+Also included are the achievable and unachievable bounds from [8].
+The new upper bound (18) of Theorem 5 yields tighter results for small confirmation depths.
 Similarly, ETC’s safety bounds as a function of confirmation depth $k$ is plotted in Fig.
-6.
-For the ETC parameters, the achievable bounds in this paper significantly outperform the achievable bound of (guo2022bitcoin, , Eqn.
-(7)).
-Figure 5. Bitcoin’s safety versus confirmation depth.
-Figure 6. ETH’s safety versus confirmation depth.
+7.
+For the ETC parameters, the achievable bounds in this paper outperform the achievable bound of [8].
+Figure 6: Bitcoin’s safety versus confirmation depth.
+Figure 7: ETH’s safety versus confirmation depth.
 $\Delta$ 00 s $2.5$ s $10$ s $40$ s $160$ s $640$ s
-$\beta=10\%$ $7$ $7$ $7$ $10$ $21$ $52$
-$\beta=20\%$ $14$ $14$ $14$ $26$ $35$ $117$
-$\beta=30\%$ $34$ $35$ $35$ $63$ $91$ $633$
-$\beta=40\%$ $150$ $154$ $161$ $277$ $552$infty
-Table 2. The confirmation depth needed to guarantee $99.9\%$ safety for various block delays and adversarial fractions.
+$\beta=10\%$ $7$ $7$ $8$ $15$ $16$ $37$
+$\beta=20\%$ $14$ $16$ $16$ $19$ $28$ $65$
+$\beta=30\%$ $34$ $37$ $39$ $46$ $65$ $633$
+$\beta=40\%$ $150$ $153$ $156$ $277$ $552$infty
+TABLE III: The confirmation depth needed to guarantee $99.9\%$ safety for various block delays and adversarial fractions.
 For a fixed $99.9\%$ safety guarantee, using the Bitcoin parameters, we compute the confirmation depth needed according to Theorem 5 for various block propagation delays and adversarial mining fractions.
-Table 2 shows that the confirmation depth rises slowly with the delay bound when the adversarial fraction is relatively low; it rises more quickly when the adversarial fraction is relatively high.
+Table III shows that the confirmation depth rises slowly with the delay bound when the adversarial fraction is relatively low; it rises more quickly when the adversarial fraction is relatively high.
 $\bar{F}(0)$ $\bar{F}(1)$ $\bar{F}(2)$ $\bar{F}(3)$ $\bar{F}(4)$
@@
 ETC $0.06$ $0.008$ $0.001$ $2\times 10^{-4}$ $3\times 10^{-5}$
-Table 3. The complementary cdf, $\bar{F}(n)$.
-Table 3 shows the probability of having more than a given number of balanced heights with 25% adversarial mining.
+TABLE IV: The complementary cdf, $\bar{F}(n)$.
+Table IV shows the probability of having more than a given number of balanced heights with 25% adversarial mining.
 For example, the probability of having more than four balanced heights is less than $10^{-9}$ for Bitcoin.
 Hence the bound in Lemma 12 is fairly tight for our purposes.
-$\lambda^{-1}$ (s) $\Delta$ (s) $10^{-3}$ $10^{-6}$ $10^{-9}$
-Bitcoin (guo2022bitcoin, ) $600$ 10 $\leq 25$ $\leq 51$ $\leq 78$
-Bitcoin new $600$ 10 $\leq 23$ $\leq 47$ $\leq 71$
-Litecoin (guo2022bitcoin, ) $150$ 10 $\leq 36$ $\leq 74$ $\leq 110$
-Litecoin new $150$ 10 $\leq 32$ $\leq 63$ $\leq 99$
-Dogecoin (guo2022bitcoin, ) $60$ 10 $\leq 85$ $\leq 171$ $\leq 258$
-Dogecoin new $60$ 10 $\leq 52$ $\leq 127$ $\leq 211$
-ETC (guo2022bitcoin, ) $13$ 2 $\leq 73$ $\leq 151$ $\leq 230$
-ETC new $13$ 2 $\leq 47$ $\leq 114$ $\leq 193$
-Table 4. Bounds of confirmation depth (latency) for different security levels with $25\%$ adversarial mining power.
-Table 4 presents the depth needed to guarantee different security levels (probability of target transaction’s safety being violated) including $10^{-3}$, $10^{-6}$, $10^{-9}$, for Bitcoin, Litecoin, Dogecoin, and ETC, respectively.
-The table shows that the new upper bound due to Theorem 5 is often substantially smaller than the previous upper bound from (guo2022bitcoin, ).
-6.2. Confirmation versus Fault Tolerance
+We note that for small confirmation depths, the probability of mining more adversarial blocks than pacers dominates the probability of balancing any height.
+As the confirmation depth increases, the latter probability begins to make a noticeable impact, hence the bump on the curves in Fig.
+6.
+$\lambda^{-1}$ (s) $\Delta$ (s) $10^{-3}$
+Bitcoin [7] $600$ 10 $\leq 25$
+Bitcoin new $600$ 10 $\leq 24$
+Litecoin [7] $150$ 10 $\leq 36$
+Litecoin new $150$ 10 $\leq 32$
+Dogecoin [7] $60$ 10 $\leq 85$
+Dogecoin new $60$ 10 $\leq 52$
+ETC [7] $13$ 2 $\leq 73$
+ETC new $13$ 2 $\leq 47$
+TABLE V: Bounds of confirmation depth (latency) to guarantee $99.9\%$ safety with $25\%$ adversarial mining power.
+Table V presents the depth needed to guarantee a $99.9\%$ for Bitcoin, Litecoin, Dogecoin, and ETC, respectively.
+The table shows that the new upper bound due to Theorem 5 is often substantially smaller than the previous upper bound from [7].
+VI-B Confirmation versus Fault Tolerance
 In this subsection, we plot the confirmation depth or time that guarantees a desired safety level as a function of the fraction of adversarial mining, while holding other parameters fixed.
-This also makes it easier to see the maximum fault tolerance different bounding methods can apply to.
-We assume the ETC parameters.
-Refer to captionRefer to caption
-Figure 7. Confirmation latency that guarantees 99.9% safety for ETC assuming different fractions of adversarial mining. Top: Confirmation in depth. Bottom: Confirmation in time (minutes).
+This also makes it easier to see the maximum fault tolerance different bounding methods can be applied to.
+We base our analysis on the parameters of Ethereum Classic (ETC).
+Figure 8: Confirmation latency that guarantees 99.9% safety for ETC assuming different fractions of adversarial mining. Top: Confirmation in depth. Bottom: Confirmation in time (minutes).
 Fig.
@@
 The top graph uses confirmation by depth (Theorems 25, and 6), whereas the bottom graph uses confirmation by time (Theorems 8 and 9).
-We include two points reported in (gazi2022practical, ), where the numerical evaluation becomes impractical for much higher adversarial mining rates.
-Also included in the graph are achievable infinite-latency fault tolerances established in (garay2015bitcoin, )(pass2017analysis, )(li2021close, ) and the ultimate tolerance established in (dembo2020everything, )(gazi2020tight, ), and also in this paper.
+We include two points reported in [5], where the numerical evaluation becomes impractical for much higher adversarial mining rates.
+Also included in the graph are achievable infinite-latency fault tolerances established in [6][7][8][11][13], and the ultimate tolerance established in [3][4], and also in this paper.
 The new upper and lower bounds are significantly closer than all previous bounds.
-Even our new closed-form upper bound is quite competitive.
-6.3. Throughput-Latency Trade-Off
+Even our new closed-form upper bound is highly competitive.
+VI-C Throughput-Latency Trade-Off
 In this subsection, we evaluate the best throughput we can guarantee for optimized parameters including block size, mining rate, and confirmation depth.
@@
 We assume the block propagation delay and the block size satisfy the affine relationship (22).
-Using empirically determined delay bounds of 10 seconds for 1.6 MB Bitcoin blocks and 2 seconds for 200 KB Bitcoin Cash blocks, [^3] The 90th percentile of Bitcoin Cash propagation time is 1 second (fechner2022calibrating, ).
+Using empirically determined delay bounds of 10 seconds for 1.6 MB Bitcoin blocks and 2 seconds for 200 KB Bitcoin Cash blocks, [^3] The 90th percentile of Bitcoin Cash propagation time has been measured as 1 second [12].
 we infer that $r=178.6$ KB/second and $\nu=0.9$ (seconds).
-Figure 8. Achievable and unachievable throughput-latency trade-off.
+Figure 9: Achievable and unachievable throughput-latency trade-off and 5 existing currencies’ throughput-latency.
 For every given expected confirmation latency $d$, we can solve the optimization problem (24) with $q=0.001$ and given natural fork rates $\lambda\Delta=3/4$, $7/12$, $5/12$, and $1/4$, respectively.
@@
 We observe that the choice of $\lambda\Delta=7/12$ yields a large fraction of the maximum possible throughput for practical latencies of $30$ minutes or more.
-6.4. Guidance to Selecting Parameters
+VI-D Guidance for Selecting Parameters
 Bitcoin’s block size ($1.6$ MB) and mining rate ($1$ every $10$ minutes) have proven to be safe choices, with a maximum throughput of about $2.8$ KB/s.
@@
 As another important factor, increasing the mining rate increases natural forks, but also reduces the expected time to the same confirmation depth.
-The theorems in this paper provide the most accurate guidance to date to selecting parameters for desired system-level performance.
+The theorems in this paper provide the most accurate guidance to date for selecting parameters for desired system-level performance.
 For example, if the block size is increased to $2$ MB, the propagation upper bound will be around $\Delta=12.5$ seconds according to (22).
@@
 In fact, $44.44\%$ adversarial mining is practically not tolerable using the original Bitcoin parameters due to the excessively long delays of thousands of block confirmations.
-7. Conclusion
+VII Conclusion
 In this paper, we have developed a concise mathematical model for blockchain systems based on the Nakamoto consensus.
-We have obtained a new set of closed-form latency-security bounds, which are the first such bounds to apply to all parameter values.
+We have obtained a new set of closed-form latency-security bounds, which are the first such bounds to achieve the ultimate fault tolerance.
 These bounds are not only easily computed but also demonstrate superior numerical performance across a broad spectrum of parameter configurations.
 Our analysis not only provides an intuitive proof of the ultimate fault tolerance but also offers new insights for the practical application the Nakamoto consensus.
-Existing safety guarantees for longest-chain-fork-choice protocols have primarily been established through “block counting” methods, which calculates, the probability that one species of blocks ismined more than another species during some periods longer than the confirmation latency.
+Existing safety guarantees for longest-chain fork-choice protocols have primarily been established through “block counting” methods, which calculates, the probability that one species of blocks ismined more than another species during some periods longer than the confirmation latency.
 Such analyses essentially empower the adversary to use non-causal knowledge of future blocks (up until confirmation) and advance them to build a winning chain.
 In contrast, the production of a balanced height depends on blocks mined within an opportunity window of $\Delta$ seconds; future blocks cannot be used to create balancing candidates or balanced heights.
-This requirement eliminates certain non-causal attack strategies, resulting in improvements and the precise determination of fault tolerance.
+This requirement eliminates certain non-causal attack strategies, resulting in improvements and achieving the ultimate fault tolerance.
 In today’s Bitcoin and future open blockchain networks, miners play a crucial role in investing in infrastructure to mitigate block propagation delays.
@@
 The theorems presented in this paper serve as useful tools, empowering all stakeholders to make informed economic decisions regarding their operations.
+Acknowledgement
+D.
+Guo thanks Dr.
+David Tse for stimulating discussions that sparked the conceptualization of some initial ideas in the paper.
+This work was supported in part by the National Science Foundation (grant No.
+2132700).
 Appendix A Proof of Theorem 1
-Figure 9. Illustration of mining times of pacers and A-blocks after time $s$.
-Under a confirmation rule of latency $t$, assume a target transaction is public by time $s$.
+Figure 10: Illustration of mining times of pacers and A-blocks after time $s$.
+Under a confirmation rule of latency $t$, a target block is mined by time $s$.
 Within this proof, we count block arrivals from time $s$, i.e., the first A-block is the first A-block mined after $s$, and the first pacer is the first pacer mined after $s$.
@@
 An illustration of these random variables is given in Fig.
-9.
+10.
 Let $E_{j}$ denote the event that the $j$-th A-block arrives before the $j$-th pacer but after time $s+t$. The probability of safety violation is at least
-(45) $P(E_{j})=P(\sum_{i=0}^{j}X_{i}<\sum_{i=0}^{j}Y_{i},\sum_{i=0}^{j}X_{i}>t)$
-(46) $\geq 1-P(\sum_{i=0}^{j}(X_{i}-Y_{i})\geq 0)-P(\sum_{i=0}^{j}X_{i}\leq t)$
-$\geq 1-P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)-P(X_{0}-Y_{0}\geq n)$
-(47) $\quad-P(\sum_{i=0}^{j}X_{i}\leq t)$
+$P(E_{j})$ $=P\Bigg{(}\sum_{i=0}^{j}X_{i}<\sum_{i=0}^{j}Y_{i},\sum_{i=0}^{j}X_{i}>t\Bigg{)}$ (56)
+$\geq 1-P\Bigg{(}\sum_{i=0}^{j}(X_{i}-Y_{i})\geq 0\Bigg{)}-P\Bigg{(}\sum_{i=0}^{j}X_{i}\leq t\Bigg{)}$ (57)
+$\geq 1-P\Bigg{(}\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n\Bigg{)}-P(X_{0}-Y_{0}\geq n)-P\Bigg{(}\sum_{i=0}^{j}X_{i}\leq t\Bigg{)}$ (58)
 where (58) holds for every integer $n$.
@@
 Thus, by the law of large numbers,
-(48) $P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)<\frac{\epsilon}{3}.$
+$P\Bigg{(}\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n\Bigg{)}<\frac{\epsilon}{3}.$ (59)
 We conclude that $P(E_{j})\rightarrow 1$ as $j\rightarrow\infty$.
 Appendix B Proof of Theorem 11
-B.1. Critical Events
-Figure 10. Illustration of all possible state transitions over time.
-As a convention, we refer to the genesis block as the 0-th balancing candidate. We describe several types of events as follows: For every $i\in\{0,1,\cdots\}$, let
+B-A Critical Events
+Figure 11: Illustration of all possible state transitions over time.
+Let a target block be given. As a convention, we refer to the genesis block as the 0-th balancing candidate. We describe several types of events as follows: For every $i\in\{0,1,\cdots\}$, let
   • $C_{i}$ denotes the event that $i$ or more candidates are eventually mined;
@@
 For convenience, we define
-(49) $A_{i}=C_{i}\cap\bar{C}_{i+1}$
+$A_{i}=C_{i}\cap\bar{C}_{i+1}$ (60)
 which denotes the event that the $i$-th candidate is the last candidate ever mined.
 As time progresses, the events are seen to occur in natural order as illustrated in Fig.
-10.
+11.
 The process begins from $U_{0}$ as the genesis block, which is also the 0-th candidate, cannot be balanced.
@@
 Hence
-(50) $P(B_{k}|C_{k})\leq\epsilon$
+$P(B_{k}|C_{k})\leq\epsilon$ (61)
 where $\epsilon$ is given by (13).
@@
 This implies that the adversary must extend a shorter chain to catch up with the highest public honest block in order for the next candidate to be mined.
-In Appendix B.2, we show that the chance that no additional candidate will ever be mined is lower bounded as
-(51) $P(A_{k}|U_{k})\geq\delta$
+In Appendix B-B, we show that the chance that no additional candidate will ever be mined is lower bounded as
+$P(A_{k}|U_{k})\geq\delta$ (62)
 where $\delta$ is given by (14). The lower bound (62) is positive only when (1) holds.
-B.2. Proof of (51)
+B-B Proof of (62)
 Here we wish to lower bound $P(A_{k}|U_{k})$, which is the probability that no more candidate is ever mined after that the $k$-th candidate is not balanced by the end of its candidacy.
@@
 Conditioned on the first event, the second event’s probability is thus lower bounded as
-(52) $P\left(\cap_{t\geq 0}\{A_{r,r+t}\leq P_{\Delta,t}\}\right)=P\left(\cap_{t\geq 0}\{A_{0,t}\leq P_{0,t}\}\right)$
-(53) $=P\left(\max_{t\geq 0}\{A_{0,t}-P_{0,t}\}=0\right)$
-(54) $=P(M=0).$
+$P\left(\cap_{t\geq 0}\{A_{r,r+t}\leq P_{\Delta,t}\}\right)=P\left(\cap_{t\geq 0}\{A_{0,t}\leq P_{0,t}\}\right)$ (63)
+$=P\left(\max_{t\geq 0}\{A_{0,t}-P_{0,t}\}=0\right)$ (64)
+$=P(M=0).$ (65)
 Based on the preceding bounds and the pmf of $M$ given by (12), we have
-(55) $P(A_{k}|U_{k})$ $\geq e^{-a\Delta}P(M=0)$
-(56) $=\delta.$
-B.3. Partially Markov Processes
+$P(A_{k}|U_{k})$ $\geq e^{-a\Delta}P(M=0)$ (66)
+$=\delta.$ (67)
+B-C Partially Markov Processes
 Consider a sample space in which each outcome corresponds to a (sample) path on the directional graph depicted by Fig.
@@
 Let $P^{0}(\cdot)$ stand for the special pmf under which all state transitions on the directed graph are Markov with the following transition probabilities: For every $k\in\{0,1,...\}$,
-(57) $P^{0}(B_{k}|C_{k})$ $=1-P^{0}(U_{k}|C_{k})=\epsilon,$
-(58) $P^{0}(A_{k}|U_{k})$ $=1-P^{0}(C_{k+1}|U_{k})=\delta,$
+$P^{0}(B_{k}|C_{k})$ $=1-P^{0}(U_{k}|C_{k})=\epsilon,$ (68)
+$P^{0}(A_{k}|U_{k})$ $=1-P^{0}(C_{k+1}|U_{k})=\delta,$ (69)
 where $\epsilon$ and $\delta$ are given by (13) and (14), respectively. These transition probabilities satisfy the prior constraints (61) and (62) with equality under condition (1).
 We next prove that every pmf $P$ which satisfies (61) and (62) must also satisfy
-(59) $P(X\leq n)\geq P^{0}(X\leq n).$
+$P(X\leq n)\geq P^{0}(X\leq n).$ (70)
 Towards this end, we define a sequence of pmfs, $P^{1},P^{2},\dots$, as follows: Let $0\,b_{1}\dots b_{i}*$ denote the event consisting of all sample paths with the prefix $0\,b_{1}\dots b_{i}$.
 Let the probability of a sample path under $P^{m}$ be defined as
-(60) $P^{m}(\{0\,b_{1}\dots b_{k-1}0\})=P(\{0\,b_{1}\dots b_{k-1}0\})$
-for $k=1,\dots,m$ and as
+$P^{m}(\{0\,b_{1}\dots b_{k-1}0\})=P(\{0\,b_{1}\dots b_{k-1}0\})$ (71)
+for $k=1,\dots,m-1$ and as
 […]
-for $k=m+1,m+2,\dots$.
-The outcome $00\,b_{1}\dots b_{k-1}$ 0 with $k>m$ implies that $C_{m}$ occurs.
+for $k=m,m+1,\dots$.
+The outcome $00\,b_{1}\dots b_{k-1}$ 0 with $k\geq m$ implies that $C_{m}$ occurs.
 It is straightforward to verify that (71), (72) defines a pmf.
@@
 In particular, conditioned on $C_{m}$, the balanced stages before stage $m$ and the balanced stages afterward are independent.
-For every pair of natural numbers $k$ and $j$, we define the following pair of random variables as functions of the sample path $00\,b_{1}\dots b_{k-1}$ 0:
-(62) $L_{j}$ $=b_{1}+\dots+b_{\min(k,j)-1}$
-(63) $R_{j}$ $=b_{j}+\dots+b_{k-1},$
+For every pair of natural numbers $l$ and $j$, we define the following pair of random variables as functions of the sample path $00\,b_{1}\dots b_{l-1}$ 0:
+$L_{j}$ $=b_{1}+\dots+b_{\min(l,j)-1}$ (73)
+$R_{j}$ $=b_{j}+\dots+b_{l-1},$ (74)
 with the convention that $L_{0}=0$. Evidently, with $01_{B_{0}}=$ 0
-(64) $X=L_{m}+1_{B_{m}}+R_{m+1}$
+$X=L_{m}+1_{B_{m}}+R_{m+1}$ (75)
 holds for every natural number $m$.
-In Appendices B.4 and B.5, we show that for every $m=0,1,\dots$ and every $n=0,1,\dots$,
-(65) $P^{m+1}(X\leq n)\geq P^{m}(X\leq n)$
+In Appendices B-D and B-E, we show that for every $m=0,1,\dots$ and every $n=0,1,\dots$,
+$P^{m+1}(X\leq n)\geq P^{m}(X\leq n)$ (76)
 and
-(66) $\lim_{m\to\infty}P^{m}(X\leq n)=P(X\leq n)$
+$\lim_{m\to\infty}P^{m}(X\leq n)=P(X\leq n)$ (77)
 hold. As the limit of an increasing sequence, $P(X\geq n)$ must satisfy (70).
 For $n=0,1,\dots$, let us denote
-(67) $f_{n}$ $=P^{0}(X\geq n|C_{1})$
-(68) $=P^{0}(R_{1}\geq n|C_{1}).$
+$f_{n}$ $=P^{0}(X\geq n|C_{1})$ (78)
+$=P^{0}(R_{1}\geq n|C_{1}).$ (79)
 Then we have the following recursion
-(69) $f_{n}=$ $P^{0}(R_{1}\geq n,U_{1}|C_{1})+P^{0}(R_{1}\geq n,B_{1}|C_{1})$
+$f_{n}=$ $P^{0}(R_{1}\geq n,U_{1}|C_{1})+P^{0}(R_{1}\geq n,B_{1}|C_{1})$ (80)
 $=$ $P^{0}(R_{2}\geq n|C_{2})P^{0}(C_{2}|U_{1})P^{0}(U_{1}|C_{1})$
-(70) $+P^{0}(R_{2}\geq n-1|B_{1})P^{0}(B_{1}|C_{1})$
-(71) $=$ $(1-\delta)(1-\epsilon)f_{n}+\epsilon f_{n-1}.$
+$+P^{0}(R_{2}\geq n-1|B_{1})P^{0}(B_{1}|C_{1})$ (81)
+$=$ $(1-\delta)(1-\epsilon)f_{n}+\epsilon f_{n-1}.$ (82)
 Solving the recursion with the fact that $f_{0}=1$ yields
-(72) $f_{n}=\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}.$
+$f_{n}=\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}.$ (83)
 It is easy to verify that
-(73) $P^{0}(X\leq n)$ $=1-P^{0}(R_{1}\geq n+1|C_{1})P(C_{1})$
-(74) $=F(n)$
+$P^{0}(X\leq n)$ $=1-P^{0}(R_{1}\geq n+1|C_{1})P^{0}(C_{1})$ (84)
+$=F(n)$ (85)
 where $F(\cdot)$ is defined in (16).
@@
 In the remainder of this section, we are concerned with the number of balanced candidates up to the $k$-th candidacy, denoted as $X^{(k)}$.
-Conditioning on whether $C_{k+1}$ is visited, we can lower bound the cdf of $X^{(k)}$:
-$P(X^{(k)}\leq n)$
-(75) $\geq P^{0}(X^{(k)}\leq n)$
-(76) $=P^{0}(X^{(k)}\leq n,X\geq n+1)+P^{0}(X\leq n)$
-(77) $=P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})+P^{0}(X\leq n)$
-where (75) can be easily seen from (65), and (77) holds because $X^{(k)}\leq n$ and $X\geq n+1$ implies the occurrence of $C_{k+1}$.
+We first establish the following result:
+$P(X^{(k)}\leq n)\geq P^{0}(X^{(k)}\leq n)$ (86)
+where (86) can be proved as follows. Since we only consider the first $k$ candidates, which means only the first $k$ stages are relevant, we have
+$P(X^{(k)}\leq n)=P^{k+1}(X^{(k)}\leq n).$
+Furthermore, we can prove that for every $m=0,1,\dots,k$ and $n=0,1,\dots,k$,
+$P^{m+1}(X^{(k)}\leq n)\geq P^{m}(X^{(k)}\leq n).$ (87)
+For $X^{(k)}$, we define $L^{\prime}_{j}$ and $R^{\prime}_{j}$ as functions of the sample path $00\,b_{1}\dots b_{l-1}$ 0 for natural number $l$:
+$L^{\prime}_{j}$ $=b_{1}+\dots+b_{\min(j,l)-1},\quad j=0,1,\dots,k,$ (88)
+$R^{\prime}_{j}$ $=b_{j}+\dots+b_{\min\{k,l\}},\quad j=1,2,\dots,k+1,$ (89)
+with boundary conditions $L^{\prime}_{0}=0$ and $R^{\prime}_{k+1}=0$. Thus, we can express $X^{(k)}$ as:
+$X^{(k)}=L^{\prime}_{m}+1_{B_{m}}+R^{\prime}_{m+1},\quad\text{for}m\leq k.$ (90)
+The remainder of the proof follows the same steps as in the proof of (76) in Appendix B-D, the details of which is omitted here.
+Using (86), we can lower bound the cdf of $X^{(k)}$:
+$P(X^{(k)}\leq n)$ $=P^{0}(X^{(k)}\leq n,X\geq n+1)+P^{0}(X\leq n)$ (91)
+$=P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})+P^{0}(X\leq n)$ (92)
+where (92) holds because $X^{(k)}\leq n$ and $X\geq n+1$ implies the occurrence of $C_{k+1}$.
 Since we already know $P^{0}(X\leq n)$ by (85), what left is to calculate $P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$.
 By law of total probability, we have
-(78) $\begin{split}&P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})\\ &=\sum_{i=0}^{n}P^{0}(X^{(k)}=i,C_{k+1})P^{0}(X-X^{(k)}\geq n+1-i|C_{k+1}).\end{split}$
+$\begin{split}P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})=\sum_{i=0}^{n}P^{0}(X^{(k)}=i,C_{k+1})P^{0}(X-X^{(k)}\geq n+1-i|C_{k+1}).\end{split}$ (93)
 Due to Markov property on $P^{0}$, $X-X^{(k)}\geq n+1-i$ conditioning on $C_{k+1}$ is identically distributed as $X\geq n+1-i$ conditioning on $C_{1}$.
@@
 $P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$
-(79) $=\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}((1-\epsilon)(1-\delta))^{k-i}(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-i+1}$
-(80) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}(\frac{\epsilon(1-\epsilon)(1-\delta)}{\epsilon+\delta-\epsilon\delta})$
-$=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{k}$
-(81) $\qquad\sum_{i=0}^{n}\binom{k}{i}(\epsilon+\delta-\epsilon\delta)^{i}(1-\epsilon)(1-\delta)^{k-i}$
-(82) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n+1}F_{b}(n;k,(\epsilon+\delta-\epsilon\delta)),$
+$=\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}((1-\epsilon)(1-\delta))^{k-i}(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-i+1}$ (94)
+$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-k+1}\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}\Big{(}\frac{\epsilon(1-\epsilon)(1-\delta)}{\epsilon+\delta-\epsilon\delta}\Big{)}^{k-i}$ (95)
+$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-k+1}\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{k}\sum_{i=0}^{n}\binom{k}{i}(\epsilon+\delta-\epsilon\delta)^{i}((1-\epsilon)(1-\delta))^{k-i}$ (96)
+$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n+1}F_{b}(n;k,(\epsilon+\delta-\epsilon\delta)),$ (97)
 where $F_{b}(\cdot;k,p)$ denotes the cdf of a binomial distribution with $k$ independent trials and success probability $p$.
 Plugging (97) back into (93) completes the proof of (15) and also Theorem 11.
-B.4. Proof of (65)
+B-D Proof of (76)
 We use $C_{m,i}$ to denote the intersection of events $C_{m}$ and $L_{m}=i$. Then
-$P^{m+1}($ $R_{m+1}<c\,|\,C_{m+1,i},B_{m})$
-(83) $=P^{0}(R_{m+1}<c\,|\,C_{m+1})$
-(84) $=P^{m+1}(R_{m+1}<c\,|\,C_{m+1,i},U_{m}).$
+$P^{m+1}(R_{m+1}<c\,|\,C_{m+1,i},B_{m})$ $=P^{0}(R_{m+1}<c\,|\,C_{m+1})$ (98)
+$=P^{m+1}(R_{m+1}<c\,|\,C_{m+1,i},U_{m}).$ (99)
 Also, partially Markov process with measure $P^{m}(\cdot)$ becomes memoryless once event $C_{m}$ occurs:
-(85) $P^{m}(A_{m}|C_{m,i})$ $=P^{0}(A_{m}|C_{m,i})$
-(86) $P^{m}(B_{m}|C_{m,i})$ $=P^{0}(B_{m}|C_{m,i})$
-(87) $P^{m}(C_{m,i})$ $=P(C_{m,i})$
-(88) $=P^{m+1}(C_{m,i}).$
+$P^{m}(A_{m}|C_{m,i})$ $=P^{0}(A_{m}|C_{m,i})$ (100)
+$P^{m}(B_{m}|C_{m,i})$ $=P^{0}(B_{m}|C_{m,i})$ (101)
+$P^{m}(C_{m,i})$ $=P(C_{m,i})$ (102)
+$=P^{m+1}(C_{m,i}).$ (103)
 Consider the right-hand side of (76):
-(89) $P^{m}(X\leq a)=P^{m}(X\leq a,C_{m}^{c})+P^{m}(X\leq a,C_{m})$
-(90) $=P^{m}(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P^{m}(C_{m,i})P^{m}(X\leq a|C_{m,i})$
-(91) $=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m}(X\leq a|C_{m,i})$
+$P^{m}(X\leq a)=P^{m}(X\leq a,C_{m}^{c})+P^{m}(X\leq a,C_{m})$ (104)
+$=P^{m}(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P^{m}(C_{m,i})P^{m}(X\leq a|C_{m,i})$ (105)
+$=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m}(X\leq a|C_{m,i})$ (106)
 where (106) is due to the fact that $P^{m}$ is identical to $P$ up until stage $m-1$ as in (102). Similarly, by (103)
-(92) $\begin{split}&P^{m+1}(X\leq a)\\ &=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m+1}(X\leq a|C_{m,i}).\end{split}$
+$\begin{split}P^{m+1}(X\leq a)=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m+1}(X\leq a|C_{m,i}).\end{split}$ (107)
 In order to establish (76), it suffices to show
-(93) $P^{m+1}(X\leq a|C_{m,i})\geq P^{m}(X\leq a|C_{m,i})$
+$P^{m+1}(X\leq a|C_{m,i})\geq P^{m}(X\leq a|C_{m,i})$ (108)
 holds for all $i\in\{0,\dots,a\}$.
-Since $C_{m}=B_{m}\cup(U_{m},C_{m+1})\cup A_{m},$, we have
+Since $C_{m}=B_{m}\cup(U_{m},C_{m+1})\cup A_{m},$we have
 $P^{m+1}(X\leq a|C_{m,i})$
-$=$ $P^{m+1}(X\leq a,B_{m}|C_{m,i})+P^{m+1}(X\leq a,A_{m}|C_{m,i})$
-(94) $+P^{m+1}(X\leq a,U_{m},C_{m+1}|C_{m,i})$
-$=$ $P^{m+1}(R_{m}\leq a-i-1,B_{m}|C_{m,i})+P^{m+1}(A_{m}|C_{m,i})$
-(95) $+P^{m+1}(R_{m}\leq a-i,U_{m},C_{m+1}|C_{m,i})$
-$=$ $P^{m+1}(R_{m}\leq a-i-1|C_{m+1})P^{m+1}(B_{m},C_{m+1}|C_{m,i})$
-$\quad+P^{m+1}(R_{m}\leq a-i|C_{m+1})P^{m+1}(U_{m},C_{m+1}|C_{m,i})$
-(96) $\quad+P^{m+1}(A_{m}|C_{m,i})$
-$=$ $P^{m+1}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{m+1}(U_{m},C_{m+1}|C_{m,i})x_{2}$
-(97) $\quad+P^{m+1}(A_{m}|C_{m,i})$
+$=$ $P^{m+1}(X\leq a,B_{m}|C_{m,i})+P^{m+1}(X\leq a,A_{m}|C_{m,i})+P^{m+1}(X\leq a,U_{m},C_{m+1}|C_{m,i})$ (109)
+$=$ $P^{m+1}(R_{m+1}\leq a-i-1,B_{m}|C_{m,i})+P^{m+1}(A_{m}|C_{m,i})+P^{m+1}(R_{m+1}\leq a-i,U_{m},C_{m+1}|C_{m,i})$ (110)
+$=$ $P^{m+1}(R_{m+1}\leq a-i-1|C_{m+1})P^{m+1}(B_{m},C_{m+1}|C_{m,i})+P^{m+1}(R_{m+1}\leq a-i|C_{m+1})P^{m+1}(U_{m},C_{m+1}|C_{m,i})$
+$+P^{m+1}(A_{m}|C_{m,i})$ (111)
+$=$ $P^{m+1}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{m+1}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{m+1}(A_{m}|C_{m,i})$ (112)
 where
-(98) $x_{1}$ $=P^{0}(R_{m}\leq a-i-1|C_{m+1})$
-(99) $x_{2}$ $=P^{0}(R_{m}\leq a-i|C_{m+1}),$
+$x_{1}$ $=P^{0}(R_{m+1}\leq a-i-1|C_{m+1})$ (113)
+$x_{2}$ $=P^{0}(R_{m+1}\leq a-i|C_{m+1}),$ (114)
 and (110) is due to (75),(112) is due to (98). Similarly, with (100) and (101), we have
-(100) $\begin{split}P^{m}&(X\leq a|C_{m,i})=P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}\\ &+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i}).\end{split}$
+$\begin{split}P^{m}&(X\leq a|C_{m,i})=P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}\\ &+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i}).\end{split}$ (115)
 Due to (61) and (62),
-(101) $P^{0}(A_{m}|C_{m,i})$ $\leq P^{m+1}(A_{m}|C_{m,i})$
-(102) $P^{0}(B_{m},C_{m+1}|C_{m,i})$ $\geq P^{m+1}(B_{m},C_{m+1}|C_{m,i}).$
+$P^{0}(A_{m}|C_{m,i})$ $\leq P^{m+1}(A_{m}|C_{m,i})$ (116)
+$P^{0}(B_{m},C_{m+1}|C_{m,i})$ $\geq P^{m+1}(B_{m},C_{m+1}|C_{m,i}).$ (117)
 Now we introduce the following claim to assist the proof, which enables us to compare differently weighted sums of three numbers.
@@
 Then all real numbers $x_{1}\leq x_{2}\leq 1$ satisfy
-(103) $\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}\leq\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
+$\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}\leq\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$ (118)
 Proof.
 The proof is straightforward:
-$\lambda_{1}x_{1}$ $+\lambda_{2}x_{2}+\lambda_{3}$
-(104) $=x_{2}+\lambda_{1}(x_{1}-x_{2})+\lambda_{3}(1-x_{2})$
-(105) $\leq x_{2}+\lambda_{1}^{\prime}(x_{1}-x_{2})+\lambda_{3}^{\prime}(1-x_{2})$
-(106) $=\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
+$\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}$ $=x_{2}+\lambda_{1}(x_{1}-x_{2})+\lambda_{3}(1-x_{2})$ (119)
+$\leq x_{2}+\lambda_{1}^{\prime}(x_{1}-x_{2})+\lambda_{3}^{\prime}(1-x_{2})$ (120)
+$=\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$ (121)
 Lemma 14 implies
-$\quad P^{m+1}(X\leq a|C_{m,i})$
-$\geq P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+$
-(107) $\quad P^{0}(A_{m}|C_{m,i})$
-(108) $=P^{m}(X\leq a|C_{m,i}),$
+$P^{m+1}(X\leq a|C_{m,i})$ $\geq P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i})$ (122)
+$=P^{m}(X\leq a|C_{m,i}),$ (123)
 which completes the proof of (76).
-B.5. Proof of (66)
-By the definition of limit, proving (66) equals to prove $\forall\lambda\geq 0,a\in\{0,1,\dots\},\exists N_{0},s.t.\forall N>N_{0}$, $|P^{N}(X\leq a)-P(X\leq a)|<\lambda$.
+B-E Proof of (77)
+By the definition of limit, proving (77) equals to prove $\forall\lambda\geq 0,a\in\{0,1,\dots\},\exists N_{0}$ such that $\forall N>N_{0}$, $|P^{N}(X\leq a)-P(X\leq a)|<\lambda$.
 Conditioning on the occurrence of event $C_{N}$,
-$P^{N}(X\leq a)$ $=P^{N}(C_{N}^{c})P^{N}(X\leq a|C_{N}^{c})$
-(109) $\quad+P^{N}(C_{N})P^{N}(X\leq a|C_{N}),$
-(110) $P(X\leq a)$ $=P(C_{N}^{c})P(X\leq a|C_{N}^{c})+P(C_{N})P(X\leq a|C_{N}).$
-By definition of (60) and (61), $P(X\leq a|C_{N}^{c})=P^{N}(X\leq a|C_{N}^{c})$ and $P(C_{N}^{c})=P^{N}(C_{N}^{c})$. Let $p_{1}=P^{N}(X\leq a|C_{N})$ and $p_{2}=P(X\leq a|C_{N}))$. We have,
-$|P^{N}$ $(X\leq a)-P(X\leq a)|$
-(111) $=P(C_{N})|p_{1}-p_{2}|$
-(112) $\leq(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1}|p_{1}-p_{2}|$
-(113) $<\lambda,$
+$P^{N}(X\leq a)$ $=P^{N}(C_{N}^{c})P^{N}(X\leq a|C_{N}^{c})+P^{N}(C_{N})P^{N}(X\leq a|C_{N}),$ (124)
+$P(X\leq a)$ $=P(C_{N}^{c})P(X\leq a|C_{N}^{c})+P(C_{N})P(X\leq a|C_{N}).$ (125)
+By definition of (71) and (72), $P(X\leq a|C_{N}^{c})=P^{N}(X\leq a|C_{N}^{c})$ and $P(C_{N}^{c})=P^{N}(C_{N}^{c})$. Let $p_{1}=P^{N}(X\leq a|C_{N})$ and $p_{2}=P(X\leq a|C_{N})$. We have,
+$|P^{N}(X\leq a)-P(X\leq a)|$ $=P(C_{N})\,|p_{1}-p_{2}|$ (126)
+$\leq(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1}|p_{1}-p_{2}|$ (127)
+$<\lambda,$ (128)
 where (127) is due to the following
-(114) $P(C_{N})\leq P^{N-1}(C_{N})\leq P^{N-2}(C_{N})\leq\cdots$ $\leq P^{0}(C_{N}),$
+$P(C_{N})\leq P^{N-1}(C_{N})\leq P^{N-2}(C_{N})\leq\cdots$ $\leq P^{0}(C_{N}),$ (129)
 (129) can be derived by definition of (71) and (72) as
-(115) $P^{i}(C_{N})$ $=P^{0}(C_{N}|C_{i+1})P^{i}(C_{i+1}|C_{i})P^{i}(C_{i})$
-(116) $\leq P^{0}(C_{N}|C_{i+1})P^{0}(C_{i+1}|C_{i})P^{i-1}(C_{i}).$
+$P^{i}(C_{N})$ $=P^{0}(C_{N}|C_{i+1})P^{i}(C_{i+1}|C_{i})P^{i}(C_{i})$ (130)
+$\leq P^{0}(C_{N}|C_{i+1})P^{0}(C_{i+1}|C_{i})P^{i-1}(C_{i}).$ (131)
 Then, by the memoryless property of the Markov process with measure $P^{0}(\cdot)$we have
-(117) $P^{0}(C_{N})$ $=P^{0}(C_{1}|U_{0})\prod_{i=1}^{N-1}P^{0}(C_{i+1}|C_{i})$
-(118) $=(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1},$
+$P^{0}(C_{N})$ $=P^{0}(C_{1}|U_{0})\prod_{i=1}^{N-1}P^{0}(C_{i+1}|C_{i})$ (132)
+$=(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1},$ (133)
 where for $i=\{1,2,\cdots,N-1\}$
-(119) $P^{0}(C_{i+1}|C_{i})$ $=1-P^{0}(A_{i}|C_{i})$
-(120) $=1-\delta(1-\epsilon).$
+$P^{0}(C_{i+1}|C_{i})$ $=1-P^{0}(A_{i}|C_{i})$ (134)
+$=1-\delta(1-\epsilon).$ (135)
 Thus $\lim\limits_{N\rightarrow\infty}P^{N}(X\leq a)=P(X\leq a)$. Hence (77) holds.
@@
 For every integer $n$, we have
-(121) $P(X+Y\geq 0)$ $\leq P(X>-n\text{ or }Y>n-1)$
-(122) $\leq P(X>-n)+P(Y>n-1)).$
+$P(X+Y\geq 0)$ $\leq P(X>-n\text{ or }Y>n-1)$ (136)
+$\leq P(X>-n)+P(Y>n-1).$ (137)
 Taking the infimum over $n$ on both sides yields the upper bound in Lemma 12.
@@
 Then for all $k\in\mathbb{Z}$, we have
-(123) $\bar{F}_{X}(-a+k)+\bar{F}_{Y}(a-k-1)\geq\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$
+$\bar{F}_{X}(-a+k)+\bar{F}_{Y}(a-k-1)\geq\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$ (138)
 Or equivalently, for $k\in\mathbb{N}$,
-(124) $P(-a<X\leq-a+k)\leq P(a-k-1<Y\leq a-1),$
-(125) $P(a-1<Y\leq a+k-1)\leq P(-a-k<X\leq-a).$
+$P(-a<X\leq-a+k)\leq P(a-k-1<Y\leq a-1),$ (139)
+$P(a-1<Y\leq a+k-1)\leq P(-a-k<X\leq-a).$ (140)
 To attain the bound, we must construct $X_{0}$ and $Y_{0}$ such that $X_{0}+Y_{0}\geq 0$ whenever $X_{0}>-a$ or $Y_{0}>a-1$, and we need to ensure the two cases are disjoint.
 Set a random variable $U$, which has a uniform distribution over $[0,1]$. Given $F$, define
-(126) $F^{-1}(x)=\sup\{a\in\mathbb{R}|F(a)<x\}.$
+$F^{-1}(x)=\sup\{a\in\mathbb{R}|F(a)<x\}.$ (141)
 We define $X_{0}$ and $Y_{0}$ to be the following:
-(127) $X_{0}=F_{X}^{-1}(U+F_{X}(-a)-\lfloor U+F_{X}(-a)\rfloor)$
-(128) $Y_{0}=F_{Y}^{-1}(F_{Y}(a-1)-U-\lfloor F_{Y}(a-1)-U\rfloor)$
+$X_{0}=F_{X}^{-1}(U+F_{X}(-a)-\lfloor U+F_{X}(-a)\rfloor)$ (142)
+$Y_{0}=F_{Y}^{-1}(F_{Y}(a-1)-U-\lfloor F_{Y}(a-1)-U\rfloor)$ (143)
 $X_{0}$ and $Y_{0}$ have cdf $F_{X}$ and $F_{Y}$ since they are rearrangements of $F^{-1}_{X}$ and $F^{-1}_{Y}$.
 Now, we show that $X_{0}(u)+Y_{0}(u)\geq 0$ for $u\leq 1-F_{X}(-a)$ $a.e.$ The same holds for $u>F_{Y}(a-1)$, and we omit this similar proof here.
-(129) $F^{-1}_{X}(x)=\sup\{a\in\mathbb{R}|F_{X}(a)<x\}.$
+$F^{-1}_{X}(x)=\sup\{a\in\mathbb{R}|F_{X}(a)<x\}.$ (144)
 implies that
-(130) $F_{x}\circ F_{x}^{-1}(x)>x\quad a.e.,$
+$F_{x}\circ F_{x}^{-1}(x)>x\quad a.e.,$ (145)
 since $X$ is integer-valued. Consider
-(131) $S_{b}=\left\{u\leq 1-F_{X}(-a)|X_{0}(u)\leq b,Y_{0}(u)\leq-b-1\right\}$
+$S_{b}=\left\{u\leq 1-F_{X}(-a)|X_{0}(u)\leq b,Y_{0}(u)\leq-b-1\right\}$ (146)
 And,
-(132) $F_{X}(b)$ $\geq F_{X}\circ X_{0}(u)$
-(133) $=F_{X}\circ F^{-1}_{X}(u+F_{X}(-a))$
-(134) $>u+F_{X}(-a)\quad a.e.,$
+$F_{X}(b)$ $\geq F_{X}\circ X_{0}(u)$ (147)
+$=F_{X}\circ F^{-1}_{X}(u+F_{X}(-a))$ (148)
+$>u+F_{X}(-a)\quad a.e.,$ (149)
 Accordingly,
-(135) $F_{Y}(-b-1)$ $\geq F_{Y}\circ Y_{0}(u)$
-(136) $=F_{Y}\circ F^{-1}_{Y}(F_{Y}(a-1)-u)$
-(137) $>F_{Y}(a-1)-u\quad a.e.$
+$F_{Y}(-b-1)$ $\geq F_{Y}\circ Y_{0}(u)$ (150)
+$=F_{Y}\circ F^{-1}_{Y}(F_{Y}(a-1)-u)$ (151)
+$>F_{Y}(a-1)-u\quad a.e.$ (152)
 Adding up (149) and (152) yields
-(138) $F_{X}(b)+F_{Y}(-b-1)>F_{X}(-a)+F_{Y}(a-1)\quad a.e.$
+$F_{X}(b)+F_{Y}(-b-1)>F_{X}(-a)+F_{Y}(a-1)\quad a.e.$ (153)
 However, $\phi(a)$ is maximal, leading to $P(S_{b})=0$
-(139) $P(u\leq 1-F_{X}(-a)|X_{0}(u)+Y_{0}(u)<0)=0$
+$P(u\leq 1-F_{X}(-a)|X_{0}(u)+Y_{0}(u)<0)=0$ (154)
 Finally,
 $P(X_{0}+Y_{0})$ $\geq 1-F_{X}(-a)+1-F_{Y}(a-1)$
-(140) $\quad=\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$
+$\quad=\bar{F}_{X}(-a)+\bar{F}_{Y}(a-1)$ (155)
 Thus $X_{0}$ and $Y_{0}$ attain the optimal bound, which concludes the lemma.
@@
 This appendix finished the proof of Theorem 2 in Sec.
-5.3.
+V-C.
 We utilize Theorem 11, proved in Appendix B, to derive a closed-form upper bound on the probability of a safety violation for confirmation by depth $k$.
@@
 The second bound has a steeper exponent whose expression is more complicated.
-D.1. Upper Bounding $Y$
+D-A Upper Bounding $Y$
 Recall that $Y^{(k)}$ is defined in (48).
@@
 $P(Y^{(k)}>-n)$
-(141) $\leq P\!\big{(}L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}$
-(142) $\leq P\!\big{(}L_{s}-k+A_{s,\tau+2\Delta}+\sup_{c\geq 0}\{A_{\tau+2\Delta,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}.$
+$\leq P\!\big{(}L_{s}-k+\sup_{c\geq 0}\{A_{s,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}$ (156)
+$\leq P\!\big{(}L_{s}-k+A_{s,\tau+2\Delta}+\sup_{c\geq 0}\{A_{\tau+2\Delta,\tau+2\Delta+c}-P^{\prime}_{0,c}\}>-n\big{)}.$ (157)
 The lead $L_{s}$ depends only on blocks mined by $s$.
@@
 Let us define
-(143) $M=\sup_{t\geq 0}\{A^{\prime}_{0,t}-P^{\prime}_{0,t}\}$
+$M=\sup_{t\geq 0}\{A^{\prime}_{0,t}-P^{\prime}_{0,t}\}$ (158)
 where $A^{\prime}$ is a Poisson process identically distributed as $A$.
@@
 We can thus rewrite (157) as
-(144) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,\tau+2\Delta}+M>-n\right).$
-Let the first H-block mined after $s+3\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on.
-Let us also define $s+T+2\Delta$ to be the first time $k$ new pacers are mined.
-Since the height of the $k$-th new pacer is at least $h_{b}+k$, we have $s+T+2\Delta\geq\tau$ by $\tau$’s definition.
+$P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,\tau+2\Delta}+M>-n\right).$ (159)
+Let the first H-block mined after $s$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on.
+Let us also define $s+T-\Delta$ to be the first time $k$ new pacers are mined.
+Since the height of the $k$-th new pacer is at least $h_{b}+k$, we have $s+T-\Delta\geq\tau$ by $\tau$’s definition.
 We then relax (159) to
-(145) $P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,s+T+4\Delta}+M>-n\right).$
+$P(Y^{(k)}>-n)$ $\leq P\left(L_{s}-k+A_{s,s+T+\Delta}+M>-n\right).$ (160)
 While $\tau$ is generally dependent on $L_{s}$, we have $T$ independent of $L_{s}$ due to the memoryless nature of mining.
 […]
-(146) $P(Y^{(k)}>-n)$ $\leq P\left(L+A_{0,2\Delta}-k+A_{2\Delta,T+6\Delta}+M>-n\right)$
-(147) $=P\left(L+M+A_{0,T+6\Delta}-k+n>0\right).$
-In (147), $L$, $M$, and $A_{0,T+4\Delta}$ are mutually independent.
+$P(Y^{(k)}>-n)$ $\leq P\left(L+A_{0,2\Delta}-k+A_{2\Delta,T+3\Delta}+M>-n\right)$ (161)
+$=P\left(L+M+A_{0,T+3\Delta}-k+n>0\right).$ (162)
+In (162), $L$, $M$, and $A_{0,T+3\Delta}$ are mutually independent.
 We shall derive their MGFs shortly and use them to upper bound the probability on the right hand side of (162).
 We take two separate approaches: In Appendix F, we let $n=ck$ with a specific choice of $c>0$ and establish Corollary 3 by showing that as long as (1) holds, (50) and (162) both vanish as $k\to\infty$.
-D.2. Chernoff Bound
+D-B Chernoff Bound
 In this appendix, obtain an explicit exponential bound and thus establish Theorem 2.
@@
 First, by Theorem 11,
-(148) $P(X^{(k)}\geq n)$ $\leq P(X\geq n)$
-(149) $\leq(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}$
+$P(X^{(k)}\geq n)$ $\leq P(X\geq n)$ (163)
+$\leq(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}$ (164)
 where $\epsilon$ and $\delta$ are defined in (13) and (14), respectively. Note that this even holds for non-integer $n$.
@@
 Using the same computation in Appendix F, we have
-(150) $P(Y^{(k)}>-n)$ $\leq P(M+L+A_{0,T+6\Delta}+n-k>0)$
-(151) $\leq e^{(e^{\nu}-1)6\bar{a}}\left(q(\nu)\right)^{k}e^{\nu n}E[e^{\nu M}]E[e^{\nu L}]$
+$P(Y^{(k)}>-n)$ $\leq P(M+L+A_{0,T+3\Delta}+n-k>0)$ (165)
+$\leq e^{(e^{\nu}-1)3\bar{a}}\left(q(\nu)\right)^{k}e^{\nu n}E[e^{\nu M}]E[e^{\nu L}]$ (166)
 where
-(152) $q(\nu)=\frac{e^{(e^{\nu}-1)\bar{a}-\nu}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}.$
+$q(\nu)=\frac{e^{(e^{\nu}-1)\bar{a}-\nu}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}.$ (167)
 We select $\nu$ to minimize $q(\nu)$, which is not difficult by setting the derivative $q^{\prime}(\nu)=0$. The unique valid solution $\nu^{*}$ satisfies
-(153) $e^{\nu^{*}}=1+\gamma(\bar{a},\bar{h})/\bar{a}$
+$e^{\nu^{*}}=1+\gamma(\bar{a},\bar{h})/\bar{a}$ (168)
 where
-(154) $\gamma(\alpha,\eta)=\frac{1}{2}\left(2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}\right).$
+$\gamma(\alpha,\eta)=\frac{1}{2}\left(2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}\right).$ (169)
 Letting $n=\zeta k$ with $\zeta$ to be decided later, we have
-(155) $P(X^{(k)}\geq\zeta k)\leq(1-\delta)e^{\zeta k\,\ln\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)}.$
+$P(X^{(k)}\geq\zeta k)\leq(1-\delta)e^{\zeta k\,\ln\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)}.$ (170)
 Using lemma 4, we evaluate the MGFs of $M$ and $L$ as
-(156) $E[e^{\nu^{*}L}]$ $=E[e^{\nu^{*}M}]$
-(157) $=\mathcal{E}\left(1+\frac{\gamma(\bar{a},\bar{h})}{\bar{a}}\right).$
+$E[e^{\nu^{*}L}]$ $=E[e^{\nu^{*}M}]$ (171)
+$=\mathcal{E}\left(1+\frac{\gamma(\bar{a},\bar{h})}{\bar{a}}\right).$ (172)
 Letting $n=\zeta k$ and plugging (168) back into (166) and also evaluating the MGFs of $M$ and $L$ , we obtain
-(158) $P(Y^{(k)}>-n)\leq f(\bar{a},\bar{h})e^{(\zeta\nu^{*}-\xi(\bar{a},\bar{h}))k}$
+$P(Y^{(k)}>-n)\leq f(\bar{a},\bar{h})e^{(\zeta\nu^{*}-\xi(\bar{a},\bar{h}))k}$ (173)
 where
-(159) $f(\alpha,\eta)=\frac{e^{6\gamma}\gamma^{2}(\alpha-\eta+\alpha\eta)^{2}}{((\alpha+\gamma)(\gamma-\eta)+\alpha\eta e^{\gamma})^{2}}$
+$f(\alpha,\eta)=\frac{e^{5\gamma}\gamma^{2}(\alpha-\eta+\alpha\eta)^{2}}{((\alpha+\gamma)(\gamma-\eta)+\alpha\eta e^{\gamma})^{2}}$ (174)
 and
-(160) $\xi(\alpha,\eta)$ $=-\gamma-\ln\frac{\eta\alpha}{(\alpha+\gamma)(\eta-\gamma)}$
-(161) $=\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma.$
+$\xi(\alpha,\eta)$ $=-\gamma-\ln\frac{\eta\alpha}{(\alpha+\gamma)(\eta-\gamma)}$ (175)
+$=\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma.$ (176)
 To strike a balance between the exponents of $P(Y^{(k)}>-\zeta k)$ and $P(X^{(k)}\geq\zeta k)$, we let
-(162) $\zeta\nu^{*}-\xi(\bar{a},\bar{h})=\zeta\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}).$
+$\zeta\nu^{*}-\xi(\bar{a},\bar{h})=\zeta\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}).$ (177)
 It gives us
-(163) $\zeta=\frac{\xi(\bar{a},\bar{h})}{\ln(1+\frac{\gamma}{\bar{a}})-\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})},$
+$\zeta=\frac{\xi(\bar{a},\bar{h})}{\ln(1+\frac{\gamma}{\bar{a}})-\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})},$ (178)
 such that
-(164) $P(Y^{(k)}>-\zeta k)\leq f(\bar{a},\bar{h})(e^{-c(\bar{a},\bar{h})})^{k},$
+$P(Y^{(k)}>-\zeta k)\leq f(\bar{a},\bar{h})(e^{-c(\bar{a},\bar{h})})^{k},$ (179)
 and
-(165) $P(X^{(k)}\geq\zeta k)\leq(1-\delta)(e^{-c(\bar{a},\bar{h})})^{k},$
+$P(X^{(k)}\geq\zeta k)\leq(1-\delta)(e^{-c(\bar{a},\bar{h})})^{k},$ (180)
 with
-(166) $c(\bar{a},\bar{h})$ $=\xi(\bar{a},\bar{h})-\nu^{*}\zeta$
-(167) $=\xi(\bar{a},\bar{h})(1-\ln(1+\frac{\gamma}{\bar{a}})(\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}))^{-1})^{-1},$
+$c(\bar{a},\bar{h})$ $=\xi(\bar{a},\bar{h})-\nu^{*}\zeta$ (181)
+$=\xi(\bar{a},\bar{h})(1-\ln(1+\frac{\gamma}{\bar{a}})(\ln(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}))^{-1})^{-1},$ (182)
 which can be seen from the exponential parameter of (173).
 Denote
-(168) $b(\bar{a},\bar{h})$ $=f(\bar{a},\bar{h})+1-\delta$
-(169) $=\frac{e^{8\gamma}\gamma^{2}(\bar{a}-\bar{h}+\bar{a}\bar{h})^{2}}{((\bar{a}+\gamma)(\gamma-\bar{h})+\bar{a}\bar{h}e^{\gamma})^{2}}+1-\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)e^{-\bar{a}}.$
+$b(\bar{a},\bar{h})$ $=f(\bar{a},\bar{h})+1-\delta$ (183)
+$=\frac{e^{5\gamma}\gamma^{2}(\bar{a}-\bar{h}+\bar{a}\bar{h})^{2}}{((\bar{a}+\gamma)(\gamma-\bar{h})+\bar{a}\bar{h}e^{\gamma})^{2}}+1-\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)e^{-\bar{a}}.$ (184)
 We have
-(170) $P(X^{(k)}\geq\zeta k)+P(Y^{(k)}>-\zeta k)\leq b(\bar{a},\bar{h})e^{-c(\bar{a},\bar{h})k},$
+$P(X^{(k)}\geq\zeta k)+P(Y^{(k)}>-\zeta k)\leq b(\bar{a},\bar{h})e^{-c(\bar{a},\bar{h})k},$ (185)
 which completes the proof of Theorem 2.
@@
 Appendix E Bounding the Lead
-The lead increases by at most one upon the mining of an A-block and decreases by at least one upon the mining of a jumper unless the lead is already zero, in which case it stays at 0 with the mining of a jumper.
+The lead increases by at most one upon the mining of an A-block and decreases by at least one upon the publication of a jumper unless the lead is already zero, in which case it stays at 0 with the publication of a jumper.
 Note that $L_{0}=0$.
 On every sample path, there exists $u\in[0,s]$ with $L_{u}=0$ such that
-(171) $L_{s}-L_{u}$ $\leq A_{u,s}-J_{u,s}.$
+$L_{s}-L_{u}$ $\leq A_{u,s}-J_{u,s-\Delta}.$ (186)
 If $L_{s}=0$, $u=s$ satisfies (186). If $L_{s}>0$, $u$ is in the last interval of zero lead before $s$. In general, $L_{u}=0$ and (186) imply that
-(172) $L_{s}$ $\leq\sup_{q\in[0,s]}\{A_{q,s}-J_{q,s}\}$
-(173) $\leq\sup_{q\in[0,s]}\{A_{q,s}-P_{q+\Delta,s}\}$
-(174) $\leq\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{s-r,s}\}+A_{s-2\Delta,s}.$
+$L_{s}$ $\leq\sup_{q\in[0,s]}\{A_{q,s}-J_{q,s-\Delta}\}$ (187)
+$\leq\sup_{q\in[0,s]}\{A_{q,s}-P_{q+\Delta,s-\Delta}\}$ (188)
+$\leq\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{s-r,s-\Delta}\}+A_{s-2\Delta,s}.$ (189)
 where we have invoked Lemma 13 to obtain (188) and also used a change of variable $r=s-\Delta-q$ to obtain (189).
-Let us introduce a reverse pacer process, $P^{s}_{0,r}$, which denotes the number of pacers mined during $(s-r,s]$, i.e.,
-(175) $P^{s}_{0,r}=P_{s-r,s}.$
-Due to the nature of the pacer process, one can show that $P^{s}_{0,r}$ stochastically dominates $P_{0,r-\Delta}$ as a process.
+Let us introduce a reverse pacer process, $P^{s-\Delta}_{0,r}$, which denotes the number of pacers mined during $(s-r,s-\Delta]$, i.e.,
+$P^{s-\Delta}_{0,r}=P_{s-r,s-\Delta}.$ (190)
+Due to the nature of the pacer process, one can show that $P^{s-\Delta}_{0,r}$ stochastically dominates $P_{0,r-\Delta}$ as a process.
 Since $A$ is a memoryless Poisson process, $A_{s-\Delta-r,s-2\Delta}$ is statistically the same process as $A_{0,r-\Delta}$.
@@
 $P(L_{s}\geq x)$
-(176) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{0,r-\Delta}\}+A_{s-2\Delta,s}\geq x)$
-(177) $\leq P(\sup_{r\in[0,s-\Delta)}\{A_{0,r-\Delta}-P_{0,r-\Delta}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
-(178) $\leq P(\sup_{q\geq 0}\{A_{0,q}-P_{0,q}\}+A^{\prime}_{t,t+2\Delta}\geq x)$
-(179) $=P(L+A^{\prime}_{t,t+2\Delta}\geq x).$
+$\leq P\bigg{(}\sup_{r\in[0,s-\Delta)}\{A_{s-\Delta-r,s-2\Delta}-P_{0,r-\Delta}\}+A_{s-2\Delta,s}\geq x\bigg{)}$ (191)
+$\leq P\bigg{(}\sup_{r\in[0,s-\Delta)}\{A_{0,r-\Delta}-P_{0,r-\Delta}\}+A^{\prime}_{t,t+2\Delta}\geq x\bigg{)}$ (192)
+$\leq P\bigg{(}\sup_{q\geq 0}\{A_{0,q}-P_{0,q}\}+A^{\prime}_{t,t+2\Delta}\geq x\bigg{)}$ (193)
+$=P(L+A^{\prime}_{t,t+2\Delta}\geq x).$ (194)
 Appendix F Proof of Corollary 3
-This proof follows Appendix D.1. It suffices to show that for every $\epsilon>0$, there exists $c>0$ such that
-(180) $P(X^{(k)}\geq ck)$ $\leq\epsilon/2$
-(181) $P(Y^{(k)}>-ck)$ $\leq\epsilon/2$
+This proof follows Appendix D-A. It suffices to show that for every $\epsilon>0$, there exists $c>0$ such that
+$P(X^{(k)}\geq ck)$ $\leq\epsilon/2$ (195)
+$P(Y^{(k)}>-ck)$ $\leq\epsilon/2$ (196)
 for sufficiently large $k$. By Theorem 11, (195) holds for large enough $k$ for every $c>0$. The remainder of this appendix is devoted to the proof of (196).
 By (162) and the Chernoff bound, we have for $\nu>0$,
-(182) $P(Y^{(k)}>-ck)$ $\leq P(L+M+A_{0,T+6\Delta}>(1-c)k)$
-(183) $\leq E[e^{\nu(M+L+A_{0,T+6\Delta}+(c-1)k)}]$
-(184) $=E[e^{\nu(A_{0,T+6\Delta}+(c-1)k)}]E[e^{\nu M}]E[e^{\nu L}]$
+$P(Y^{(k)}>-ck)$ $\leq P(L+M+A_{0,T+3\Delta}>(1-c)k)$ (197)
+$\leq E[e^{\nu(M+L+A_{0,T+3\Delta}+(c-1)k)}]$ (198)
+$=E[e^{\nu(A_{0,T+3\Delta}+(c-1)k)}]E[e^{\nu M}]E[e^{\nu L}]$ (199)
 where $E[e^{\nu L}]=E[e^{\nu M}]=\mathcal{E}(e^{\nu})$ is finite for every $\nu>0$ by Lemma 4, where $\mathcal{E}$ is defined as (11).
-Only the first term on the right hand side of (184) depends on $k$. Given $t\geq 0$, $A_{0,t}$ is a Poisson random variable with mean $at$, whose MGF is
-(185) $E[e^{\nu A_{0,t}}]=e^{(e^{\nu}-1)at}.$
+Only the first term on the right-hand side of (199) depends on $k$. Given $t\geq 0$, $A_{0,t}$ is a Poisson random variable with mean $at$, whose MGF is
+$E[e^{\nu A_{0,t}}]=e^{\theta t}$ (200)
+where we have defined
+$\theta=\left(e^{\nu}-1\right)a$ (201)
 […]
-(186) $E[e^{\eta Z}]=\frac{1}{1-\eta/h}.$
-Using (185) and (186), we can write
-$E[$ $e^{\nu(A_{0,T+4\Delta}+(c-1)k)}]$
-(187) $=E[E[e^{\nu(A_{0,T+6\Delta})}|T]]\cdot e^{\nu(c-1)k}$
-(188) $=E[e^{(e^{\nu}-1)a(T+6\Delta)}]\cdot e^{\nu(c-1)k}$
-(189) $=E[e^{(e^{\nu}-1)aT}]\cdot e^{(e^{\nu}-1)6a\Delta}\cdot e^{\nu(c-1)k}$
-(190) $=\left(E[e^{(e^{\nu}-1)aZ}]\right)^{k}\cdot e^{(e^{\nu}-1)(6+k)a\Delta}\cdot e^{\nu(c-1)k}$
-(191) $=e^{(e^{\nu}-1)6a\Delta}\left(\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}\right)^{k}$
-(192) $=e^{(e^{\nu}-1)6a\Delta}(g(\nu,c))^{k}$
+$E[e^{\eta Z}]=\frac{1}{1-\eta/h}.$ (202)
+Using (200), (202), and the definition of $T$, we can write
+$E[$ $e^{\nu(A_{0,T+3\Delta}+(c-1)k)}]$
+$=E[E[e^{\nu(A_{0,T+3\Delta})}|T]]\cdot e^{\nu(c-1)k}$ (203)
+$=E[e^{\theta(T+3\Delta)}]\cdot e^{\nu(c-1)k}$ (204)
+$=\left(E[e^{\theta Z}]\right)^{k}\cdot e^{\theta(3+k)\Delta}\cdot e^{\nu(c-1)k}$ (205)
+$=e^{3\theta\Delta}(g(\nu,c))^{-k}$ (206)
 where
-(193) $g(\nu,c)$ $=\frac{e^{(e^{\nu}-1)a\Delta}e^{\nu(c-1)}}{1-(e^{\nu}-1)a/h}$
-(194) $=\frac{e^{(e^{\nu}-1)\bar{a}}e^{\nu(c-1)}}{1-(e^{\nu}-1)\bar{a}/\bar{h}}$
-where $\bar{a}=a\Delta$ and $\bar{h}=h\Delta$.
-In the remainder of this appendix, we will specify $\nu$ and $c$ as functions of $(\bar{a},\bar{h})$ which satisfy
-(195) $g(\nu,c)<1,$
-so that the right hand side of (192) and thus also of (184) vanishes as $k\to\infty$. Throughout this appendix, we denote
-(196) $c=\frac{1}{4}\left(1-\bar{a}\frac{1+\bar{h}}{\bar{h}}\right)$
-(197) $\theta$ $=\bar{a}(e^{\nu}-1).$
-We have $c>0$ for mining rate and propagation delay satisfy the fault tolerance as (1). It is also quite obvious that $12c<$ 1. Now we choose
-(198) $\theta=\frac{1}{2}\min\{1,\frac{\bar{a}}{2},\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}\}$
-which is positive and yields the corresponding $\nu=\ln(1+\frac{\theta}{a})>0$. We rewrite (193) into
-(199) $g(\nu,c)=\frac{e^{\theta}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
-By (198), we have $\theta<1$ and $\frac{\theta}{\bar{a}}<\frac{1}{2}$, so that
-(200) $g(\nu,c)$ $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+\frac{\theta}{\bar{a}})^{1-c}}$
-(201) $\leq\frac{1+\theta+\theta^{2}}{(1-\frac{\theta}{\bar{h}})(1+(1-2c)\frac{\theta}{\bar{a}})}$
-(202) $=\frac{1+\theta+\theta^{2}}{1+\theta+\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}}$
-where we have used the fact that $(1-2c)/\bar{a}-1/\bar{h}=1+2c/\bar{a}$. Since $\theta$ is strictly less than $\frac{c\bar{h}}{\bar{a}\bar{h}+1-2c}$, we have
-(203) $\theta^{2}<\frac{2c}{\bar{a}}\theta-\frac{1-2c}{\bar{a}\bar{h}}\theta^{2}.$
-Together, (202) and (203) guarantee (195). With the chosen $\nu$ and $c$, for sufficiently large $k$, (181) holds. The proof of Corollary 3 is thus complete.
+$g(\nu,c)$ $=\left(1-\frac{\theta}{h}\right)\left(1+\frac{\theta}{a}\right)^{1-c}e^{-\theta\Delta}.$ (207)
+We let
+$c=\frac{a}{2}\left(\frac{1}{a}-\frac{1}{h}-\Delta\right),$ (208)
+which is strictly positive when (7) holds, i.e., when the parameters are within the fault tolerance. Using Taylor series expansion, we have
+$g(\nu,c)$ $=1+\left(\frac{1-c}{a}-\frac{1}{h}-\Delta\right)\theta+o(\theta)$ (209)
+$=1+\frac{1}{2}\left(\frac{1}{a}-\frac{1}{h}-\Delta\right)\theta+o(\theta).$ (210)
+The coefficient of the first-order term in (210) is strictly positive due to (7).
+Hence there exists $\theta>0$ for which $g(\nu,c)>1$, so that the right hand side of (206) and thus also the right hand side of (199) vanish as $k\to\infty$.
+For sufficiently large $k$, (196) holds.
+The proof of Corollary 3 is thus complete.
 Appendix G Proof of Theorem 5
@@
 Starting from (162) in Appendix D, we use the law of total probability to write
-$P($ $A_{0,T+6\Delta}\leq-n+k-(M+L)))$
-$=$ $\sum_{i=0}^{k-n+1}P(L=i)(\sum_{j=0}^{k-n+1-i}P(M=j)$
-(204) $\quad P(A_{0,T+6\Delta}\leq-n+k-(i+j)))$
-$=$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
+$P($ $A_{0,T+3\Delta}\leq-n+k-(M+L))$
+$=$ $\sum_{i=0}^{k-n}P(L=i)(\sum_{j=0}^{k-n-i}P(M=j)P(A_{0,T+3\Delta}\leq-n+k-(i+j)))$ (211)
 […]
-$=$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
 […]
 Note the pdf of the $k$-st new pacer’s mining time $T$ , $f_{T}(t)$ can be driven by taking the derivative of its cdf , which is
-(207) $F_{2}(t-k\Delta;k,h),$
-where $F_{2}$ denotes the cdf of Erlang distribution. Now we only have to calculate the last term $P(A^{\prime}_{0,t+6\Delta}\leq-n+k-(i+j))$, which equals
-(208) $F_{1}(-n+k-(i+j);a(t+6\Delta))).$
+$F_{2}(t-k\Delta;k,h),$ (214)
+where $F_{2}$ denotes the cdf of Erlang distribution. Now we only have to calculate the last term $P(A_{0,t+3\Delta}\leq-n+k-(i+j))$, which equals
+$F_{1}(-n+k-(i+j);a(t+3\Delta))).$ (215)
 Finally, we have
 $P(Y^{(k)}\leq-n)$
-$\geq$ $\sum_{i=0}^{k-n+1}P(L=i)\sum_{j=0}^{k-n+1-i}P(M=j)$
 […]
-$=$ $\sum_{i=0}^{k-n+1}e(i)\sum_{j=0}^{k-n+1-i}e(j)\int_{0}^{\infty}F_{1}(-n+k-(i+j);a(t+6\Delta))$
-(210) $\cdot f_{2}(t-k\Delta;k,h)\,dt.$
+[…]
 Thus, according to (52), the probability of violation is upper bounded by
-(211) $\inf_{n}\{2-P(Y^{(k)}\leq-n)-P(X^{(k)}\leq n-1)\},$
+$\inf_{n}\{2-P(Y^{(k)}\leq-n)-P(X^{(k)}\leq n-1)\},$ (218)
 which equals (18). This completes the proof of Theorem 5.
 Appendix H Proof of Theorem 6
-Consider the same attack as the private-mining attack in Definition 14 except that we let the propagation delays of all H-blocks be zero before time $s-\Delta$ for simplicity.
-The adversary mines privately from $s-\Delta$ and delays all new H-blocks maximally.
-Let the first H-block mined after $s-\Delta$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on.
-Let $P^{\prime}_{r}$ denote the number of new pacers mined from $s-\Delta$ up until time $s+r-\Delta$.
+Consider the same attack as the private-mining attack in Definition 12 except that we let the propagation delays of all H-blocks be zero before time $s$ for simplicity.
+The adversary mines privately from $s$ and delays all new H-blocks maximally.
+Let the first H-block mined after $s$ be called the first new pacer (this takes an exponential time), then the first H-block mined at least $\Delta$ seconds after that be called the second new pacer, and so on.
+Let $P^{\prime}_{r}$ denote the number of new pacers mined from $s$ up until time $s+r$.
 Evidently, $P^{\prime}_{r}$ is identically distributed as $P_{0,r+\Delta}$.
-A violation occurs if there exists $d\geq s+\tau$ such that 1) the lead $L_{s-\Delta}$ plus the number of A-blocks $A_{s-\Delta,d}$ is no fewer than $P^{\prime}_{d-s+\Delta}$, the number of new pacers mined between $s-\Delta$ and $d-\Delta$ (any additional ones mined after that is not yet public); and 2) at least one pacer arrives by time $d$ to carry the target transaction, i.e., $P^{\prime}_{d-s+\Delta}\geq 1$.
+A violation occurs if there exists $d\geq s+\tau$ such that 1) the lead $L_{s}$ plus the number of A-blocks $A_{s,d}$ is no fewer than $P^{\prime}_{d-s-\Delta}$, the number of new pacers mined between $s$ and $d-\Delta$ (any additional ones mined after that is not yet public); and 2) at least one pacer arrives by time $d$ to be the target block, i.e., $P^{\prime}_{d-s-\Delta}\geq 1$.
 Hence the probability of violation is lower bounded by
-(212) $P$ $\left(\cup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right),$
-where $\tau$ is the time it takes for $k$ such new pacers to be mined after $s$. And the probability of (212) can be further bounded from below by
-$P\left(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
-(213) $\quad-P\left(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
-The second term is easily determined as
-(214) $P(\bigcup_{d\geq s+\tau}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{\tau+\Delta}=0)$
-(215) $=e^{-h(\tau+2\Delta)}.$
-Since $L_{s-\Delta}$, $A_{s-\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we have the following similar in Appendix D,
-$P(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
-(216) $=P(L_{s-\Delta}+\max_{d\geq s+\tau}\{A_{s-\Delta,d}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
-(217) $=P(L_{s-\Delta}+\max_{d\geq+\tau}\{A^{\prime}_{0,d-s-\Delta}-P^{\prime}_{d-s+\Delta}\}\geq 0)$
-(218) $\geq P(L_{s-\Delta}+A^{\prime}_{0,\tau-\Delta}+\max_{d\geq\tau}\{A^{\prime}_{\tau-\Delta,d-\Delta}-P^{\prime}_{0,d-\tau}\}-P^{\prime}_{\Delta}-k\geq 0)$
-(219) $=P(P^{\prime\prime}_{\Delta,2\Delta}\leq A^{\prime}_{0,\tau-\Delta}+M+L_{s-\Delta}-k),$
-where (218) is due to that there are $k$ new pacers during $(s,\tau]$. $M$ in (219) is defined in (143). $P^{\prime}$ and $P$ are independent identically distributed. Thus, (219) can be written as
-(220) $\begin{split}P&(\bigcup_{d\geq s+\tau}\left\{L_{s-\Delta}+A_{s-\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})\\ &\geq 1-P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-M-L_{s-\Delta}).\end{split}$
-Now, we want to calculate $P(A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)$ Conditioning on $P^{\prime\prime}_{\Delta,2\Delta}$,
-(221) $\begin{split}P(&A^{\prime}_{0,\tau-\Delta}\leq k-1+P^{\prime}_{\Delta,2\Delta}-L_{s-\Delta}-M)\\ &=P(A^{\prime}_{0,\tau-\Delta}\leq k-1-L_{s-\Delta}-M)f_{1}(0,h\Delta)\\ &\;\;+P(A_{0,\tau}\leq k-L_{s-\Delta}-M)(1-f_{1}(0,h\Delta)).\end{split}$
-We have
-$P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$
-(222) $=$ $\sum_{i=0}^{k}\sum_{j=0}^{k-i}P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)P(M=j)P(L_{s-\Delta}=i),$
-where $P(A^{\prime}_{0,\tau-\Delta}\leq k-L_{s-\Delta}-M)$ can be calculated as
+$P$ $\left(\cup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\geq 1\right\}\right),$ (219)
+where $\tau$ is the time it takes for $k$ such new pacers to be mined after $s$.
+For $k\geq 1$, $P^{\prime}_{d-s-\Delta}\geq 1$.
+Since $L_{s}$, $A_{s,d}$ and $P^{\prime}_{d-s-\Delta}$ are independent, we have the following similar in Appendix D,
+$P(\bigcup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
+$=P(L_{s}+\max_{d\geq s+\tau}\{A_{s,d}-P^{\prime}_{d-s-\Delta}\}\geq 0)$ (220)
+$=P(L_{s}+\max_{d\geq\tau}\{A^{\prime}_{0,d}-P^{\prime}_{d-\Delta}\}\geq 0)$ (221)
+$\geq P(L_{s}+A^{\prime}_{0,\tau}+\max_{d\geq\tau}\{A^{\prime}_{\tau,d}-P^{\prime}_{0,d-\tau}\}-k\geq 0)$ (222)
+$=P(0\leq A^{\prime}_{0,\tau}+M+L_{s}-k),$ (223)
+where (222) is due to that there are $k$ new pacers during $(s,\tau]$. $M$ in (223) is defined in (158). $P^{\prime}$ and $P$ are i.i.d. Thus, (223) can be written as
+$\begin{split}P&(\bigcup_{d\geq s+\tau}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s+\Delta}\right\})\\ &\geq 1-P(A^{\prime}_{0,\tau}\leq k-1-M-L_{s}).\end{split}$ (224)
+Now, we want to calculate $P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$ We have
+$P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$
+$=$ $\sum_{i=0}^{k}\sum_{j=0}^{k-i}P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)P(M=j)P(L_{s}=i),$ (225)
+where $P(A^{\prime}_{0,\tau}\leq k-1-L_{s}-M)$ can be calculated as
 […]
 where $\tau$ has cdf $F_{2}(t-k\Delta;k,h)$, and
-(224) $P(A_{0,t-\Delta}\leq k-i-j)=F_{1}(k-i-j;a(t-\Delta)).$
+$P(A_{0,t}\leq k-1-i-j)=F_{1}(k-1-i-j;at).$ (227)
 Finally, the probability that the target transaction’s safety is violated is no large than (19). To ease computation, we can also relax the infinite integral to finite integral as
 […]
-[…]
-(225) $\quad+\bar{F}_{2}(T;k,h).$
 Appendix I Proof of Theorem 8
@@
 However, the starting time of the error events of interest is different than in the case of confirmation by depth.
-I.1. An Enlarged Violation Event
+I-A An Enlarged Violation Event
 Definition 18 (Unbalanced jumper).
@@
 Lemma 15.
-Suppose $r^{\prime}\geq r\geq s\geq 0$. If an $r$-credible chain and an $r^{\prime}$-credible chain disagree on a target transaction that is public at time $s$, then
-(226) $L_{s}+A_{s,r^{\prime}}\geq\hat{J}_{s,r}$
+Suppose $r^{\prime}\geq r\geq s\geq 0$. If an $r$-credible chain and an $r^{\prime}$-credible chain disagree on a target block that is mined by time $s$, then
+$L_{s}+A_{s,r^{\prime}}\geq\hat{J}_{s,r}$ (229)
 Proof.
@@
 We assume $s<r-\Delta$ in the remainder of this proof.
-Let block $b$ be the first H-block mined after time $s$.
-Since the target transaction is public at time $s$, Definition 10 implies that there must be disagreements on all heights of jumpers mined from $s$ to $r-\Delta$, i.e., on no fewer than $J_{s,r-\Delta}$ heights.
-Let us denote the highest honest block before time $s$ as height $\eta$.
+Since the target block is mined by time $s$, there must be disagreements on all heights of jumpers mined from $s$ to $r-\Delta$, i.e., on no fewer than $J_{s,r-\Delta}$ heights.
+Let us denote the public height at time $s$ as height $\eta$.
 Every unbalanced jumper mined during $(s,r-\Delta]$ must meet either one of the following two conditions:
@@
 The adversary needs to mine at least on $\hat{J}_{s,r}$ heights of unbalanced jumpers.
-Due to the pre-mining phase, at time $s$, the highest A-block has a height $\eta+L_{s}$, which is $L_{s}$ higher than the highest honest block, which means at most $L_{s}$ of A-blocks can be mined before time $s$ on the heights of unbalanced jumpers.
+Due to the pre-mining phase, at time $s$, the highest A-block has a height $\eta+L_{s}$, which is at most $L_{s}$ higher than the highest honest block, which means at most $L_{s}$ of A-blocks can be mined before time $s$ on the heights of unbalanced jumpers.
 The rest A-blocks can only be mined after time $s$, the number of which is $A_{s,r^{\prime}}$.
@@
 Fix $\epsilon\in(0,\Delta)$, barring the following event
-(227) $F_{s,t}^{\epsilon}=\bigcup_{d\geq t+s}\left\{L_{s}+A_{s,d}\geq J_{s,d-\epsilon-\Delta}-X\right\}$
-all $r^{\prime}$-credible chains agree on the target transaction for all $r^{\prime}\geq t+s$.
+$F_{s,t}^{\epsilon}=\bigcup_{d\geq t+s}\left\{L_{s}+A_{s,d}\geq J_{s,d-\epsilon-\Delta}-X\right\}$ (230)
+all $r^{\prime}$-credible chains agree on the target block for all $r^{\prime}\geq t+s$.
 Proof.
@@
 If an $r^{\prime}$-credible chain and an $r$-credible chain disagree, then we have the following by Lemma 15:
-(228) $L_{s}+A_{s,r^{\prime}}$ $\geq\hat{J}_{s,r}$
-(229) $\geq\hat{J}_{s,r^{\prime}-\epsilon}$
-(230) $\geq J_{s,r^{\prime}-\epsilon-\Delta}-X$
+$L_{s}+A_{s,r^{\prime}}$ $\geq\hat{J}_{s,r}$ (231)
+$\geq\hat{J}_{s,r^{\prime}-\epsilon}$ (232)
+$\geq J_{s,r^{\prime}-\epsilon-\Delta}-X$ (233)
 where (233) is due to that the number of balanced candidates mined during $(s,r^{\prime}-\epsilon]$ is smaller than $X$.
@@
 The lemma is then established by induction on $n$.
+∎
 By Lemma 13, we have
-(231) $J_{s,d-\epsilon-\Delta}-X$ $\geq P_{s+\Delta,d-\epsilon-\Delta}-X.$
+$J_{s,d-\epsilon-\Delta}-X$ $\geq P_{s+\Delta,d-\epsilon-\Delta}-X.$ (234)
 By (230) and (234), we have
-(232) $F_{s,t}^{\epsilon}\subset\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\}.$
+$F_{s,t}^{\epsilon}\subset\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\}.$ (235)
 Hence
-(233) $P(F_{s,t}^{\epsilon})$ $\leq P(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\})$
-(234) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}+X\geq 0),$
+$P(F_{s,t}^{\epsilon})$ $\leq P(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\})$ (236)
+$=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}+X\geq 0),$ (237)
 where equality is by the definition of the union.
 Let us define
-(235) $Y_{t}=\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}.$
+$Y_{t}=\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}.$ (238)
 We apply Lemma 12 to obtain
-(236) $P(F_{s,t}^{\epsilon})\leq\inf_{n\in\{1,2,...\}}\{P(X\geq n)+P(Y_{t}>-n)\}$
+$P(F_{s,t}^{\epsilon})\leq\inf_{n\in\{1,2,...\}}\{P(X\geq n)+P(Y_{t}>-n)\}$ (239)
 where it suffices to take the infimum over the natural number set only.
-I.2. Probability Bound
+I-B Probability Bound
 We next upper bound $P(Y_{t}>-n)$ using properties of the adversarial mining process, the jumper process, as well as Kroese’s results on the difference of renewal processes.
@@
 We define
-(237) $T=\min\{t\in[s,s+\Delta]:P_{t}=P_{s+\Delta}\},$
+$T=\min\{t\in[s,s+\Delta]:P_{t}=P_{s+\Delta}\},$ (240)
 which represents the last time a pacer is mined by $s+\Delta$.
@@
 We evaluate $P(Y_{t}\leq-n)$ by conditioning on and then averaging over $T$:
-(238) $P(Y_{t}\leq-n)=\mathbb{E}[P(Y_{t}\leq-n|T)]$
-(239) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{T,d-\epsilon-\Delta}\}\leq-n|T)]$
-(240) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-\epsilon-\Delta-T}\}\leq-n|T)].$
+$P(Y_{t}\leq-n)$
+$=\mathbb{E}[P(Y_{t}\leq-n|T)]$ (241)
+$=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{T,d-\epsilon-\Delta}\}\leq-n|T)]$ (242)
+$=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-\epsilon-\Delta-T}\}\leq-n|T)].$ (243)
 By (240), $d-\epsilon-\Delta-T\geq d-s-\epsilon-2\Delta$, so (243) implies
 $P(Y_{t}\leq-n)$
-(241) $\geq\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n|T)]$
-(242) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n)$
+$\geq\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n|T)]$ (244)
+$=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n)$ (245)
 where we drop conditioning on $T$ because all other random variables in (244) are independent of $T$.
@@
 $P(Y_{t}\leq-n)$
-(243) $\geq$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,q}-P^{\prime}_{0,q-\epsilon-2\Delta}\}\leq-n)$
-(244) $=$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime}_{t-\epsilon-2\Delta,q-\epsilon-2\Delta}\}\leq-n).$
+$\geq P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,q}-P^{\prime}_{0,q-\epsilon-2\Delta}\}\leq-n)$ (246)
+$=P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime}_{t-\epsilon-2\Delta,q-\epsilon-2\Delta}\}\leq-n).$ (247)
 We further replace $A^{\prime}_{t,q}$ by $A^{\prime\prime}_{0,q-t}$ in (247).
-We also note that $P^{\prime}_{t-2\Delta-\epsilon,q-2\Delta-\epsilon}$ is statistically dominated by $P^{\prime\prime}_{0,q-t}$.
+We also note that $P^{\prime}_{t-2\Delta-\epsilon,q-2\Delta-\epsilon}$ statistically dominates $P^{\prime\prime}_{0,q-t}$.
 So, we rewrite (247) as
 $\quad P(Y_{t}\leq-n)$
-(245) $\geq P\left(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
-(246) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+\sup_{q\geq t}\{A^{\prime\prime}_{0,q-t}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
-(247) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+M\leq-n\right)$
-(248) $=1-P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq L_{s}+A_{0,t}^{\prime}+M+n-1)$
+$\geq P\left(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$ (248)
+$=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+\sup_{q\geq t}\{A^{\prime\prime}_{0,q-t}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$ (249)
+$=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+M\leq-n\right)$ (250)
+$=1-P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq L_{s}+A_{0,t}^{\prime}+M+n-1)$ (251)
 where
-(249) $M=\sup_{r\geq 0}\{A^{\prime\prime}_{0,r}-P^{\prime\prime}_{0,r}\}.$
-As the maximum difference between two renewal processes, $M$ has the pmf (11) by (Kroese_difference, , Theorem 3.4) (see also (li2021close, , Sec. 6)). given by (11).
+$M=\sup_{r\geq 0}\{A^{\prime\prime}_{0,r}-P^{\prime\prime}_{0,r}\}.$ (252)
+As the maximum difference between two renewal processes, $M$ has the pmf (12) by [22, Theorem 3.4] (see also [6, Sec. 6]). given by (12).
 Thus, by (239) and (251), we have
-$P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
-(250) $+P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
+$\begin{split}P(F_{s,t}^{\epsilon})&\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.\end{split}$ (253)
 Taking the limit of $\epsilon\to 0$, we use continuity to obtain
-$\lim_{\epsilon\rightarrow 0}P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
-(251) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
+$\begin{split}\lim_{\epsilon\rightarrow 0}P(F_{s,t}^{\epsilon})&\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.\end{split}$ (254)
 The second term on the right-hand side of (254) is essentially the probability that the number of pacers mined during $(0,t-2\Delta]$ is fewer than the number of blocks mined by the adversary over a duration of $t$, plus the adversarial advantage after t, plus the adversary’s pre-mining gain, plus the number of balanced heights.
@@
 Hence (254) still holds with $L_{s}$ replaced by $L+A^{\prime}_{t,t+2\Delta}$ to yield
-$\lim_{\epsilon\to 0}P(F_{s,t}^{\epsilon})\leq$ $\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
-(252) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)\}$
+$\begin{split}\lim_{\epsilon\to 0}P(F_{s,t}^{\epsilon})\leq&\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)\}\end{split}$ (255)
 where $P^{\prime}$, $A^{\prime}$, $M$, and $L$ are independent.
@@
 We have in general
-(253) $P(P^{\prime}_{0,t}\leq i-1)$ $=P(i\Delta+X_{1}+\cdots+X_{i}>t)$
+$P(P^{\prime}_{0,t}\leq i-1)$ $=P(i\Delta+X_{1}+\cdots+X_{i}>t)$ (256)
 […]
-(254) $P(P^{\prime}_{0,t}\leq i-1)=1-F_{2}(t-i\Delta;i,h)$
+$P(P^{\prime}_{0,t}\leq i-1)=1-F_{2}(t-i\Delta;i,h)$ (257)
 where $F_{2}(\cdot;i,h)$ denotes the Erlang cdf with shape parameter $i$ and rate $h$.
@@
 $P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)$
-$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}P(A_{0,t+2\Delta}^{\prime}=i)P(M=j)P(L=k)$
-(255) $\cdot P(P_{0,t-2\Delta}^{\prime}\leq i+j+k+n-1)$
-$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)$
-(256) $\cdot(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h)),$
+$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}P(A_{0,t+2\Delta}^{\prime}=i)P(M=j)P(L=k)\cdot P(P_{0,t-2\Delta}^{\prime}\leq i+j+k+n-1)$ (258)
+$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\cdot(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h)),$ (259)
 where $f_{1}(\cdot;\omega)$ denotes the pmf of a Poisson distribution with mean $\omega$.
-(The cdf of $P^{\prime}_{0,t}$ has been evaluated in (li2021close, , Sec.
-6.1).) It is easy to see that (256) is equal to $G(n,t)$ defined in (25).
+(The cdf of $P^{\prime}_{0,t}$ has been evaluated in [6, Sec.
+6.1].) It is easy to see that (259) is equal to $G(n,t)$ defined in (26).
 By (255), the safety of the said transaction confirmed using latency $t$ can be violated with a probability no higher than
-(257) $\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+G(n,t)\},$
+$\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+G(n,t)\},$ (260)
 which is equal to (25).
@@
 $\quad P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)$
-$\leq 1-P(A_{0,t+2\Delta}^{\prime}\leq i_{0})P(M\leq j_{0})P(L\leq k_{0})$
-$\quad+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n,A_{0,t+2\Delta}^{\prime}\leq i_{0},$
-(258) $\quad M\leq j_{0},L\leq k_{0})$
-(259) $\leq\overline{G}(n,t),$
+$\leq 1-P(A_{0,t+2\Delta}^{\prime}\leq i_{0})P(M\leq j_{0})P(L\leq k_{0})+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n,A_{0,t+2\Delta}^{\prime}\leq i_{0},M\leq j_{0},L\leq k_{0})$ (261)
+$\leq\overline{G}(n,t),$ (262)
 where $\overline{G}(n,t)$ is calculated as (27).
 Finally, we have that the safety of a given transaction confirmed using latency $t$ can be violated with a probability no higher than
-$\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+\overline{G}(n,t)\}$
-(260) $=\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+\overline{G}(n,t)\}.$
+$\begin{split}&\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+\overline{G}(n,t)\}=\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+\overline{G}(n,t)\}.\end{split}$ (263)
 Appendix J Proof of Theorem 9
 In this section, we lower bound the probability of safety violation achieved by a specific attack.
-The attack is the same as described in Appendix H, except that all H-blocks mined after $s-2\Delta$ are maximally delayed by $\Delta$, instead of after $s-\Delta$.
-Accordingly, the first new pacer is the first H-block mined after $s-2\Delta$.
-Hence the probability that the adversary wins is lower bounded by
-(261) $P$ $\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right).$
+The attack is the same as described in Appendix H.
+The first new pacer is the first H-block mined after $s$.
+Thus the probability that the adversary wins is lower bounded by
+$P$ $\left(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\geq 1\right\}\right).$ (264)
 It is not difficult to see that (264) is further lower bounded by
-$P\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
-(262) $-P\left(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
-The second term in (J) is easily determined as
-(263) $P(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{t+\Delta}=0)$
-(264) $=e^{-h(t+2\Delta)}.$
-It remains to bound the first term in (J).
-Since $L_{s-2\Delta}$, $A_{s-2\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we can replace $A_{s-2\Delta,d}$ with identically distributed and independent $A^{\prime}_{0,d-s+2\Delta}$ and replace $P^{\prime}_{d-s+\Delta}$ with $P^{\prime\prime}_{0,d-s+2\Delta}$ in (261) to obtain
-$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
-(265) $=P(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A^{\prime}_{0,d-s+2\Delta}\geq P^{\prime\prime}_{0,d-s+2\Delta}\right\})$
-(266) $=P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime}_{t+\Delta,q+2\Delta}\geq P^{\prime\prime}_{0,t+2\Delta}+P^{\prime\prime}_{t+2\Delta,q+2\Delta}\right\}).$
-Evidently, $A^{\prime}_{t+\Delta,q+2\Delta}$ is identically distributed as $A^{\prime\prime}_{0,q-t+\Delta}$.
-Moreover, as a delayed renewal processes with identical inter-arrival times, $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ has statistically a larger delay than $P^{\prime}_{\Delta,q-t+\Delta}$.
-Thus $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ is stochastically dominated by $P^{\prime\prime}_{\Delta,q-t+\Delta}=P^{\prime\prime}_{0,q-t+\Delta}$.
+$\begin{split}P\left(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\}\right)-P\left(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s-\Delta}=0\right\}\right).\end{split}$ (265)
+The second term in (265) is easily determined as
+$P(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s-\Delta}=0\right\})$ $=P(P^{\prime}_{t-\Delta}=0)$ (266)
+$=e^{-h(t-\Delta)}.$ (267)
+It remains to bound the first term in (265).
+Since $L_{s}$, $A_{s,d}$ and $P^{\prime}_{d-s-\Delta}$ are independent, we can replace $A_{s,d}$ with identically distributed and independent $A^{\prime}_{0,d-s}$ and replace $P^{\prime}_{d-s-\Delta}$ with i.i.d.
+$P^{\prime\prime}_{0,d-s}$ in (264) to obtain
+$P$ $(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
+$=P(\bigcup_{d\geq s+t}\left\{L_{s}+A^{\prime}_{0,d-s}\geq P^{\prime\prime}_{0,d-s}\right\})$ (268)
+$=P(\bigcup_{q\geq t}\left\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}\geq P^{\prime\prime}_{0,t+\Delta}+P^{\prime\prime}_{t+\Delta,q}\right\}).$ (269)
+Evidently, $A^{\prime}_{t,q}$ is identically distributed as $A^{\prime\prime}_{0,q-t}$.
+Moreover, as a delayed renewal processes with identical inter-arrival times, $P^{\prime\prime}_{t+\Delta,q}$ has statistically a larger delay than $P^{\prime}_{\Delta,q-t}$.
+Thus $P^{\prime\prime}_{t+\Delta,q}$ is stochastically dominated by $P^{\prime\prime}_{\Delta,q-t}=P^{\prime\prime}_{0,q-t}$.
 Hence (269) implies that
-$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
-(267) $\geq P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime\prime}_{0,q-t+\Delta}\geq P^{\prime}_{0,t+2\Delta}+P^{\prime\prime}_{0,q-t+\Delta}\right\})$
-(268) $\geq P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$
+$P$ $(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
+$\geq P(\bigcup_{q\geq t}\left\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}\geq P^{\prime}_{0,t+\Delta}+P^{\prime\prime}_{0,q-t}\right\})$ (270)
+$\geq P(P^{\prime}_{0,t+\Delta}\leq A^{\prime}_{0,t}+M+L_{s})$ (271)
 where $M$ is defined in (252).
 The pmf of $M$ equals (12).
-Similar to (li2021close, ), $P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$ can be calculated using the pmfs of the relevant independent random variables.
+Similar to [6], $P(P^{\prime}_{0,t+\Delta}\leq A^{\prime}_{0,t}+M+L_{s})$ can be calculated using the pmfs of the relevant independent random variables.
 Thus, given a latency $t$, the safety of a given transaction can be violated with a probability no less than given in (28).